### v2

# Five-loop Anomalous Dimensions of Cubic Scalar Theory from Operator Product Expansion

Abstract
Recently, an Operator Product Expansion method has been proposed for renormalization in quantum field theories. This method extracts UV divergences from two-point propagator-type integrals, and it is believed to be applicable to generic QFTs. In the present work, we apply the OPE method to the renormalization of scalar theory with cubic interactions in six-dimensions. We have successfully computed the 5-loop corrections to the anomalous dimensions of the $\phi^{Q}$ operator, as well as the large $N$ expansion of scaling dimensions at the Wilson-Fisher fixed point to $1/N^{5}$ order. This computation sets a new loop-order record for the anomalous dimension of the $\phi^{Q}$ operator in cubic scalar theory, and further demonstrates the efficiency and capacity of the OPE method in renormalization.
Keywords:
Anomalous Dimension, OPE, Cubic Scalar Theory, Renormalization
1 Introduction
During the development of the renormalization group method, scalar field theories have served as an ideal playground for testing multi-loop computation techniques and providing insights into more complicated field theories such as non-Abelian gauge theories. Yet the renormalization of scalar field theories is itself of great importance, as explored in Wilson’s theory of critical phenomena Wilson:1971dc , that in certain statistical systems, the critical exponents near second-order phase transitions are related to the anomalous dimensions of the field or composite operators constructed from fields PELISSETTO2002549 ; Moshe:2003xn . Two notable scalar field theories are the six-dimensional cubic scalar field theory and the four-dimensional quartic scalar theory, both of which are renormalizable theories. These two theories belong to the same $O(N)$ universality class, and it has been conjectured that the $O(N)$ quartic scalar theory at its UV fixed point is equivalent to the $O(N)$ cubic scalar theory at its IR fixed point Fei:2014yja ; Fei:2014xta , namely they approach to the same critical theory in the fixed points. Verifying this conjecture with high-loop precision for the scaling dimensions of fields or composite operators represents an intriguing problem, but it demands perturbative computation of renormalization functions to higher-loop orders. Because of the technical advancements, recently there has been a growing interest in pushing renormalization computations to the next loop order.
It has been a long history for renormalization computations of quartic and cubic scalar theories, dating back to the early era studies of critical exponents up to $\epsilon^{2}$ orders for generalized Ising and Baxter models Wilson:1971dc , as well as the generalized Heisenberg model Wilson:1971vs . The 3-loop computations of the beta functions, field renormalization and mass renormalization of the quartic scalar theory were completed just one year after the 1-loop and 2-loop results BREZIN1973227 ; Brezin:1973igb , with 4-loop results following five years later Kazakov:1979ik . The full 5-loop renormalization of the quartic scalar theory required over a decade of collaborative efforts from numerous researchers Chetyrkin:1981jq ; Gorishnii:1983gp ; Kazakov:1983dyk ; Gorishnii:1983jr ; Kleinert:1991rg ; Adzhemyan:2013jra . Throughout this development, techniques such as the Gegenbauer polynomial $x$-space, IR-rearrangement Vladimirov:1979zm ; Chetyrkin:1980pr ; Caswell:1981ek , $R$-operation ($R^{\ast}$-operation) Chetyrkin:1982nn ; CHETYRKIN1984419 ; Larin:2002sc , as well as the modern Feynman diagram generating and Feynman integral reduction algorithms were refined and played the major roles. Early attempts were also made to push renormalization computations to higher-loop order, for instance the primitive divergences of the 7-loop beta function Broadhurst:1995km , or extend them to composite operators, for instance the 4-loop anomalous dimension of operators with derivatives Derkachov:1997pf . However, the complete 6-loop computation of the quartic scalar theory became feasible only approximately 15 years after the 5-loop result, firstly with the 6-loop field renormalization result Batkovich:2016jus , followed by the beta function and mass renormalization for the $O(1)$ model Kompaniets:2016hct , and subsequently for the $O(N)$ model Kompaniets:2017yct . A major breakthrough came with the development of the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which enabled the successful computation of the 7-loop beta function, mass renormalization, and field renormalization Schnetz:2016fhy , and even the 8-loop field renormalization function Schnetz:2025opm . The history of renormalization computations for the cubic scalar theory has closely paralleled that of the quartic scalar theory, which is not surprising since both have benefited from the same technical advancements. However, due to greater complexity, progress in the cubic scalar theory has consistently lagged by 1 or 2 loops compared to its quartic counterpart at similar stages. Early 1-loop and 2-loop analysis of the cubic scalar theory revealed its asymptotic freedom Macfarlane:1974vp ; Ma:1975vn . Over the subsequent 40 years, 3-loop deAlcantaraBonfim:1980pe ; deAlcantaraBonfim:1981sy ; Fei:2014xta and 4-loop Gracey:2015tta renormalization computations were gradually completed, with the 5-loop result finished only very recently Kompaniets:2021hwg ; Borinsky:2021jdb . The frontier record once again came from the graphical function method, by which the 6-loop beta function, field renormalization and mass renormalization functions of the cubic scalar theory have been successfully acquired this very year Schnetz:2025wtu .
Besides the computations of the beta functions and anomalous dimensions of fields and mass, the renormalization of composite operators constitutes an equal important yet more challenging problem. Phenomenologically, this is relevant to specific critical exponents in phase transition problems. One class of composite operators of interest is the fixed-charge symmetric traceless operator $\phi^{Q}$. In statistical physics, the anomalous dimensions of $\phi^{Q}$ for certain values of $Q$ are relevant to phase transitions in diverse systems PELISSETTO2002549 , such as liquid crystals, graphite-intercalation compounds, polymers, etc. DePrato:2003yd ; Calabrese:2004ca . Interest in computing the scaling dimensions of the $\phi^{Q}$ operator in cubic and quartic scalar theories grows from 2019, when the semi-classical method is introduced to compute the scaling dimension of $\phi^{Q}$ for generic $Q$ at the large $Q$ limit. This leads to all-loop results at leading and sub-leading orders in the $Q$ expansion for the quartic scalar theory Arias-Tamargo:2019xld ; Badel:2019oxl ; Antipin:2020abu ; Giombi:2020enj and the cubic scalar theory Arias-Tamargo:2020fow ; Antipin:2021jiw . Progress in the semi-classical approach stimulates the growing attention to the perturbative method approach towards renormalization computation of the $\phi^{Q}$ operator, aiming to validate the semi-classical results and also to obtain results for all orders of $Q$. For the quartic scalar theory, the scaling dimension of the $\phi^{Q}$ operator has been computed perturbatively at 4-loop Jack:2021ypd and 5-loop order Jin:2022nqq . While the 6-loop result is also available Bednyakov:2022guj , in which the data of $Q=1,\ldots,5$ was computed via perturbative Feynman diagram method and the data of $Q=6,7$ was imported from the semi-classical method. Concurrently, for the cubic scalar theory, perturbative computations have been carried out at 3-loop Jack:2021ziq and 4-loop order Huang:2024hsn .
High-loop order renormalization computations would never be possible without the continuous evolution of both traditional and modern techniques. One of the important techniques to mention is the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which takes the advantage of computing scalar integrals in coordinate space. It pushes renormalization computations to 8-loops for the four-dimensional quartic scalar theory and 6-loops for the six-dimensional cubic scalar theory, and for the first time, with some tricks it also pushes the renormalization computations to 7-loops in the Gross-Neveu model involving spinning particles Huang:2025ree . The graphical function method is designed to evaluate scalar integrals with no more than four external points, which is certainly not the case for integrals of correlation functions involving $\phi^{Q}$ operators. To address this, we have proposed an Operator Product Expansion (OPE) method for renormalization computations by analyzing the UV divergence structures of correlation functions in the large momentum expansion limit Huang:2024hsn . This OPE method has proven highly useful and efficient for determining anomalous dimensions Huang:2024hsn , and this fact has also been shown in the condensate calculations Marino:2024uco ; Liu:2025bqq . In the OPE framework, the momenta of the two operators are taken to the infinite limit. This can be considered pictorially as two hard fields propagating in a soft field background Huang:2016bmv ; Huang:2022qnx , reducing the integral evaluation to a single-scale problem characterized by the large momentum parameter. By identifying the contributing terms in the OPE, the essential UV divergences required for renormalization computations are found to be determined by two-point hard sub-graphs. Concurrently, the Wilson coefficients in the OPE are proportional to a combination of $Z$-factors and the contributions of these two-point hard sub-graphs. By requiring the UV finiteness of the Wilson coefficients, the $Z$-factors can be determined once the UV divergences of the hard graphs are known. This approach thus reduces renormalization computations to the evaluation of two-point propagator-type integrals. Note that the generality of OPE across any quantum field theories enables diverse applications of this OPE method, to theories with scalars or particles with spins, and also to the situations involving operator mixing and operators with derivatives, with Lorentz indices, etc.
In this work, we apply the OPE method developed in Huang:2024hsn to compute the 5-loop scaling dimension of the $\phi^{Q}$ operators in the six-dimensional $O(N)$ cubic scalar theory, thereby extending beyond the 4-loop result reported in that paper. This paper is structured as follows. In §2, we provide a concise review of the cubic scalar theory, its renormalization functions, and the OPE method for renormalization computation. Treatments for organizing and simplifying Feynman graphs of the correlation functions involving the $\phi^{Q}$ operator are also presented. The §3 is devoted to a detailed demonstration of the OPE method using 2-loop examples, showing how the field renormalization function, beta function and the anomalous dimensions of the $\phi^{Q}$ operator can be computed efficiently. In §4, general discussion for the computation of 5-loop anomalous dimension of $\phi^{Q}$ operator has been presented, and various 5-loop results can be found in this section and the Appendix. In §5, conclusion and discussion are presented.
2 Review of the theory and the method
In this section we review the six-dimensional $O(N)$ cubic scalar theory, with emphasis on its renormalization computations and the techniques for organizing and simplifying Feynman diagrams. To make this paper self-contained, we also provide a brief review of the OPE method for renormalization computations Huang:2024hsn , especially for its application in the context of cubic scalar theory.
2.1 The cubic scalar theory and its renormalization
The theory under consideration is the six-dimensional $O(N)$ cubic scalar theory, defined by the Lagrangian,
$\mathcal{L}=\frac{1}{2}(\partial\phi_{i})^{2}+\frac{1}{2}(\partial\sigma)^{2}-\frac{g}{2!}\sigma\phi_{i}^{2}-\frac{h}{3!}\sigma^{3}~,$ (1)
where $\sigma$ and $\phi_{i},i=1,\ldots,N$ denote two types of real scalar fields, with the $\phi_{i}$ fields transforming as an $O(N)$ vector. This theory is perturbatively renormalizable in six-dimensional space-time. The primary focus of this paper is the renormalization analysis of the symmetric traceless $\phi^{Q}$ operator, defined as,
$\phi^{Q}:=T_{i_{1}\cdots i_{Q}}\phi^{i_{1}}\cdots\phi^{i_{Q}}~,$ (2)
where $T_{i_{1}\cdots i_{Q}}$ is a symmetric traceless tensor and $Q$ is a fixed charge number. We will work in the $\overline{\mbox{MS}}$ regularization scheme tHooft:1973mfk in $d=(6-\epsilon)$-dimension. The classical scaling dimension of $\phi^{Q}$ is Q(d/21)Q(d/2-1), while the total scaling dimension is
$\Delta_{Q}=Q\left(\frac{d}{2}-1\right)+\gamma_{Q}+Q\gamma_{\phi}~.$ (3)
The difference between the total scaling dimension and the classical scaling dimension is defined as the anomalous dimension of the $\phi^{Q}$ operator. The gamma function $\gamma_{Q}$ of $\phi^{Q}$ operator can be computed perturbatively by Feynman diagrams of inserting the $\phi^{Q}$ operator in the 1PI Green function with $Q$ external legs. However, explicit evaluation of these correlation functions is complicated by the cumbersome $O(N)$ tensor structures inherent to the operator, which cause the number of terms in Feynman rules to grow exponentially with increasing $Q$. To address this problem, it is convenient to introduce an auxiliary $O(N)$ zero-norm complex vector $q$ satisfying $^{2}=0, qq$$\cdot\bar{q}=1$, and define the tensor $T$ as Antipin:2020abu ; Jin:2022nqq
$T_{i_{1}\ldots i_{Q}}:=\frac{1}{Q!}q_{i_{1}}q_{i_{2}}\cdots q_{i_{Q}}~.$ (4)
This allows us to redefine the $\phi^{Q}$ operator in terms of a $\varphi^{Q}$ operator,
$\phi^{Q}:=\frac{1}{Q!}\varphi^{Q}~,$ (5)
where the complex field $\varphi$ is a redefinition of the scalar field $\varphi:=q\cdot\phi$, with its complex conjugate field $\bar{\varphi}:=\bar{q}\cdot\phi$. This definition effectively strips $O(N)$ indices from the correlation functions, while the anomalous dimension of $\phi^{Q}$ operator is equivalent to that of $\varphi^{Q}$ operator. The latter can now be computed by $\varphi^{Q}$-$Q\varphi$ correlation functions, which are free of $O(N)$ index structures. In the $\varphi$-field representation, the $\sigma\phi^{2}$ vertices along the line of the operator and the external legs are reformulated as $\sigma\varphi\bar{\varphi}$ vertices, while other $\sigma\phi^{2}$ vertices appearing in loop structures of $\varphi$ fields contribute factors of $N$ after summing over color indices. Due to charge conservation, scalar charge flows along $\varphi$ fields in the corresponding Feynman diagrams.
For purpose of renormalization computation, following standard procedures, we introduce renormalized fields as,
$\varphi_{0}=Z^{\frac{1}{2}}_{\varphi}\varphi~~~,~~~\sigma_{0}=Z^{\frac{1}{2}}_{\sigma}\sigma~,$ (6)
and renormalized coupling constants as,
$g_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{g}g~~~,~~~h_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{h}h~~~\mbox{with}~~~\widetilde{Z}_{g}=Z_{g}Z^{-\frac{1}{2}}_{\sigma}Z^{-1}_{\varphi}~~~,~~~\widetilde{Z}_{h}=Z_{h}Z^{-\frac{3}{2}}_{\sigma}~,$ (7)
where $g_{0},h_{0}$ denote bare coupling constants and $\varphi_{0},\sigma_{0}$ denote bare fields. We also define the renormalized operator $\varphi^{Q}_{R}=Z_{\varphi^{Q}}\varphi_{0}^{Q}$. Since $\varphi_{0}^{Q}=Z_{\varphi}^{\frac{Q}{2}}\varphi^{Q}$, we get
$\varphi_{R}^{Q}=\widetilde{Z}_{\varphi^{Q}}\varphi^{Q}~~~\mbox{with}~~~\widetilde{Z}_{\varphi^{Q}}=Z_{\varphi^{Q}}Z_{\varphi}^{\frac{Q}{2}}~.$ (8)
Then $\widetilde{Z}_{\varphi^{Q}}$ can be assigned as the Feynman rule of the $\varphi_{R}^{Q}$-$Q\varphi$ vertex, and the anomalous dimension of the $\varphi^{Q}$ operator can be determined by the $Z_{\varphi^{Q}}$-factor.
Provided that the $Z_{\varphi}$, $Z_{\sigma}$, $Z_{g}$ and $Z_{h}$ have been determined, we can compute beta functions via the standard field theory method. The beta functions $\beta(g)=\frac{\partial g}{\partial\ln\mu}$ and $\beta(h)=\frac{\partial h}{\partial\ln\mu}$ for the renormalized couplings can be derived from $\widetilde{Z}_{h},\widetilde{Z}_{g}$ by solving the following matrix,
$\begin{pmatrix}\frac{1}{g}+\frac{\partial\ln\widetilde{Z}_{g}}{\partial g}&\frac{\partial\ln\widetilde{Z}_{g}}{\partial h}\\ \frac{\partial\ln\widetilde{Z}_{h}}{\partial g}&\frac{1}{h}+\frac{\partial\ln\widetilde{Z}_{h}}{\partial h}\\ \end{pmatrix}\begin{pmatrix}\frac{\partial g}{\partial\ln\mu}\\ \frac{\partial h}{\partial\ln\mu}\\ \end{pmatrix}=\begin{pmatrix}-\frac{\epsilon}{2}\\ -\frac{\epsilon}{2}\\ \end{pmatrix}~$. (9)
Then the field anomalous dimensions of $\varphi$ and $\sigma$ can be computed by definition as,
$\gamma_{\varphi}=\frac{1}{2}\left(\frac{\partial\ln Z_{\varphi}}{\partial g}\beta(g)+\frac{\partial\ln Z_{\varphi}}{\partial h}\beta(h)\right)~~~,~~~\gamma_{\sigma}=\frac{1}{2}\left(\frac{\partial\ln Z_{\sigma}}{\partial g}\beta(g)+\frac{\partial\ln Z_{\sigma}}{\partial h}\beta(h)\right)~.$ (10)
Furthermore, once $Z_{\varphi^{Q}}$ has been determined, the anomalous dimension of the $\varphi^{Q}$ operator can be computed by
$\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}=-\frac{\partial\ln Z_{\varphi^{Q}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}Z_{\varphi}^{-\frac{Q}{2}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}+Q\left(\frac{1}{2}\frac{\partial\ln Z_{\varphi}}{\partial\ln\mu}\right)~,$ (11)
where the first term computes $\gamma_{Q}:=\gamma_{\varphi^{Q}}$ and the second term computes $Q\gamma_{\varphi}$.
With the knowledge of beta functions and the anomalous dimensions, in order to study the large $N$ expansion of the anomalous dimensions around the Wilson-Fisher fixed point, one need to solve the fixed points from beta functions,
$\beta(g^{\ast})=0~~~,~~~\beta(h^{\ast})=0~.$ (12)
After substituting the solutions $g^{\ast},h^{\ast}$ into the anomalous dimensions and expanding the results in the series of $1/N$, one produces the large $N$ expansion of the cubic scalar theory at the IR fixed point.
The renormalization computation problems require the determination of the $Z$-factors. In general, a $Z$-factor of up to $L$-loop contributions takes the schematic form as
$Z_{{\tiny\texttt{X}}}=1+\sum_{n_{\ell}=1}^{L}\sum_{n_{\epsilon}=1}^{n_{\ell}}\frac{z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}}{\epsilon^{n_{\epsilon}}}~,$ (13)
in which the symbol X can represent the fields $\varphi,\sigma$, the couplings $h,g$ or the operator $\varphi^{Q}$ of arbitrary value $Q$. The $Z_{{\tiny\texttt{X}}}$-factor can be obtained by determining the coefficients $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ up to the appropriate loop orders. All such $Z$-factors can be conveniently computed using two-point propagator-type integrals, following the spirit of the OPE method to renormalization computations. We will review the OPE method in the next subsection.
2.2 The OPE method for the renormalization
In general quantum field theory, the product of two time-ordered operators admits an asymptotic expansion in the limit where their coordinates coincide, known as the Operator Product Expansion. The expansion basis is a set of composite operators of the theory, ordered according to their dimensions, while the expansion coefficients are functions of the coordinate difference, which are in general singular. When considering the OPE in momentum space by Fourier transformation with respect to the coordinate difference, the expansion is performed in the limit of large momentum $p\to\infty$, schematically expressed as,
$\widehat{O}_{\tiny\mbox{II}}^{\tiny\mbox{ren}}(k-p)\widehat{O}_{\tiny\mbox{I}}^{\tiny\mbox{ren}}(p)~\sim~C_{i}^{\tiny\mbox{ren}}(g_{i};p)\widehat{O}_{i}^{\tiny\mbox{ren}}(k)+\cdots$ (14)
where $k$ is a finite momentum, and the momenta of the two operators tend to the opposite directions of infinite limits [^1] Pictorially, we can consider the two operators with large momentum as two hard fields propagating in the background of soft fields. This idea has also been realized in the study of the boundary operators of BCFW recursion relations Huang:2022qnx . By shifting two external momenta $p_{1},p_{2}$ with a complex auxiliary momentum $zq$ as $p_{1}\to p_{1}+zq$, $p_{2}\to p_{2}-zq$, at the large $z$ limit we get two hard fields propagating in the background of soft fields.. The superscript $^{\tiny\mbox{ren}}$ in the operator $\widehat{O}$ represents a renormalized operator in dimensional regularization scheme.
Alternatively, we may consider the OPE of bare operators, schematically written as,
$\widehat{O}_{\tiny\mbox{II}}^{\tiny\mbox{bare}}(k-p)\widehat{O}_{\tiny\mbox{I}}^{\tiny\mbox{bare}}(p)~\sim~C_{i}^{\tiny\mbox{bare}}(g_{i0};p)\widehat{O}_{i}^{\tiny\mbox{bare}}(k)+\cdots$ (15)
where the Wilson coefficients are now functions of $p^{2}$ and the bare coupling constants. Since renormalized and bare operators are related by their corresponding $Z$-factor,
$\widehat{O}^{{\tiny\mbox{ren}}}=Z_{\widehat{O}}\widehat{O}^{{\tiny\mbox{bare}}}~,$ (16)
from eqn.(14) and eqn.(15) we get the relation,
$C_{i}^{\tiny\mbox{ren}}(g_{i};p)=\frac{Z_{\widehat{O}_{\tiny\mbox{I}}}Z_{\widehat{O}_{\tiny\mbox{II}}}}{Z_{\widehat{O}_{i}}}C_{i}^{\tiny\mbox{bare}}(g_{i0};p)~.$ (17)
Special attention should be paid to the OPE of fundamental fields, since the $Z$-factor of a fundamental field $\varphi$ or $\sigma$ does not correspond directly to $Z_{\varphi}$ or $Z_{\sigma}$. Instead, it is defined through $\varphi=Z_{\varphi}^{-\frac{1}{2}}\varphi_{0}$ or $\sigma=Z_{\sigma}^{-\frac{1}{2}}\sigma_{0}$. Furthermore, one usually computes the amputated correlation functions, and an extra factor $Z_{\varphi}$ or $Z_{\sigma}$ should be included for each fundamental field in the OPE of amputated correlation functions. With these two considerations, a fundamental field $\varphi$ or $\sigma$ contributes a $Z^{\frac{1}{2}}_{\varphi}$ or $Z^{\frac{1}{2}}_{\sigma}$ factor to the numerator of the Wilson coefficient relation (17) in the context of amputated OPE. However, the $Z$-factor in the denominator corresponds to an operator but not an external leg, so it still contributes to the $Z^{-\frac{1}{2}}_{\varphi}$ or $Z^{-\frac{1}{2}}_{\sigma}$ factor.
Due to the fact that the Wilson coefficients of renormalized operators $C_{i}^{\tiny\mbox{ren}}(g_{i};p)$ are UV finite, once the Wilson coefficients of bare operators $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed, the ratios of $Z$-factors can be determined using UV finiteness conditions generated from the cancellation of $\epsilon$ poles. It could be possible that some operators in the basis have the same dimension, leading to the operator mixing situation. In this case, the relation (16) is generalized to the vectors of mixed operators related by matrices of $Z$-factors. The $Z$-matrices can be determined by the UV finiteness conditions from all the OPE coefficients corresponding to the mixed operators. Importantly, the above discussion is based on general operators, which means that it is applicable to either simple operators, or operators with Lorentz indices, with fermion and gauge fields, etc.
The Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ of operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ can be computed by the large momentum expansion of bare correlation functions. For an operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ composed of $n$ fields, the corresponding Wilson coefficient is given by
$C_{i}^{\tiny\mbox{bare}}(g_{i0};p)=\Big{\langle}\widehat{O}_{\tiny\mbox{II}}^{\tiny\mbox{bare}}(k-p)\widehat{O}_{\tiny\mbox{I}}^{\tiny\mbox{bare}}(p)\varphi_{0}(k_{1})\varphi_{0}(k_{2})\cdots\varphi_{0}(k_{n})\Big{\rangle}\Big{|}_{k_{i}\to 0~,~i=1,\ldots,n}~,$ (18)
where $k=-\sum_{i=1}^{n}k_{i}$. Pictorially, this corresponding to splitting the Feynman graph of the original correlation function into hard and soft sub-graphs by cutting $n$ internal propagators and/or external legs, with the Wilson coefficient receiving contributions from all possible $n$-cut hard sub-graphs. If the operator contains fields with derivatives, for instance $\partial_{\mu}\varphi_{0}(k)$, we can compute the coefficient by acting $i\frac{\partial}{\partial k^{\mu}}$ to the bare correlation function. For hard sub-graphs, since all the $k_{i},i=1,\ldots,n$ have been set to zero and there remains only two operators with hard momenta, we effectively get two-point graphs. Hence the Wilson coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed by the two-point propagator-type integrals. This greatly simplifies the computation, especially when dealing with correlation functions with large number of external legs.
The above discussion outlines the core idea of computing renormalization functions via OPE method. To determine beta functions and anomalous dimensions, we first need the $Z$-factors, which are extracted from UV finiteness conditions generated from (17), i.e.,
$\frac{Z_{\widehat{O}_{\tiny\mbox{I}}}Z_{\widehat{O}_{\tiny\mbox{II}}}}{Z_{\widehat{O}_{i}}}C_{i}^{\tiny\mbox{bare}}(g_{i0};p)=\mbox{UV~finite}~.$ (19)
To establish these UV finiteness conditions, we first need to compute the coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ from the corresponding two-point integrals. Then using the general ansatz (13) for $Z$-factors, we can expand the complete expression as series expansion of $\epsilon$, and generate equations by setting coefficients of all $\epsilon$ poles to zero. The $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ coefficients of the $Z_{\tiny\texttt{X}}$-factor can be obtained by solving these equations, and consequently the $Z_{\tiny\texttt{X}}$-factor can be successfully determined to the desired loop order, provided the two-point integrals can be computed to that order.
Now let us apply the OPE method to the renormalization computation of the cubic scalar theory, working in coordinate space. As in the standard field theory computation, the $Z_{\varphi}$ and $Z_{\sigma}$-factors are computed from two-point correlation functions. This computation can be naturally framed within the OPE method. Consider the OPE,
$\varphi\bar{\varphi}~\sim~c_{\mathbf{1}}~\mathbf{1}+\cdots~~~,~~~\sigma\sigma~\sim~c^{\prime}_{\mathbf{1}}~\mathbf{1}+\cdots~,$ (20)
where $\mathbf{1}$ is the unit operator of zero dimension, and $c$ generically represents Wilson coefficient, which is in general different between the two relations. Since they are the OPE of two fundamental fields, the large momentum expansion of them produces,
$\langle\varphi(x_{1})\bar{\varphi}(x_{2})\rangle~\cong~Z_{\varphi}G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}))~~~,~~~\langle\sigma(x_{1})\sigma(x_{2})\rangle~\cong~Z_{\sigma}G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}))~,$ (21)
where we use notation $G(g_{i0};\varphi(x_{1}),\varphi(x_{2}),\varphi(\mbox{\textst{$x$}}_{3}),\ldots)$ to represent the bare Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$. It is computed by the two-point integrals, derived from the bare correlation function with $x$ denoting the corresponding external leg removed, as in eqn.(18). Since they are already two-point functions, no additional legs need removal, so $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}))$ are computed directly from Feynman graphs of two-point correlation functions. Then from UV finiteness conditions of (21), the $Z$-factors of fields $\varphi$ and $\sigma$ can be determined.
The $Z$-factors of coupling constants $g,h$ can be computed from the OPE,
$\varphi\bar{\varphi}~\sim~c_{\sigma}~\sigma+\cdots~~~,~~~\sigma\sigma~\sim~c^{\prime}_{\sigma}~\sigma+\cdots~.$ (22)
In the large momentum expansion, it produces $Z$-factor ratios $Z^{\frac{1}{2}}_{\varphi}Z^{\frac{1}{2}}_{\varphi}/Z^{-\frac{1}{2}}_{\sigma}$ and $Z^{\frac{1}{2}}_{\sigma}Z^{\frac{1}{2}}_{\sigma}/Z^{-\frac{1}{2}}_{\sigma}$ respectively. So we obtain,
$\langle\varphi(x_{1})\bar{\varphi}(x_{2})\sigma(x_{3})\rangle~\cong~Z_{\sigma}^{\frac{1}{2}}Z_{\varphi}G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~,$ (23)
$\langle\sigma(x_{1})\sigma(x_{2})\sigma(x_{3})\rangle~\cong~Z^{\frac{3}{2}}_{\sigma}G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~.$ (24)
Further using the relation (7) we get [^2] In the previous paper Huang:2024hsn , we have taken the convention that the $g_{0}$ or $h_{0}$ in the denominator is absorbed into the bare correlation function for three-point functions. This leads to slightly different expressions in that paper, where $G$ represents the bare correlation function divided by $g_{0}$ or $h_{0}$ respectively.
$\langle\varphi(x_{1})\bar{\varphi}(x_{2})\sigma(x_{3})\rangle~\cong~\frac{Z_{g}g}{g_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~,$ (25)
$\langle\sigma(x_{1})\sigma(x_{2})\sigma(x_{3})\rangle~\cong~\frac{Z_{h}h}{h_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~.$ (26)
The $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ are directly computed from the two-point integrals of corresponding three-point functions by removing one leg of $\sigma$ field, and the $Z_{g},Z_{h}$ can be determined from UV finiteness conditions of (25), (26). With results of $Z_{\varphi},Z_{\sigma}$ and $Z_{g},Z_{h}$, the beta functions and anomalous dimensions of fields can be readily computed by definitions (9), (10).
The $Z$-factor of the $\varphi^{Q}$ operator can be computed from the OPE,
$\varphi^{Q}\bar{\varphi}~\sim~c_{i}~\varphi^{Q-1}~.$ (27)
After large momentum expansion it produces the ratio of $Z$-factors as,
$\frac{Z_{\varphi^{Q}}Z^{\frac{1}{2}}_{\varphi}}{Z_{\varphi^{Q-1}}}=\frac{Z_{\varphi^{Q}}Z^{\frac{Q}{2}}_{\varphi}}{Z_{\varphi^{Q-1}}Z^{\frac{Q-1}{2}}_{\varphi}}=\frac{\widetilde{Z}_{\varphi^{Q}}}{\widetilde{Z}_{\varphi^{Q-1}}}~.$ (28)
Thus, the $Z_{\varphi^{Q}}$-factor can be computed recursively using the UV finiteness conditions,
$\frac{\widetilde{Z}_{\varphi^{Q}}}{\widetilde{Z}_{\varphi^{Q-1}}}G_{Q}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})=\mbox{UV~finite}~,$ (29)
where $G_{Q}$ is computed from Feynman graphs of the $\varphi^{Q}$-$Q\varphi$ correlation function by removing (Q1)(Q-1) external legs. They are two-point integrals with large momentum localized at the $\varphi^{Q}$ operator and one remaining external leg. Using the results of $Z_{\varphi^{Q}}$ and the previously computed beta functions, the anomalous dimension of the $\varphi^{Q}$ operator can be computed by definition (11).
2.3 From Feynman diagrams to two-point integrals
The UV finiteness conditions discussed in the previous subsection require computing two-point integrals derived from two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions. In this subsection, we review some techniques which will help to simplify the problem of generating Feynman diagrams and transfer them to two-point graphs for subsequent evaluation of two-point integrals.
The $\Phi$-field representation
The cubic scalar theory we consider involves two types of fields $\varphi,\sigma$, and two types of cubic interactions. The complex field $\varphi$ carries a scalar charge that flows along the line of $\varphi,\bar{\varphi}$ fields, which can be represented by an arrow indicating the direction of charge flow. When generating Feynman diagrams, we need to account for all possible field configurations in the internal lines, which becomes a rather difficult task when the number of internal lines is large. It has also been observed that for a given external field configuration, different internal field configurations only contribute to varying combinations of couplings, while the underlying integrals remain identical. Thus there is no need to generate all Feynman diagrams and redundantly compute the same integrals. To clarify the separation between the structure of couplings and the integrals, we will adopt the $\Phi$-field representation. In this framework, we combine the $\varphi_{i},i=1,\ldots,N$ fields and the $\sigma$ field into a single $(N+1)$ component scalar field as,
$\Phi_{I}=(\Phi_{0},\Phi_{1},\ldots,\Phi_{N}):=(\sigma,\varphi_{1},\ldots,\varphi_{N})~,$ (30)
and the cubic interactions can be universally encoded as
$\frac{1}{6}\lambda_{IJK}\Phi_{I}\Phi_{J}\Phi_{K}~~~\mbox{where}~~~\lambda_{000}=h~~\mbox{and}~~\lambda_{0ij}=g\delta_{ij}~~\forall i,j\neq 0~,$ (31)
with all other components of the symmetric coupling $\lambda_{IJK}$ vanishing. In the $\Phi$-field representation, we essentially generate the topologies of Feynman diagrams without specifying the explicit configuration of component fields. In this case, we have separated the information of the scalar integral part from the factor of couplings. The Feynman scalar integral can be derived from a given topology by assigning propagators according to Feynman rules, while the factors of couplings of all Feynman diagrams sharing the same topology are encoded in the combination of $\lambda_{IJK}$ couplings, which we define as the coupling factor.
The symmetry factor and the coupling factor
When writing the Feynman integral of correlation functions, we first generate all corresponding topologies with private Mathematica code. For each topology, we get the scalar integral from the combination of propagators, multiplied by the standard symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ from the total symmetry of the graph and external legs. Each topology effectively represents a collection of numerous Feynman diagrams that share the same external state configuration but differ in their internal state configurations. By summing over the indices of the $\lambda_{IJK}$, we get the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of the contributing Feynman graphs.
To clarify the concept of the coupling factor, let us take the two-point two-loop topologies $\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ depicted in Fig.(2) as an example.
$Ng^{4}$$Ng^{3}h$$Ng^{3}h$$h^{4}$$g^{3}h$$g^{3}h$$g^{4}$$g^{4}$
Figure 1: Examples of coupling factors in the $\Phi$-field representations and their corresponding component field representations. The internal $\Phi$ fields represent all possible internal state configurations in component fields.
For this topology, we can consider the external field configuration of two external $\sigma$ field represented by the dashed lines, or the field configuration of two external $\varphi,\bar{\varphi}$ fields represented by the arrowed lines, as depicted in the LHS of Fig.(1). Each vertex is assigned a $\lambda_{IJK}$, and after summing over the internal indices, we get the respective coupling factors as,
$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\sigma\sigma})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{0I_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}0}=Ng^{4}+2Ng^{3}h+h^{4}~,$ (32)
$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\varphi\varphi})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{iI_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}i}=2g^{3}h+2g^{4}~.$ (33)
Alternatively, we can draw all contributing Feynman diagrams with the explicit internal field configurations, as shown in the RHS of Fig.(1). It can be verified that summing over all four Feynman diagrams of each external field configuration indeed produces the coupling factors evaluated above. Since generating topologies is significantly more efficient than generating complete Feynman diagrams, and the evaluation of coupling factors involves only algebraic operations, the $\Phi$-field representation substantially enhances the computational capability.
The Feynman graphs of $\varphi^{Q}$-$Q\varphi$ correlation functions
In the UV finiteness conditions, the function $G$ is computed schematically as
$G=\sum\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{{\tiny\mbox{topo}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{{\tiny\mbox{topo}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{{\tiny\mbox{topo}}})~,$ (34)
where the summation runs over all possible topologies, and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{{\tiny\mbox{topo}}})$ is computed from the two-point integrals generated by these topologies. Remark that the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of a topology depends on the explicit external field configuration, but the symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}$ are irrelevant to the external fields, they are computed from the topologies. For two-point correlation functions, $G$ is directly evaluated using all 1PI two-point topologies. For three-point correlation functions, we start from all 1PI three-point topologies and compute the independent two-point graphs by removing one external leg from these topologies in all possible ways. For the $\varphi^{Q}$-$Q\varphi$ correlation function, we start from all 1PI $\varphi^{Q}\to Q\varphi$ topologies, and $G_{Q}$ can be computed from all independent two-point graphs generated by removing (Q1)(Q-1) external legs from these topologies in all possible ways.
By investigating the structure of 1PI $\varphi^{Q}\to Q\varphi$ topologies, it has been proven that only a part of these topologies contributes to the two-point integrals Huang:2024hsn . These contributing topologies are called irreducible graphs, characterized by the property that when the operator vertex is deleted, all remaining external legs stay connected in a single graph (except those that reduce to a single line). The other topologies are called reducible graphs, with the property that when deleting the operator vertex causes the external legs to distribute into two or more separate connected components. Hence we only need to generate the irreducible graphs of the 1PI $\varphi^{Q}\to Q\varphi$ topologies, and compute $G_{Q}$ from their corresponding two-point graphs after all possible removals of (Q1)(Q-1) legs. The summation of these graphs gives,
$G_{Q}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})=\mathbb{U}_{\footnotesize\mbox{v}}\left(\frac{1}{x_{Q}^{2\lambda}}\right)+\sum_{n=2}^{Q}C_{Q-1}^{n-1}\mathtt{F}_{n}(\mbox{\textst{$x$}}_{\sigma_{1}},\cdots,\mbox{\textst{$x$}}_{\sigma_{n-1}},x_{Q})~.$ (35)
The first term comes from the tree-level contribution and evaluates to 1, while
$\mathtt{F}_{n}(\mbox{\textst{$x$}}_{\sigma_{1}},\cdots,\mbox{\textst{$x$}}_{\sigma_{n-1}},x_{Q})=(n-1)!\sum_{n_{\ell}=n-1}^{L}G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})~.$ (36)
The $G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}$ receives contributions from all $n_{\ell}$-loop irreducible topologies. As in the standard procedure, it can be evaluated using the symmetry factor, coupling factor and UV divergence as,
$G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})=\sum_{{\tiny\mbox{irreducible}}\atop{\tiny\mbox{topologies}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{n_{\ell}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\cdots\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{n_{\ell}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\cdots\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{n_{\ell}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\cdots\mbox{\textst{{\scriptsize$\varphi$}}}})~,$ (37)
where the summation is over all non-isomorphic graphs generated by removing (Q1)(Q-1) external legs of the irreducible 1PI topologies of the $\varphi^{Q}$-$Q\varphi$ correlation function.
Let us summarize this section. After generating the independent topologies of $\varphi^{Q}$-$Q\varphi$ correlation functions, one computes the symmetry factor and coupling factor of each topology, and the UV divergences of two-point integrals in the dimensional regularization scheme to an appropriate $\epsilon$ order. Then the $G$ and $G_{Q}$ can be readily computed by summing over all two-point graphs. The $Z$-factors could be determined from UV finiteness conditions, and eventually the anomalous dimensions can be computed from these $Z$-factors.
3 Two-loop demonstration
For illustrative purposes, in this section we will present detailed computations of the two-loop anomalous dimensions of the $\varphi^{Q}$ operator. This requires computing the field renormalization $Z_{\varphi}$ and $Z_{\sigma}$-factors, the coupling renormalization $Z_{h}$ and $Z_{g}$-factors and the $\varphi^{Q}$ operator renormalization $Z_{\varphi^{Q}}$-factor. These factors are determined from the corresponding two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions respectively. All results related to Feynman diagrams are computed by private Mathematica code, while the UV divergences of two-point integrals are computed by HyperlogProcedures Maple package HP . The strategy of computation is formalized as a sequential algorithm, following the computations of the symmetry factors of topologies, the coupling factors of Feynman diagrams, the UV divergence of two-point integrals and finally the determination of $Z$-factors from UV finiteness conditions.
3.1 Two-point correlation functions and the field renormalization
Let us begin with the field renormalization computation. The $Z$-factors $Z_{\varphi}$ and $Z_{\sigma}$ can be computed from the two-point correlation functions $\langle\varphi\bar{\varphi}\rangle$ and $\langle\sigma\sigma\rangle$ respectively as in eqn.(21). The one-loop and two-loop topologies in $\Phi$-field representation are sketched in Fig.(2).
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$
Figure 2: The 1-loop and 2-loop 1PI topologies of two-point functions in $\Phi$-field representation.
Each topology represents a collection of diagrams with distinct allowable internal field configurations of $\sigma,\varphi_{i}$, and these collections are characterized by coupling factors. The symmetry factors and the coupling factors of these topologies are listed in the table below,
topology $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma_{\Phi\Phi})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\sigma\sigma})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\varphi\bar{\varphi}})$
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi}$ 1/2 $h_{0}^{2}+Ng_{0}^{2}$ $2g_{0}^{2}$
$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi}$ 1/2 $N(2g_{0}^{4}+g_{0}^{2}h_{0}^{2})+h_{0}^{4}$ $Ng_{0}^{4}+2g_{0}^{4}+g_{0}^{2}h_{0}^{2}$
$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ 1/2 $N(g_{0}^{4}+2g_{0}^{3}h_{0})+h_{0}^{4}$ $2g_{0}^{4}+2g_{0}^{3}h_{0}$
There is no need to cut any external legs of these topologies, and their corresponding UV divergences can be directly computed by TwoPoint command in HyperlogProcedures. Feynman integrals in the $\overline{\mbox{MS}}$ dimensional regularization scheme can be evaluated as power series of parameter $\epsilon$ with the results(an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral has not been explicitly shown),
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi})=-\frac{1}{3\epsilon}-\frac{4}{9}+\left(\frac{\pi^{2}}{72}-\frac{13}{27}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (38)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi})=\frac{1}{18\epsilon^{2}}+\frac{43}{216\epsilon}+\left(\frac{1171}{2592}-\frac{\pi^{2}}{216}\right)+\left(-\frac{\zeta_{3}}{108}-\frac{43\pi^{2}}{2592}+\frac{25195}{31104}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (39)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi})=-\frac{1}{3\epsilon^{2}}-\frac{1}{\epsilon}+\left(\frac{\pi^{2}}{36}-\frac{15}{8}\right)+\left(-\frac{7\zeta_{3}}{36}+\frac{\pi^{2}}{12}-\frac{769}{288}\right)\epsilon+\mathcal{O}(\epsilon^{2})~.$ (40)
Then the bare correlation functions are readily obtained as,
$G(g_{0},h_{0};\Phi,\Phi)=1-\sum_{L={\tiny\mbox{L1}},\atop{\tiny\mbox{L2-I}},{\tiny\mbox{L2-II}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{L}_{\Phi\Phi})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{L}_{\Phi\Phi})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{L}_{\Phi\Phi})~,$ (41)
where the field $\Phi$ takes the corresponding component $\sigma$ or $\varphi$ in $\mathbb{C}_{\footnotesize\mbox{f}}$ factor. Attention should be paid to the $\pm$ signs of individual terms in the summation. The bare correlation function is computed by summing contributions from all tree-level and loop Feynman diagrams. Note that the TwoPoint command only computes the scalar integrals, omitting all $i$ factors from, for instance, the vertices or propagators. When summing these contributions, we must recover the factor $i$ for each loop order. Each vertex carries a factor of i-i. Each propagator contains a factor of $i$ in its numerator, but after Wick rotation to Euclidean space, the denominator acquires an additional factor of 1-1, resulting in an overall factor of i-i for each propagator. Each loop integration gains a factor of $i$ from the Wick rotation $\ell_{0}\to i\ell_{E0}$. Supposing a loop graph has $n_{v}$ vertices, $n_{e}$ propagators, $n_{\ell}$ loops and $n_{p}$ external legs, then the total factor would be $(-i)^{n_{v}}(-i)^{n_{e}}i^{n_{\ell}}$. From graph theory, we have the relations $n_{v}+n_{\ell}-n_{e}=1$ and $3n_{v}=2n_{e}+n_{p}$, and after calculation we get the final factor $(-1)^{n_{p}}i$ for each loop order. For two-point functions, $n_{p}=2$, so each loop order has a factor $i$, whereas the tree-level contrition has a factor i-i. If we factor out an overall i-i, the two-point correlation function can be evaluated as the tree-level contribution minus the sum of loop contributions.
The renormalized two-point correlation functions eqn.(21) should be UV finite, so we get the UV finiteness conditions
$Z_{\varphi}G(g_{0},h_{0};\varphi(x_{0}),\bar{\varphi}(x_{1}))=\mbox{UV~finite}~~~,~~~Z_{\sigma}G(g_{0},h_{0};\sigma(x_{0}),\sigma(x_{1}))=\mbox{UV~finite}~,$ (42)
which means they should be free of $\epsilon$ poles. In order to get the equations of UV finiteness conditions, we rewrite the bare couplings $g_{0},h_{0}$ in the bare correlation functions to $g,h$ by relations eqn.(7), then expand the full expression at the limit $g,h\to 0$ followed by the limit $\epsilon\to 0$. Note that in the rewriting of couplings with eqn.(7), each $g$ or $h$ is accompanied by a factor $\mu^{\frac{\epsilon}{2}}$, so from the coupling factors we know that at $L$-loop order there will be a $\mu^{L\epsilon}$ factor. Combined with the $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor, at $L$-loop order there will be an overall $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor. Expansion of the factor around $\epsilon\to 0$ produces powers of $\ln\frac{\mu^{2}}{p^{2}}$ terms. Since each $\ln\frac{\mu^{2}}{p^{2}}$ order is independent, the UV finiteness conditions also require the coefficient of each $\ln\frac{\mu^{2}}{p^{2}}$ order is free of $\epsilon$ poles. Hence the higher Log order terms will produce additional equations. Inspecting the UV divergence structures of $L$-loop integrals, the leading divergence is $1/\epsilon^{L}$}, thus at the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ order there will in general be $L$ equations for the cancellation of $L$ poles. At the $\ln^{k}\frac{\mu^{2}}{p^{2}}$ order, the leading divergence is $1/\epsilon^{L-k}$, so there will be (Lk)(L-k) equations. We can either use the equations from the leading Log terms [^3] The leading Log term denotes terms with $\ln^{0}(\mu^{2}/p^{2})$ dependence, and the sub-Log term denotes terms with $\ln(\mu^{2}/p^{2})$ dependence, etc., or include those equations from higher Log terms. If we only use equations of the leading Log terms, then all equations of $Z_{\varphi},Z_{\sigma},Z_{g},Z_{h}$ should be solved simultaneously to completely determine the $Z$-factors. However, if we also use the equations of sub-Log terms, then each $Z$-factor can be determined by solving its own equations of UV finiteness conditions. This will be demonstrated in the following computations.
Now let us take a look at the UV finiteness conditions. For $Z_{\varphi}$-factor, at one-loop, after the above prescribed expansions, we get
$\left(\frac{1}{\epsilon}\left(z_{\phi}^{11}+\frac{g^{2}}{3}\right)+\left(1+\frac{7g^{2}}{9}\right)+\mathcal{O}(\epsilon)\right)+\left(\frac{g^{2}}{3}+\frac{7g^{2}}{9}\epsilon+\mathcal{O}(\epsilon^{2})\right)\ln\frac{\mu^{2}}{p^{2}}+\cdots$ (43)
The coefficients of sub-Log terms are already UV finite, and they will not provide any additional equations. In order for the leading Log term to be finite, the $1/\epsilon$ pole should cancel, and we get
$z_{\varphi}^{11}=-\frac{g^{2}}{3}~.$ (44)
At two-loop order, we need to solve $z^{21}_{\varphi},z^{22}_{\varphi}$. Again after expansions, the vanishing of $1/\epsilon$ and $1/\epsilon^{2}$ poles at leading Log term produces two equations,
$0=z_{\varphi}^{21}+z_{\varphi}^{11}+\frac{7}{3}\left(\frac{2}{3}g^{2}z_{g}^{11}-\frac{1}{3}g^{2}z_{\varphi}^{11}-\frac{1}{3}g^{2}z_{\sigma}^{11}\right)+\left(\frac{g^{2}}{3}-\frac{67g^{4}N}{432}+\frac{293g^{4}}{216}+\frac{5g^{3}h}{3}-\frac{67g^{2}h^{2}}{432}\right)~,$
$0=z_{\varphi}^{22}+\left(\frac{2}{3}g^{2}z_{g}^{11}-\frac{1}{3}g^{2}z_{\varphi}^{11}-\frac{1}{3}g^{2}z_{\sigma}^{11}\right)+\left(\frac{5g^{4}}{18}+\frac{g^{3}h}{3}-\frac{g^{4}N}{36}-\frac{g^{2}h^{2}}{36}\right)~,$
and the vanishing of $1/\epsilon$ pole at sub-Log term produces one equation,
$0=\left(\frac{2}{3}g^{2}z_{g}^{11}-\frac{1}{3}g^{2}z_{\varphi}^{11}-\frac{1}{3}g^{2}z_{\sigma}^{11}\right)+\left(\frac{5g^{4}}{9}+\frac{2g^{3}h}{3}-\frac{g^{4}N}{18}-\frac{g^{2}h^{2}}{18}\right)~.$
In the traditional approach, the renormalization computation should be performed order by order. This is similar to the situation that solving the $Z$-factors using only equations of leading Log terms. For instance, by using two equations of leading Log term shown above, the determination of $z_{\varphi}^{21},z_{\varphi}^{22}$ will also rely on the results of all one-loop $Z$-factors, including $Z_{g},Z_{h},Z_{\varphi},Z_{\sigma}$, which must be determined a priori. Alternatively, if considering the UV finiteness conditions of all Log terms of a correlation function, the generated equations are sufficient to completely solve the corresponding $Z$-factor. This means that the computation does not rely on other $Z$-factors of lower-loop orders. For instance, the equation from sub-Log term shown above can be solved to obtain a combination of one-loop $Z$-factors. This one-loop $Z$-factor combination is exactly the one appearing in the equations of leading Log term, and after substituting it back to those two equations, we can completely determine the two-loop $Z_{\varphi}$-factor. Furthermore, as a by-product, the equations of sub-Log terms provide constraints to other $Z$-factors at lower-loop orders, which can be served as consistent checks for lower-loop $Z$-factors computed by other correlation functions.
Consequently, we determine the two-loop $Z_{\varphi}$-factors as,
$z_{\varphi}^{21}=\frac{11g^{4}N}{432}-\frac{13g^{4}}{216}-\frac{g^{3}h}{9}+\frac{11g^{2}h^{2}}{432}~~~,~~~z_{\varphi}^{22}=-\frac{g^{4}N}{36}+\frac{5g^{4}}{18}+\frac{g^{3}h}{3}-\frac{g^{2}h^{2}}{36}~.$ (45)
The determination of $Z_{\sigma}$-factor follows exactly the same procedures, and the only difference is the coupling factors. After the same computations, eventually we determine the two-loop $Z_{\sigma}$-factor as
$z_{\sigma}^{11}=-\frac{g^{2}N}{6}-\frac{h^{2}}{6}~,$ (46)
$z_{\sigma}^{21}=\frac{11g^{2}h^{2}N}{432}-\frac{g^{4}N}{216}-\frac{g^{3}hN}{9}-\frac{13h^{4}}{432}~~~,~~~z_{\sigma}^{22}=\frac{g^{4}N}{9}+\frac{g^{3}hN}{3}-\frac{g^{2}h^{2}N}{36}+\frac{5h^{4}}{36}~.$ (47)
The higher-loop $Z_{\varphi},Z_{\sigma}$-factors can be determined following the same procedures. However there would be much more Feynman graphs, and the evaluation of two-point integrals would become considerably more difficult.
3.2 Three-point correlation functions and the beta functions
The determination of $Z_{h},Z_{g}$-factors is quite similar as that of $Z_{\varphi},Z_{\sigma}$-factors. We start from OPE and arrive at the UV finiteness conditions as eqn.(25), (26). Once the large momentum expansion of the bare correlation functions are computed, we can determine $Z_{g}$ and $Z_{h}$ from the UV finiteness conditions. Remind that the UV divergence is computed by all possible three-point topologies with one leg removed. Following the computation strategy, the one-loop and two-loop topologies in $\Phi$-field representation are sketched in Fig.(3), as well as their independent cut graphs.
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$
Figure 3: The 1-loop and 2-loop 1PI topologies of three-point functions in $\Phi$-field representation. A gray arrow points to the cut graphs of each topology. They are generated by the non-isomorphic computation of all possible removals of one external leg.
The $Z_{g}$-factor is determined by $\langle\varphi\bar{\varphi}\sigma\rangle$ with a $\sigma$ field removed, and the $Z_{h}$-factor is determined by $\langle\sigma\sigma\sigma\rangle$ with a $\sigma$ field removed. The symmetry factors and coupling factors of cut graphs are listed in Table.(1).
topology $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\sigma\sigma\mbox{\textst{{\scriptsize$\Phi$}}}})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\varphi\bar{\varphi}\mbox{\textst{{\scriptsize$\Phi$}}}})$
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1 $g^{3}N+h^{3}$ $g^{3}+g^{2}h$
$\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1/2 $2g^{5}N+g^{2}h^{3}N+h^{5}$ $g^{5}N+2g^{4}h+g^{3}h^{2}$
$\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1 $2g^{5}N+g^{2}h^{3}N+h^{5}$ $2g^{5}+g^{4}hN+g^{2}h^{3}$
$\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1 $g^{5}N+g^{4}hN+g^{3}h^{2}N+h^{5}$ $g^{5}N+g^{5}+g^{4}h+g^{2}h^{3}$
$\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 2 $g^{5}N+g^{4}hN+g^{3}h^{2}N+h^{5}$ $g^{5}+2g^{4}h+g^{3}h^{2}$
$\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1/2 $3g^{4}hN+h^{5}$ $2g^{5}+2g^{3}h^{2}$
Table 1: The symmetry factors and coupling factors of cut graphs for 2-loop three-point topologies.
The UV divergence is computed by TwoPoint command for two-point integral of cut graphs, with the results [^4] Again an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral has not been explicitly shown,
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{\epsilon}+3+\left(\frac{23}{4}-\frac{\pi^{2}}{16}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (48)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{67}{72\epsilon}-\left(\frac{2671}{864}-\frac{5\pi^{2}}{288}\right)-\left(\frac{80911}{10368}-\frac{335\pi^{2}}{3456}-\frac{5\zeta_{3}}{144}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (49)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{67}{72\epsilon}-\left(\frac{2743}{864}-\frac{5\pi^{2}}{288}\right)-\left(\frac{86167}{10368}-\frac{335\pi^{2}}{3456}-\frac{5\zeta_{3}}{144}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (50)
and
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{21}{8\epsilon}+\left(\frac{275}{32}-\frac{5\pi^{2}}{96}\right)+\left(\frac{2785}{128}-\frac{35\pi^{2}}{128}-\frac{5\zeta_{3}}{48}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (51)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{21}{8\epsilon}+\left(\frac{259}{32}-\frac{5\pi^{2}}{96}\right)+\left(\frac{7139}{384}-\frac{35\pi^{2}}{128}+\frac{19\zeta_{3}}{48}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (52)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{2\epsilon}+\left(\frac{83}{24}-\zeta_{3}\right)+\left(\frac{3983}{288}-\frac{5\pi^{2}}{96}-\frac{16\zeta_{3}}{3}-\frac{\pi^{4}}{120}\right)\epsilon+\mathcal{O}(\epsilon^{2})~.$ (53)
Special attention should be paid to the tree-level contribution, which is evaluated by the two-point integral of a three-point vertex with one leg removed. Evaluated in the coordinate space, the result is not 1, but
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L0}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=1+\epsilon+\left(\frac{3}{4}-\frac{\pi^{2}}{48}\right)\epsilon^{2}+\left(\frac{1}{2}-\frac{\pi^{2}}{48}-\frac{\zeta_{3}}{24}\right)\epsilon^{3}+\left(\frac{5}{16}-\frac{\pi^{2}}{64}-\frac{\zeta_{3}}{24}+\frac{\pi^{4}}{23040}\right)\epsilon^{4}+\mathcal{O}(\epsilon^{5})~.$ (54)
Appropriate $\epsilon$ order terms should be included for the $L$-loop renormalization computation. Then the bare correlation functions for two-loop computation are readily obtained as [^5] For three-point function, $n_{p}=3$. As discussed in the two-point case, each loop order has a factor i-i, while at the tree-level the factor is also i-i. If factoring out an overall factor i-i, the three-point correlation function can be evaluated by tree contribution plus all loop contributions.,
$G(g_{0},h_{0};\Phi,\Phi,\mbox{\textst{$\Phi$}})=g_{\Phi}(1+\epsilon)+\sum_{L={\tiny\mbox{L1}},{\tiny\mbox{L2-I}}_{1},{\tiny\mbox{L2-I}}_{2},\atop{\tiny\mbox{L2-II}}_{1},{\tiny\mbox{L2-II}}_{2},{\tiny\mbox{L2-III}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{L}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{L}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{L}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})~,$ (55)
where the field $\Phi$ takes the corresponding component $\sigma$ or $\varphi$ in $\mathbb{C}_{\footnotesize\mbox{f}}$ factors, and $g_{\Phi}=g$ for $\langle\varphi\bar{\varphi}\sigma\rangle$ while $g_{\Phi}=h$ for $\langle\sigma\sigma\sigma\rangle$.
The UV finiteness conditions of eqn.(25) and eqn.(26) can be realized by,
$\frac{Z_{g}g}{g_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\varphi,\bar{\varphi},\mbox{\textst{$\sigma$}})=\mbox{UV finite}~~,~~\frac{Z_{h}h}{h_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\sigma,\sigma,\mbox{\textst{$\sigma$}})=\mbox{UV~finite}~.$ (56)
After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ by relations eqn.(7), then expanding the full expression at the limit $g,h\to 0$ followed by the limit $\epsilon\to 0$, we get equations for $Z$-factors by the vanishing of $\epsilon$ poles in each $\ln\frac{\mu^{2}}{p^{2}}$ term [^6] The bare coupling in the denominator will introduce a $\mu^{\frac{\epsilon}{2}}$, which cancels the $\mu^{\frac{\epsilon}{2}}$ factor in numerator. In the bare correlation function, at $L$-loop order there will be a $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor from integration, and a $\mu^{\frac{\epsilon}{2}(2L+1)}$ factor from the couplings. If we take out an overall factor $\mu^{\frac{\epsilon}{2}}$ which will not affect the $\epsilon$ expansion, then each $L$-loop is assigned with a $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor contributing to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion at $\epsilon\to 0$ limit..
For $Z_{g}$-factor, at one-loop there is only one equation from the vanishing of $1/\epsilon$ pole of the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term, which yields,
$g^{3}+g^{2}h+gz_{g}^{11}=0~~~\to~~~z_{g}^{11}=-g^{2}-gh~.$ (57)
At two-loop order, there will be two equations from the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term,
$0=z_{g}^{22}+z_{g}^{21}+(9g^{2}+6gh+1)z_{g}^{11}+3g\Big{(}hz_{h}^{11}-(2g+h)z_{\phi}^{11}-(g+2h)z_{\sigma}^{11}\Big{)}$
$~~~~~~~~~~~~~~+\left(\frac{311g^{4}N}{144}+\frac{469g^{4}}{72}-\frac{67g^{3}hN}{72}+\frac{439g^{3}h}{36}+\frac{761g^{2}h^{2}}{144}+g^{2}+\frac{61gh^{3}}{36}+gh\right)~,$ (58)
$0=z_{g}^{22}+(2gh+3g^{2})z_{g}^{11}+g\Big{(}hz_{h}^{11}-(2g+h)z_{\phi}^{11}-(g+2h)z_{\sigma}^{11}\Big{)}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(\frac{5g^{4}N}{12}+\frac{7g^{4}}{6}-\frac{g^{3}hN}{6}+\frac{7g^{3}h}{3}+\frac{11g^{2}h^{2}}{12}+\frac{gh^{3}}{3}\right)~,$ (59)
and one equation from the sub-Log term,
$0=(3g+2h)z_{g}^{11}+\Big{(}hz_{h}^{11}-(2g+h)z_{\phi}^{11}-(g+2h)z_{\sigma}^{11}\Big{)}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(\frac{5g^{3}N}{6}+\frac{7g^{3}}{3}-\frac{g^{2}hN}{3}+\frac{14g^{2}h}{3}+\frac{11gh^{2}}{6}+\frac{2h^{3}}{3}\right)~.$ (60)
Again we confirm that the one-loop $Z$-factors can be solved as a whole by the equation of sub-Log term. Consequently, from the equations of leading Log terms we get
$z_{g}^{21}=\frac{7g^{3}hN}{72}-\frac{11g^{4}N}{144}-\frac{49g^{4}}{72}-\frac{19g^{3}h}{36}-\frac{101g^{2}h^{2}}{144}-\frac{gh^{3}}{36}~~,~~z_{g}^{22}=\frac{5g^{4}N}{12}+\frac{7g^{4}}{6}-\frac{g^{3}hN}{6}+\frac{7g^{3}h}{3}+\frac{11g^{2}h^{2}}{12}+\frac{gh^{3}}{3}~.$
Similarly, solving the equations of UV finiteness conditions for $Z_{h}$-factors, we determine the two-loop $Z_{h}$-factor. We get $z_{h}^{11}=-(g^{3}N+h^{3})/h$, and
$z_{h}^{21}=\frac{7g^{2}h^{2}N}{48}-\frac{g^{5}N}{12h}-\frac{9g^{4}N}{8}-\frac{3g^{3}hN}{8}-\frac{23h^{4}}{48}~~,~~z_{h}^{22}=\frac{g^{5}N}{h}+\frac{3g^{4}N}{2}+\frac{3g^{3}hN}{2}-\frac{g^{2}h^{2}N}{4}+\frac{5h^{4}}{4}~.$
The higher-loop $Z_{h},Z_{g}$-factors can be similarly determined by including more graphs and more orders of $\epsilon$ series of two-point integrals.
With knowledge of $Z_{h},Z_{g},Z_{\phi}$ and $Z_{\sigma}$, we can compute beta functions by solving equations (9). After expanding the solutions at the limits $g,h\to 0$, we get two-loop results as,
$\beta(g)$ $=$ $-\frac{\epsilon}{2}g+\left(\frac{g^{3}N}{12}-\frac{2g^{3}}{3}-g^{2}h+\frac{gh^{2}}{12}\right)$ (61)
$+\left(\frac{11g^{4}hN}{36}+\frac{13gh^{4}}{432}-\frac{43g^{5}N}{216}-\frac{67g^{5}}{54}-\frac{5g^{4}h}{6}-\frac{11g^{3}h^{2}N}{432}-\frac{157g^{3}h^{2}}{108}-\frac{g^{2}h^{3}}{18}\right)~,$~~
$\beta(h)$ $=$ $-\frac{\epsilon}{2}h+\left(\frac{g^{2}hN}{4}-g^{3}N-\frac{3h^{3}}{4}\right)+\left(\frac{31g^{2}h^{3}N}{144}-\frac{g^{5}N}{6}-\frac{161g^{4}hN}{72}-\frac{5g^{3}h^{2}N}{12}-\frac{125h^{5}}{144}\right)~.$~~ (62)
The 2-loop anomalous dimension of field $\varphi$ or $\sigma$ is then given by definition (10) as,
$\gamma_{\varphi}=\frac{g^{2}}{6}+\frac{13g^{4}}{216}+\frac{g^{3}h}{9}-\frac{11g^{4}N}{432}-\frac{11g^{2}h^{2}}{432}~~~,~~~\gamma_{\sigma}=\frac{h^{2}}{12}+\frac{g^{2}N}{12}+\frac{g^{4}N}{216}+\frac{g^{3}hN}{9}-\frac{11g^{2}h^{2}N}{432}+\frac{13h^{4}}{432}~.$ (63)
3.3 Anomalous dimensions of the $\varphi^{2}$ and $\varphi^{3}$ operators
The $\widetilde{Z}_{\varphi^{Q}}$-factors for the $\varphi^{Q}$ operator are determined by the UV finiteness conditions (29), and the corresponding two-point integrals are evaluated by (35). When $Q=2$, all contributing topologies are irreducible topologies. The tree-level, 1-loop and 2-loop topologies of $\varphi^{2}$-$2\varphi$ correlation function have been sketched in Fig.(4), together with their corresponding cut graphs.
$\Gamma^{\tiny\mbox{L0}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L0}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-III}_{1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}i}$$\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{2}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-III}_{2}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$
Figure 4: The tree-level, 1-loop and 2-loop 1PI topologies of $\varphi^{2}$-$2\varphi$ correlation function, where the in subscript denotes a $\phi^{Q}$ operator. A gray arrow points to the cut graphs of each topology. They are generated by the non-isomorphic computation of all possible removals of one external leg. The UV divergences are computed from these cut graphs.
Note that due to the symmetry of graphs, a topology could generate more than one non-isomorphic cut graphs, similar as we have mentioned in the three-point function computation. For the cut graphs related to the same irreducible topology, the symmetry factors as well as the coupling factors are the same as those of the topology. Now Let us compute the $Z$-factor of $\varphi^{2}$, from eqn.(35) we get the explicit expression for $\varphi^{2}$ as
$G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})=1+\sum_{j=1}^{L}G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})~.$ (64)
There is only one 1-loop irreducible topology $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi{\varphi}}$, which generates one cut graph $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}}$. With results,
$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g^{2}_{0}~~~,~~~\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{\epsilon}+2+\epsilon\left(\frac{5}{2}-\frac{\pi^{2}}{24}\right)+\mathcal{O}(\epsilon^{2})~,$ (65)
we get the required UV contribution of all 1-loop irreducible topologies for $\varphi^{2}\to 2\varphi$ as
$G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{1-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})=\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})~.$ (66)
Note that in (65), an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ with $L=1$ is not explicitly shown in the result of UV divergence, while there will also be a factor $\mu^{\epsilon L}$ from the couplings $\mathbb{C}_{\footnotesize\mbox{f}}$ at $L$-loop. Hence for (66), if recovering the $(1/p^{2})^{\frac{\epsilon}{2}}$ factor and turning $g_{0},h_{0}\to g,h$, there will be an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{\frac{\epsilon}{2}}$ factor, which contributes to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion. Generally, for $G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}$ there will be an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{j\frac{\epsilon}{2}}$ factor. For presentation purpose, the factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ will always be treated as implicit in the expressions of UV divergences.
The 1-loop $\widetilde{Z}_{\varphi^{2}}$ can be determined from the UV finiteness of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, where we have used the fact that the 1PI contribution of $\widetilde{Z}_{\varphi^{1}}=1$. After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ by relations eqn.(7), then expanding the full expression at the limits $g,h\to 0$ followed by the limit $\epsilon\to 0$, we get the condition for the vanishing of $1/\epsilon$ pole in the leading Log term as,
$0=z_{\varphi^{2}}^{11}+g^{2}~~~\to~~~z_{\varphi^{2}}^{11}=-g^{2}~,$ (67)
and the sub-Log term provides no more conditions. Similarly, in order to determine the 2-loop $\widetilde{Z}_{\varphi^{2}}$, the contribution of $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ should be included. There are 5 independent irreducible topologies, which generates 7 cut graphs. The symmetry factors and coupling factors can be simply computed as,
$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~,$ (68)
and
$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}N+g_{0}^{2}h_{0}^{2}~~,~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}+g_{0}^{3}h_{0}~~,~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=2g_{0}^{4}~~,~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}~.$ (69)
The UV divergences are again computed by TwoPoint command in HyperlogProcedures, which are
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{55}{72\epsilon}+\left(\frac{\pi^{2}}{72}-\frac{1759}{864}\right)+\epsilon\left(\frac{\zeta_{3}}{36}+\frac{55\pi^{2}}{864}-\frac{42487}{10368}\right)+\mathcal{O}(\epsilon^{2})~,$ (70)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}_{1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{155}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(\frac{5\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{2915}{384}\right)+\mathcal{O}(\epsilon^{2})~,$ (71)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}_{2}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{171}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(-\frac{\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{1313}{128}\right)+\mathcal{O}(\epsilon^{2})~,$ (72)
and
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-III}_{1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{55}{72\epsilon}+\left(\frac{\pi^{2}}{72}-\frac{1687}{864}\right)+\epsilon\left(\frac{\zeta_{3}}{36}+\frac{55\pi^{2}}{864}-\frac{38095}{10368}\right)+\mathcal{O}(\epsilon^{2})~,$ (73)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-III}_{2}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{55}{72\epsilon}+\left(\frac{\pi^{2}}{72}-\frac{1759}{864}\right)+\epsilon\left(\frac{\zeta_{3}}{36}+\frac{55\pi^{2}}{864}-\frac{42487}{10368}\right)+\mathcal{O}(\epsilon^{2})~,$ (74)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{155}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(\frac{5\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{2915}{384}\right)+\mathcal{O}(\epsilon^{2})~,$ (75)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon}+\left(\frac{71}{24}-\zeta_{3}\right)+\epsilon\left(-\frac{\pi^{4}}{120}-\frac{13\zeta_{3}}{3}-\frac{\pi^{2}}{24}+\frac{2807}{288}\right)+\mathcal{O}(\epsilon^{2})~.$ (76)
Then the required UV contribution of all 2-loop irreducible topologies for $\varphi^{2}\to 2\varphi$ is
$G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})=\sum_{L={\tiny\mbox{L2-I}},{\tiny\mbox{L2-II}}_{1},{\tiny\mbox{L2-II}}_{2},\atop{\tiny\mbox{L2-III}}_{1},{\tiny\mbox{L2-III}}_{2},{\tiny\mbox{L2-IX}},{\tiny\mbox{L2-X}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})~.$ (77)
After the series expansion of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, we get the UV finiteness conditions for the vanishing of $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term as,
$0=z_{\varphi^{2}}^{21}+\left(2g^{2}+1\right)z_{\varphi^{2}}^{11}+\left(g^{2}-\frac{7g^{4}N}{144}+\frac{193g^{4}}{72}+\frac{g^{3}h}{4}-\frac{7g^{2}h^{2}}{144}\right)~,$ (78)
$0=z_{\varphi^{2}}^{22}+g^{2}z_{\varphi^{2}}^{11}+\left(\frac{g^{4}N}{12}-\frac{g^{4}}{6}-g^{3}h+\frac{g^{2}h^{2}}{12}\right)~.$ (79)
The $z_{\varphi^{2}}^{11}$ has already been determined in the 1-loop computation, thus by solving above two equations we get,
$z_{\varphi^{2}}^{21}=\frac{7g^{4}N}{144}-\frac{49g^{4}}{72}-\frac{g^{3}h}{4}+\frac{7g^{2}h^{2}}{144}~~~,~~~z_{\varphi^{2}}^{22}=-\frac{g^{4}N}{12}+\frac{7g^{4}}{6}+g^{3}h-\frac{g^{2}h^{2}}{12}~.$ (80)
Note that the vanishing of $1/\epsilon$ pole in the sub-Log term provides an additional equation $0=g^{2}z_{\varphi^{2}}^{11}+g^{4}$. Hence even if we did not compute $z_{\varphi^{2}}^{11}$ at 1-loop, equations of leading Log and sub-Log terms are sufficient to solve $z_{\varphi^{2}}^{11},z_{\varphi^{2}}^{21},z_{\varphi^{2}}^{22}$, and completely determine the 2-loop $\widetilde{Z}_{\varphi^{2}}$-factor.
The determination of $\widetilde{Z}_{\varphi^{3}}$-factor follows exactly the same procedure, where now we should use the UV finiteness conditions of $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$. The topologies of $\varphi^{3}$-$3\varphi$ correlation function are sketched in Fig.(5).
$\Gamma^{\tiny\mbox{L0}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-XII}_{1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-XII}_{2}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-XII}_{3}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$
Figure 5: The tree-level, 1-loop and 2-loop irreducible 1PI topologies of $\varphi^{3}$-$3\varphi$ correlation function. The topologies in the first row are equivalent to those corresponding to the $\varphi^{2}$-$2\varphi$ correlation function. There are only 2 new irreducible topologies shown in the second row. A gray arrow points to the cut graphs of each new irreducible topology, which are generated by the non-isomorphic computation of all possible removals of two external legs.
Notice that the loop topologies in the first row are equivalent to the $\varphi^{2}\to 2\varphi$ irreducible topologies we have already computed. There are only 2 new irreducible topologies $\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi{\varphi}\varphi}$, $\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi{\varphi}\varphi}$ at 2-loops for $\varphi^{3}\to 3\varphi$. Again from eqn.(35) we get the explicit expression for $\varphi^{3}$ as
$G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})=1+C_{2}^{1}\sum_{j=1}^{L}G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})+(2!)C_{2}^{2}\sum_{j=2}^{L}G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{3})~.$ (81)
Since $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{1-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ and $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ have already been computed in $\varphi^{2}$ case as (66) and (77), in order to compute 2-loop $\widetilde{Z}_{\varphi^{3}}$-factor, we only need to compute the additional contributions from these 2 new irreducible topologies of $\varphi^{3}\to 3\varphi$. They generate in total 4 cut graphs as shown in Fig.(5). Their symmetry factors and coupling factors are given by
$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~~~,~~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{3}h_{0}~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}~.$ (82)
The UV divergences computed by TwoPoint command are,
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon}+\frac{17}{8}+\epsilon\left(\frac{171}{32}-\frac{\pi^{2}}{24}\right)+\mathcal{O}(\epsilon^{2})~,$ (83)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XII}_{1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{163}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(-\frac{\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{1161}{128}\right)+\mathcal{O}(\epsilon^{2})~,$ (84)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XII}_{2}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=-\frac{1}{2\epsilon^{2}}-\frac{15}{8\epsilon}+\left(\frac{\pi^{2}}{24}-\frac{137}{32}\right)+\epsilon\left(\frac{\zeta_{3}}{12}+\frac{5\pi^{2}}{32}-\frac{971}{128}\right)+\mathcal{O}(\epsilon^{2})~,$ (85)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XII}_{3}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{171}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(-\frac{\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{1313}{128}\right)+\mathcal{O}(\epsilon^{2})~.$ (86)
Then the required UV contribution of all 2-loop irreducible topologies for $\varphi^{3}\to 3\varphi$ is
$G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{3})=\sum_{L={\tiny\mbox{L2-XI}},{\tiny\mbox{L2-XII}}_{1}\atop{\tiny\mbox{L2-XII}}_{2},{\tiny\mbox{L2-XII}}_{3}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})~.$ (87)
With the previously computed $\widetilde{Z}_{\varphi^{2}}$-factor, we can take the series expansion for $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$, and get the following equations from the vanishing of $1/\epsilon$, $1/\epsilon^{2}$ poles in the leading Log term and $1/\epsilon$ pole in the sub-Log term,
$0=z_{\varphi^{3}}^{21}+(4g^{2}+1)z_{\varphi^{3}}^{11}+\left(3g^{2}-\frac{7g^{4}N}{48}+\frac{355g^{4}}{24}+\frac{5g^{3}h}{4}-\frac{7g^{2}h^{2}}{48}\right)~,$ (88)
$0=z_{\varphi^{3}}^{22}+3g^{2}z_{\varphi^{3}}^{11}+\left(\frac{g^{4}N}{4}+\frac{5g^{4}}{2}-3g^{3}h+\frac{g^{2}h^{2}}{4}\right)~~~,~~~0=2g^{2}z_{\varphi^{3}}^{11}+6g^{4}~.$ (89)
By solving these 3 equations we completely determine the 2-loop $\widetilde{Z}_{\varphi^{3}}$-factor as
$z_{\varphi^{3}}^{11}=-3g^{2}~~,~~z_{\varphi^{3}}^{21}=\frac{7g^{4}N}{48}-\frac{67g^{4}}{24}-\frac{5g^{3}h}{4}+\frac{7g^{2}h^{2}}{48}~~,~~z_{\varphi^{3}}^{22}=-\frac{g^{4}N}{4}+\frac{13g^{4}}{2}+3g^{3}h-\frac{g^{2}h^{2}}{4}~.$ (90)
With results (67), (80) and (90), the 2-loop $\widetilde{Z}_{\varphi^{2}}$, $\widetilde{Z}_{\varphi^{3}}$-factors are finally determined as,
$\widetilde{Z}_{\varphi^{2}}=1+\frac{z_{\varphi^{2}}^{11}}{\epsilon}+\frac{z_{\varphi^{2}}^{21}}{\epsilon}+\frac{z_{\varphi^{2}}^{22}}{\epsilon^{2}}~~~,~~~\widetilde{Z}_{\varphi^{3}}=1+\frac{z_{\varphi^{3}}^{11}}{\epsilon}+\frac{z_{\varphi^{3}}^{21}}{\epsilon}+\frac{z_{\varphi^{3}}^{22}}{\epsilon^{2}}~.$ (91)
The gamma functions of $\varphi^{Q}$ can be computed by $\gamma_{\varphi^{Q}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}$, hence we get
$\gamma_{\varphi^{2}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial h}\beta(h)=-g^{2}+\left(\frac{7g^{4}N}{72}-\frac{49g^{4}}{36}-\frac{g^{3}h}{2}+\frac{7g^{2}h^{2}}{72}\right)~,$ (92)
$\gamma_{\varphi^{3}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial h}\beta(h)=-3g^{2}+\left(\frac{7g^{4}N}{24}-\frac{67g^{4}}{12}-\frac{5g^{3}h}{2}+\frac{7g^{2}h^{2}}{24}\right)~,$ (93)
where $\beta(g),\beta(h)$ have been provided in (61), (62). By the ansatz that the 2-loop $\gamma_{\varphi^{Q}}$ takes the form,
$\gamma_{\varphi^{Q}}=c_{1}Q(Q-1)(Q-2)+c_{2}Q(Q-1)~,$ (94)
and from results $\gamma_{\varphi^{2}},\gamma_{\varphi^{3}}$ we determine
$c_{1}=-\frac{g^{4}}{4}-\frac{g^{3}h}{6}~~~,~~~c_{2}=-\frac{g^{2}}{2}+\frac{7g^{4}N}{144}-\frac{49g^{4}}{72}-\frac{g^{3}h}{4}+\frac{7g^{2}h^{2}}{144}~.$ (95)
Then the scaling dimension of the $\varphi^{Q}$ operator can be calculated by its definition as,
$\Delta_{Q}=\left(\frac{d}{2}-1\right)Q+\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}~,$ (96)
where the anomalous dimension of the field $\varphi$ has been computed in (63).
4 Five-loop anomalous dimensions
The procedures described in previous section for 2-loop computations can be straightforwardly generalized to higher-loop calculations following exactly the same manner, with the primary difference being the need to consider a larger number of graphs. The $Z_{\varphi},Z_{\sigma}$-factors are determined from their corresponding two-point topologies in the manner analogous to (41). Similarly, the $Z_{h},Z_{g}$-factors are determined from the two-point integrals of cut graphs, generated by removing one external leg from the respective three-point topologies, following the approach outlined in (55). The beta functions are then computed from $\widetilde{Z}_{g},\widetilde{Z}_{h}$ using the definition (9). The field anomalous dimensions $\gamma_{\varphi}$ and $\gamma_{\sigma}$ can be computed from the $Z_{\varphi},Z_{\sigma}$-factors combined with the beta functions, following the definition (10). The $\widetilde{Z}_{\varphi^{Q}}$ factor is determined from the two-point integrals of cut graphs generated by removing (Q1)(Q-1) external legs from irreducible 1PI $\varphi^{Q}\to Q\varphi$ topologies, in a manner similar to that described in eqn.(35). The anomalous dimension of the $\varphi^{Q}$ operator is then given by $\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}$, where $\gamma_{\varphi^{Q}}$ is computed from $\widetilde{Z}_{\varphi^{Q}}$ and the beta functions by its definition. Determining the 5-loop $\gamma_{\varphi^{Q}}$ for general $Q$ requires the results of $L$-loop $\gamma_{\varphi^{Q_{L}}}$ with $Q_{L}=2,\ldots,L+1$ up to 5-loops. Since UV divergences are computed from scalar integrals, we do not need to consider the specific field configurations of Feynman diagrams, but only their topologies.
All topologies for two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions, as well as their cut graphs are generated by private Mathematica codes. The number of 1PI topologies for two-point and three-point functions, and irreducible correlation functions is listed in Table.(2).
$Q=2$ $Q=3$ $Q=4$ $Q=5$ $Q=6$ Total 2-point 3-point
$L=1$ 1 - - - - 1 1 1
$L=2$ 5 2 - - - 7 2 3
$L=3$ 35 26 7 - - 68 9 17
$L=4$ 303 358 168 25 - 854 46 125
$L=5$ 3136 5300 3709 1103 128 13376 322 1170
$Q=2$ $Q=3$ $Q=4$ $Q=5$ $Q=6$ Total 2-point 3-point
$L=1$ 1 - - - - 1 1 1
$L=2$ 7 4 - - - 11 2 5
$L=3$ 56 67 19 - - 142 9 35
$L=4$ 540 999 586 107 - 2232 46 303
$L=5$ 5905 15238 13846 5060 644 40693 322 3136
Table 2: (Top Table): the number of 1PI topologies of two-point, three-point correlation functions and the irreducible $\varphi^{Q}$-$Q\varphi$ correlation functions. (Bottom Table): the number of cut graphs, generated by the removal of one external leg for three-point function and (Q1)(Q-1) external legs for irreducible $\varphi^{Q}$-$Q\varphi$ correlation functions. These are the graphs to be computed by TwoPoint command for UV divergences.
The number grows rapidly as the number of loops increases. Cut graphs are generated through non-isomorphic computations of all possible external leg removals, and each of these cut graphs should be computed to determine their UV divergences. The number of cut graphs is listed in Table.(2). It can be seen that the number of cut graphs to be computed is about 3 to 4 times larger than the number of the original topologies.
The UV divergences are computed via TwoPoint command in HyperlogProcedures Maple package. The computational capability and time consumption of these evaluations are sensitive to both the loop order and the order of the dimensional regulator $\epsilon$. To apply the UV finiteness conditions for $L_{\tiny\mbox{max}}$-loop computation, an $L$-loop divergence in dimensional regularization should be computed up to the $\epsilon^{L_{\tiny\mbox{max}}-1-L}$ order. Increasing $L_{\tiny\mbox{max}}$ by one order requires UV divergence computations with an additional order in $\epsilon$ for all loop integrals, which significantly increases time consumption. For 5-loop computations, an $L$-loop divergence should be computed to the $\epsilon^{4-L}$ order, and the current version of HyperlogProcedures is capable of evaluating all required two-point integrals. Time consumption varies depending on the computer’s CPU and memory usage. On a normal desktop with a CPU base frequency 3.20GHz and 64GB memory, the approximate time required to evaluate all loop integrals is listed in Table.(3).
$Q=2$ $Q=3$ $Q=4$ $Q=5$ $Q=6$ Total 2-point 3-point
$L=1$ 0 - - - - 0 0 0
$L=2$ 0.02 h 0 - - - 0.02 h 0 0
$L=3$ 0.6 h 0.2 h 0 - - 0.8 h 0.03 h 0.15 h
$L=4$ 7 h 5.5 h 1.5 h 0 - 14 h 2.35 h 5.5 h
$L=5$ 16 d 24 d 10 d 1 d 0.2 h 51 d 5 d 8 d
Table 3: The approximate time consumption for 5-loop computations in a normal desktop. The symbols h and d are abbreviation of computer core hours and days. The working time less than 0.01 hour is simply denoted by 0.
Time consumption also increases rapidly as the loop order rises. However, for 5-loop computations, it remains within a reasonably acceptable range. In contrast, the symmetry factors and coupling factors for these topologies are generated almost instantaneously compared to the time required for evaluating UV divergences.
We have computed $Z_{\varphi}$, $Z_{\sigma}$, $Z_{h}$, $Z_{g}$ and all $\widetilde{Z}_{\varphi^{Q_{L}}},Q_{L}=2,\ldots,L+1$ up to 5-loops. From these results, we have derived the beta functions $\beta(g),\beta(h)$, the field anomalous dimensions $\gamma_{\varphi}$, $\gamma_{\sigma}$, and the scaling dimensions of the $\varphi^{Q}$} operator up to 5-loops. Results of lower than 5-loops have been compared with existing literature with agreements Fei:2014xta ; Gracey:2015tta ; Kompaniets:2021hwg .
The 5-loop correction to the anomalous dimension $\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}$ is presented in Appendix §A. The scaling dimension of the $\varphi^{Q}$ operator at the Wilson-Fisher fixed point can be computed as
$\Delta_{Q\tiny\mbox{FP}}=\left(2-\frac{\epsilon}{2}\right)Q+\left(\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}\right)\Big{|}_{g=g^{\ast},h=h^{\ast}}~,$ (97)
where $g^{\ast},h^{\ast}$ are the fixed-point solutions of the beta functions. Our 5-loop fixed-point solutions, provided in Appendix §B, are in perfect agreement with the known results in the literature Kompaniets:2021hwg . By substituting these fixed-point solutions into the anomalous dimensions, we obtain the $\Delta_{Q}$ in the large $N$ limit. The leading, sub-leading and sub-sub-leading order contributions of the large $N$ expansion are as follows,
$\Delta_{Q\tiny\mbox{FP}}=\left(2-\frac{\epsilon}{2}\right)Q+\frac{Q}{N}\Delta_{Q\tiny\mbox{FP}}^{1}+\frac{Q}{N^{2}}\Delta_{Q\tiny\mbox{FP}}^{2}+\frac{Q}{N^{3}}\Delta_{Q\tiny\mbox{FP}}^{3}+\frac{Q}{N^{4}}\Delta_{Q\tiny\mbox{FP}}^{4}+\frac{Q}{N^{5}}\Delta_{Q\tiny\mbox{FP}}^{5}+\mathcal{O}(\frac{1}{N^{6}})~,$ (98)
where
$\Delta_{Q{\tiny\mbox{FP}}}^{1}$ $=$ $\bigg{[}4-3Q\bigg{]}\epsilon+\bigg{[}-\frac{8}{3}+\frac{7Q}{4}\bigg{]}\epsilon^{2}+\bigg{[}-\frac{7}{9}+\frac{11Q}{16}\bigg{]}\epsilon^{3}$ (99)
$+\bigg{[}\zeta_{3}-\frac{8}{27}+\left(\frac{19}{64}-\frac{3\zeta_{3}}{4}\right)Q\bigg{]}\epsilon^{4}+\bigg{[}\frac{\pi^{4}}{120}-\frac{43}{324}-\frac{2\zeta_{3}}{3}+\left(\frac{7\zeta_{3}}{16}-\frac{\pi^{4}}{160}+\frac{35}{256}\right)Q\bigg{]}\epsilon^{5}~,$
$\Delta_{Q{\tiny\mbox{FP}}}^{2}$ $=$ $\bigg{[}176-132Q\bigg{]}\epsilon+\bigg{[}-\frac{1568}{3}+\frac{857Q}{2}-45Q^{2}\bigg{]}\epsilon^{2}$ (100)
$+\bigg{[}\frac{4105}{18}+72\zeta_{3}-\left(108\zeta_{3}+\frac{3743}{24}\right)Q+\left(36\zeta_{3}+\frac{93}{4}\right)Q^{2}\bigg{]}\epsilon^{3}$
$+\bigg{[}74\zeta_{3}+\frac{3\pi^{4}}{5}+\frac{19315}{216}+\left(21\zeta_{3}-\frac{9\pi^{4}}{10}-\frac{7465}{72}\right)Q+\left(\frac{3\pi^{4}}{10}-51\zeta_{3}+\frac{167}{8}\right)Q^{2}\bigg{]}\epsilon^{4}$
$+\bigg{[}\frac{12641}{324}-\frac{2281\zeta_{3}}{6}+\frac{37\pi^{4}}{60}+36\zeta_{5}+\left(\frac{577\zeta_{3}}{2}-54\zeta_{5}+\frac{7\pi^{4}}{40}-\frac{44459}{864}\right)Q$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(18\zeta_{5}-\frac{65\zeta_{3}}{4}+\frac{339}{32}-\frac{17\pi^{4}}{40}\right)Q^{2}\bigg{]}\epsilon^{5}~,$
$\Delta_{Q{\tiny\mbox{FP}}}^{3}$ $=$ $\bigg{[}7744-5808Q\bigg{]}\epsilon+\bigg{[}-66872+59520Q-9000Q^{2}\bigg{]}\epsilon^{2}$ (101)
$+\bigg{[}\frac{193141}{2}-4212\zeta_{3}-\bigg{(}3996\zeta_{3}+88575\bigg{)}Q+\bigg{(}4536\zeta_{3}+20538\bigg{)}Q^{2}-1350Q^{3}\bigg{]}\epsilon^{3}$
$+\bigg{[}61920\zeta_{3}-14040\zeta_{5}-\frac{351\pi^{4}}{10}-\frac{107537}{216}+\bigg{(}14580\zeta_{5}-22146\zeta_{3}-\frac{63913}{6}-\frac{333\pi^{4}}{10}\bigg{)}Q$
$~~~~~~~~~~~~~~~~~~~~~~+\bigg{(}\frac{2101}{2}-18558\zeta_{3}-1080\zeta_{5}+\frac{189\pi^{4}}{5}\bigg{)}Q^{2}+\bigg{(}\frac{1107}{2}+1728\zeta_{3}+540\zeta_{5}\bigg{)}Q^{3}\bigg{]}\epsilon^{4}$
$+\bigg{[}516\pi^{4}-540\zeta_{3}^{2}-166574\zeta_{3}-\frac{35397\zeta_{5}}{2}-\frac{130\pi^{6}}{7}-\frac{14477777}{864}$
$~~~~~~~~+\bigg{(}\frac{9276203}{432}-2106\zeta_{3}^{2}+\frac{245019\zeta_{3}}{2}-\frac{981\zeta_{5}}{2}+\frac{135\pi^{6}}{7}-\frac{3691\pi^{4}}{20}\bigg{)}Q$
$~~~~~~~~~~~~~~~~~~~~~~+\bigg{(}3348\zeta_{3}^{2}+\frac{549\zeta_{3}}{2}+3663\zeta_{5}-\frac{10\pi^{6}}{7}-\frac{3093\pi^{4}}{20}-\frac{190189}{24}\bigg{)}Q^{2}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\bigg{(}\frac{7551}{8}-702\zeta_{3}^{2}-2322\zeta_{3}-1215\zeta_{5}+\frac{5\pi^{6}}{7}+\frac{72\pi^{4}}{5}\bigg{)}Q^{3}\bigg{]}\epsilon^{5}~,$
and expressions $\Delta_{Q{\tiny\mbox{FP}}}^{4},\Delta_{Q{\tiny\mbox{FP}}}^{5}$ are presented in Appendix §C. As a confirmation of these results, the 3-loop contribution up to $\epsilon^{3}$ order is consistent with literature findings Vasiliev:1981yc ; Vasiliev:1981dg ; Vasiliev:1982dc ; Jack:2021ziq  [^7] The value $a_{t}^{(3)}$ in paper Jack:2021ziq that computes $\Delta_{Q}^{(3)}$ is incorrect. After using the correct value $a_{t}^{(3)}=\frac{g^{6}}{3}+g^{5}h$, the $\Delta_{Q}^{(3)}$ and the $1/N^{3}$ term in their paper are consistent with the 3-loop order part of our results.. The corresponding terms $\frac{Q^{i+1}\epsilon^{i}}{N^{i}},\frac{Q^{i+1}\epsilon^{i+1}}{N^{i}},i=1,\ldots,5$ are also consistent with those reported in the semi-classical literature Antipin:2021jiw . The 4-loop contribution up to $1/N^{2}$ order has been previously provided in Huang:2024hsn , while the 5-loop contributions up to $1/N^{5}$ order presented here are new. Our results have been provided in ancillary files to the arXiv submission of this article.
5 Conclusion
In this paper, we apply the OPE method of renormalization proposed in Huang:2024hsn to evaluate the 5-loop scaling dimensions of the $\phi^{Q}$ operator in the six-dimensional cubic scalar theory. The result is derived from over 40 000 two-point propagator-type integrals computed using the graphical function method, combined with non-trivial summation among these integrals. We have successfully obtained the 5-loop correction to the anomalous dimension of the $\phi^{Q}$ operator, as well as the large $N$ expansion of the scaling dimension at the Wilson-Fisher fixed point up to the $1/N^{5}$ order. It stands for a new record for the scaling dimensions of $\phi^{Q}$ operator in cubic scalar theory. The result shows consistency with lower-loop perturbative calculations reported in other literature and aligns with the corresponding terms derived via semi-classical methods. Furthermore, the non-trivial cancellation of $\epsilon$ poles in the final result provides a strong consistency check validating its correctness. This work further demonstrates the efficiency and power of the OPE method in renormalization computations.
Efforts to apply the OPE method to address the challenges posed by more complex problems are currently underway. A promising project aims to surpass the 6-loop record in quartic scalar theory by computing the 7-loop scaling dimensions of the $\phi^{Q}$ operator. Such research would not only advance the precision of renormalization computations but also drive the evolution of modern techniques for multi-loop integral evaluations. Also as previously noted, the OPE method is highly general and extendable beyond scalar theories. It is interesting to investigate the renormalization computations by OPE method in other situations, such as the Gross-Neveu-Yukawa theory with fermions, or scalar-QED theory with gauge fields, etc. In these theories, Feynman integrals are no longer purely scalar integrals, but involve non-trivial numerators. While further developments may enable the computation of such non-trivial master integrals via graphical function method adapted to numerator structures, the most viable current approach relies on the computational algorithms Smirnov:2008iw ; Maierhofer:2017gsa ; Wu:2023upw ; Guan:2024byi based on the IBP method Tkachov:1981wb ; Chetyrkin:1981qh ; Laporta:2000dsw . Another potential challenge would be applying the OPE method to the anomalous dimensions of more intricate operators, such as those with higher derivatives, Lorentz indices, or in scenarios with operator mixing Derkachov:1997ch ; Cao:2021cdt . While the application of the OPE method remains straightforward in these cases Huang:2024hsn , practical computations demand more sophisticated techniques. It is interesting to explore the diverse applications of the OPE method in renormalization, and further enhance its capabilities.
Acknowledgments
We would like to thank Oliver Schnetz for stimulating conversations on the graphical function method and the usage of HyperlogProcedures package. This work is supported by the Science Challenge Project (No. TZ2025012), and NSAF No.U2330401.
Appendix A Five-loop corrections to the scaling dimensions
In this section, we collect our results of the scaling dimension $\Delta_{Q}$ in six-dimensional cubic scalar theory. Note that the 4-loop result has been presented in Huang:2024hsn . The 5-loop correction of $\Delta_{Q}$ is given by
$\Delta_{Q}^{{\tiny\mbox{5-loop}}}=Q\sum_{i=2}^{10}g^{i}h^{10-i}\delta^{5}_{i}~,$ (102)
where,
$\delta^{5}_{2}$ $=$ $-\frac{3336089}{13436928}+\frac{25882183Q}{161243136}+\left(-\frac{685}{7776}+\frac{15341Q}{248832}\right)\zeta_{3}+\left(-\frac{1177}{1866240}+\frac{23Q}{55296}\right)\pi^{4}$ (103)
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(\frac{311}{1296}-\frac{271Q}{1728}\right)\zeta_{5}+\left(\frac{5}{40824}-\frac{5Q}{54432}\right)\pi^{6}+\left(-\frac{1}{108}+\frac{Q}{144}\right)\zeta_{3}^{~2}~,$
$\delta^{5}_{3}$ $=$ $-\frac{133429553}{161243136}+\frac{26603591Q}{13436928}-\frac{228283Q^{2}}{279936}+\left(\frac{231613}{62208}-\frac{28829Q}{6912}+\frac{13091Q^{2}}{15552}\right)\zeta_{3}+\left(-\frac{241}{46080}\right$. (104)
$\left.+\frac{121Q}{11520}-\frac{2599Q^{2}}{622080}\right)\pi^{4}+\left(-\frac{29}{16}+\frac{421Q}{576}+\frac{719Q^{2}}{1728}\right)\zeta_{5}+\left(-\frac{5}{3402}+\frac{5Q}{4536}\right)\pi^{6}+\left(\frac{1}{9}-\frac{Q}{12}\right)\zeta_{3}^{~2}~,$
$\delta^{5}_{4}$ $=$ $\frac{4152458077}{241864704}+\frac{1256311N}{20155392}-\frac{1679472427Q}{80621568}-\frac{2125781NQ}{53747712}+\frac{24768415Q^{2}}{4478976}-\frac{20291Q^{3}}{41472}$ (105)
$+\left(\frac{104597}{3456}-\frac{65N}{2592}-\frac{2478701Q}{62208}+\frac{3919NQ}{248832}+\frac{169543Q^{2}}{15552}-\frac{131Q^{3}}{144}\right)\zeta_{3}+\left(-\frac{31489}{1866240}\right$.
$\left.-\frac{11N}{622080}+\frac{307Q}{13824}+\frac{11NQ}{497664}-\frac{2023Q^{2}}{311040}+\frac{Q^{3}}{1080}\right)\pi^{4}+\left(-\frac{33983}{2592}+\frac{N}{324}+\frac{6473Q}{288}-\frac{NQ}{432}\right$.
$\left.-\frac{3041Q^{2}}{288}+\frac{85Q^{3}}{72}\right)\zeta_{5}+\left(\frac{205}{20412}-\frac{25Q}{1944}+\frac{25Q^{2}}{6804}-\frac{5Q^{3}}{13608}\right)\pi^{6}+\left(\frac{1195}{216}-\frac{323Q}{72}-\frac{37Q^{2}}{36}\right$.
$\left.+\frac{4Q^{3}}{9}\right)\zeta_{3}^{~2}+\left(-\frac{5467}{96}+\frac{6055Q}{96}-\frac{147Q^{2}}{16}\right)\zeta_{7}~,$
$\delta^{5}_{5}$ $=$ $-\frac{705928679}{26873856}+\frac{25722139N}{161243136}+\frac{519118363Q}{13436928}-\frac{3257NQ}{7776}-\frac{38531567Q^{2}}{2239488}+\frac{129851NQ^{2}}{746496}$ (106)
$+\frac{10871Q^{3}}{3072}-\frac{641Q^{4}}{2304}+\left(-\frac{1604333}{31104}-\frac{1145N}{1536}+\frac{581737Q}{6912}+\frac{23861NQ}{20736}-\frac{1196249Q^{2}}{31104}\right$.
$\left.-\frac{3719NQ^{2}}{10368}+\frac{7727Q^{3}}{1152}-\frac{11Q^{4}}{24}\right)\zeta_{3}+\left(\frac{4411}{155520}+\frac{607N}{124416}-\frac{11137Q}{207360}-\frac{17NQ}{2592}+\frac{1997Q^{2}}{62208}\right$.
$\left.+\frac{43NQ^{2}}{23040}-\frac{161Q^{3}}{34560}\right)\pi^{4}+\left(\frac{12157}{576}+\frac{43N}{216}-\frac{15383Q}{576}-\frac{49NQ}{192}+\frac{9767Q^{2}}{1728}+\frac{107NQ^{2}}{1728}\right$.
$\left.+\frac{5Q^{3}}{9}-\frac{Q^{4}}{3}\right)\zeta_{5}+\left(\frac{305}{54432}+\frac{25Q}{36288}-\frac{95Q^{2}}{15552}+\frac{5Q^{3}}{3402}\right)\pi^{6}+\left(-\frac{223}{144}-\frac{371Q}{96}+\frac{1723Q^{2}}{288}\right$.
$\left.-\frac{10Q^{3}}{9}\right)\zeta_{3}^{~2}+\left(\frac{917}{16}-\frac{581Q}{6}+\frac{2401Q^{2}}{48}-\frac{35Q^{3}}{3}+\frac{7Q^{4}}{6}\right)\zeta_{7}~,$
$\delta^{5}_{6}$ $=$ $\frac{101332127}{1889568}-\frac{377140433N}{80621568}+\frac{20375N^{2}}{13436928}-\frac{151156525Q}{1679616}+\frac{923790881NQ}{161243136}-\frac{84245N^{2}Q}{80621568}+\frac{3532885Q^{2}}{69984}$ (107)
$-\frac{3589273NQ^{2}}{2239488}-\frac{14759Q^{3}}{1024}+\frac{9473NQ^{3}}{82944}+\frac{545Q^{4}}{256}-\frac{7Q^{5}}{48}+\left(\frac{1035907}{7776}+\frac{4147N}{62208}+\frac{41N^{2}}{46656}-\frac{6811451Q}{31104}\right$.
$\left.+\frac{44599NQ}{62208}-\frac{197N^{2}Q}{248832}+\frac{1258969Q^{2}}{10368}-\frac{36589NQ^{2}}{62208}-\frac{19235Q^{3}}{576}+\frac{17NQ^{3}}{144}+\frac{65Q^{4}}{24}\right)\zeta_{3}$
$+\left(\frac{50047}{466560}-\frac{719N}{311040}-\frac{17N^{2}}{1866240}-\frac{35393Q}{311040}+\frac{7NQ}{3072}+\frac{17N^{2}Q}{2488320}+\frac{349Q^{2}}{20736}-\frac{103NQ^{2}}{622080}+\frac{73Q^{3}}{17280}-\frac{NQ^{3}}{2160}\right)\pi^{4}$
$+\left(\frac{26623}{144}+\frac{7325N}{648}-\frac{296567Q}{864}-\frac{25771NQ}{1728}+\frac{186683Q^{2}}{864}+\frac{851NQ^{2}}{192}-\frac{1435Q^{3}}{24}-\frac{65NQ^{3}}{144}+\frac{35Q^{4}}{6}\right)\zeta_{5}$
$+\left(\frac{415}{163296}-\frac{185N}{40824}+\frac{5Q}{2268}+\frac{55NQ}{9072}-\frac{295Q^{2}}{54432}-\frac{25NQ^{2}}{13608}+\frac{25Q^{3}}{13608}+\frac{5NQ^{3}}{27216}\right)\pi^{6}$
$+\left(-\frac{15473}{432}-\frac{367N}{216}+71Q+\frac{53NQ}{24}-\frac{5899Q^{2}}{144}-\frac{11NQ^{2}}{18}+\frac{223Q^{3}}{36}+\frac{NQ^{3}}{36}\right)\zeta_{3}^{~2}$
$+\left(-\frac{8827}{24}+\frac{19817Q}{32}-\frac{17129Q^{2}}{48}+\frac{413Q^{3}}{4}-\frac{35Q^{4}}{3}\right)\zeta_{7}~,$
$\delta^{5}_{7}$ $=$ $-\frac{1727695009}{13436928}+\frac{1219348333N}{80621568}-\frac{2802343N^{2}}{53747712}+\frac{1827570041Q}{6718464}-\frac{38406623NQ}{2239488}+\frac{42845N^{2}Q}{2239488}$ (108)
$-\frac{360106991Q^{2}}{2239488}+\frac{22560479NQ^{2}}{4478976}+\frac{35173N^{2}Q^{2}}{4478976}+\frac{1125739Q^{3}}{41472}-\frac{20573NQ^{3}}{20736}+\frac{3223Q^{4}}{4608}+\frac{133NQ^{4}}{2304}-\frac{7Q^{5}}{16}$
$+\left(-\frac{48245}{3888}+\frac{1642177N}{31104}+\frac{3011N^{2}}{62208}+\frac{128345Q}{972}-\frac{4686905NQ}{62208}-\frac{331N^{2}Q}{5184}-\frac{1930387Q^{2}}{10368}+\frac{406627NQ^{2}}{15552}\right$.
$\left.+\frac{1165N^{2}Q^{2}}{62208}+\frac{14215Q^{3}}{192}-\frac{1451NQ^{3}}{576}-\frac{141Q^{4}}{16}\right)\zeta_{3}+\left(-\frac{177979}{311040}-\frac{71161N}{622080}-\frac{95N^{2}}{248832}+\frac{162677Q}{155520}\right$.
$\left.+\frac{103393NQ}{622080}+\frac{17N^{2}Q}{23040}-\frac{19751Q^{2}}{34560}-\frac{3713NQ^{2}}{62208}-\frac{91N^{2}Q^{2}}{311040}+\frac{1781Q^{3}}{17280}+\frac{89NQ^{3}}{17280}\right)\pi^{4}+\left(-\frac{2030531}{1728}\right$.
$\left.-\frac{47359N}{576}-\frac{N^{2}}{144}+\frac{1301549Q}{576}+\frac{203777NQ}{1728}+\frac{N^{2}Q}{96}-\frac{824411Q^{2}}{576}-\frac{26117NQ^{2}}{576}-\frac{N^{2}Q^{2}}{288}+\frac{6655Q^{3}}{18}\right$.
$\left.+\frac{445NQ^{3}}{72}-\frac{110Q^{4}}{3}\right)\zeta_{5}+\left(\frac{3725}{27216}+\frac{55N}{3402}-\frac{8185Q}{36288}-\frac{3085NQ}{108864}+\frac{11695Q^{2}}{108864}+\frac{1405NQ^{2}}{108864}-\frac{205Q^{3}}{13608}\right$.
$\left.-\frac{25NQ^{3}}{13608}\right)\pi^{6}+\left(\frac{9959}{72}+\frac{629N}{72}-\frac{22685Q}{96}-\frac{3493NQ}{288}+\frac{34195Q^{2}}{288}+\frac{1273NQ^{2}}{288}-\frac{685Q^{3}}{36}-\frac{11NQ^{3}}{18}\right)\zeta_{3}^{~2}$
$+\left(\frac{12887}{12}-\frac{147N}{16}-\frac{18291Q}{8}+\frac{147NQ}{16}+\frac{9793Q^{2}}{6}-\frac{3731Q^{3}}{8}+\frac{140Q^{4}}{3}\right)\zeta_{7}~,$
$\delta^{5}_{8}$ $=$ $\frac{26586982301}{60466176}-\frac{779370073N}{60466176}+\frac{26353453N^{2}}{80621568}-\frac{3827N^{3}}{20155392}-\frac{8024021965Q}{10077696}+\frac{79115735NQ}{5038848}+\frac{1162525N^{2}Q}{20155392}$ (109)
$+\frac{2029N^{3}Q}{17915904}+\frac{514446007Q^{2}}{1119744}-\frac{4617385NQ^{2}}{559872}-\frac{25159N^{2}Q^{2}}{165888}-\frac{2234915Q^{3}}{20736}+\frac{237013NQ^{3}}{82944}+\frac{127N^{2}Q^{3}}{27648}$
$+\frac{15847Q^{4}}{1536}-\frac{163NQ^{4}}{768}-\frac{3Q^{5}}{4}+\left(-\frac{8364127}{15552}-\frac{929849N}{23328}-\frac{18653N^{2}}{31104}-\frac{43N^{3}}{31104}+\frac{7945621Q}{7776}+\frac{2483159NQ}{31104}\right$.
$\left.+\frac{33251N^{2}Q}{62208}+\frac{245N^{3}Q}{248832}-\frac{8598169Q^{2}}{15552}-\frac{163897NQ^{2}}{3456}-\frac{1465N^{2}Q^{2}}{20736}+\frac{9743Q^{3}}{108}+\frac{899NQ^{3}}{144}-\frac{149Q^{4}}{24}\right$.
$\left.+\frac{3NQ^{4}}{8}\right)\zeta_{3}+\left(\frac{18179}{15552}+\frac{89461N}{311040}+\frac{2899N^{2}}{622080}+\frac{N^{3}}{622080}-\frac{314647Q}{155520}-\frac{145487NQ}{311040}-\frac{4417N^{2}Q}{622080}-\frac{N^{3}Q}{829440}\right$.
$\left.+\frac{31291Q^{2}}{31104}+\frac{7607NQ^{2}}{34560}+\frac{481N^{2}Q^{2}}{207360}-\frac{151Q^{3}}{1152}-\frac{29NQ^{3}}{864}\right)\pi^{4}+\left(\frac{6140413}{2592}+\frac{784319N}{5184}+\frac{43N^{2}}{864}-\frac{4206853Q}{864}\right$.
$\left.-\frac{406439NQ}{1728}-\frac{7N^{2}Q}{64}+\frac{2890897Q^{2}}{864}+\frac{7891NQ^{2}}{72}+\frac{103N^{2}Q^{2}}{1728}-\frac{44575Q^{3}}{48}-\frac{1735NQ^{3}}{96}+\frac{260Q^{4}}{3}\right)\zeta_{5}$
$+\left(-\frac{9425}{40824}-\frac{7015N}{163296}+\frac{7565Q}{18144}+\frac{4255NQ}{54432}-\frac{3995Q^{2}}{18144}-\frac{325NQ^{2}}{7776}+\frac{965Q^{3}}{27216}+\frac{365NQ^{3}}{54432}\right)\pi^{6}$
$+\left(-\frac{5089}{216}-\frac{5359N}{432}+\frac{2083Q}{48}+\frac{4207NQ}{144}-\frac{1003Q^{2}}{48}-\frac{2935NQ^{2}}{144}+\frac{155Q^{3}}{72}+\frac{581NQ^{3}}{144}\right)\zeta_{3}^{~2}$
$+\left(-\frac{200585}{96}-\frac{1631N}{24}+\frac{415247Q}{96}+\frac{371NQ}{6}-\frac{148813Q^{2}}{48}+\frac{11123Q^{3}}{12}-\frac{280Q^{4}}{3}\right)\zeta_{7}~,$
$\delta^{5}_{9}$ $=$ $-\frac{4988268911}{10077696}+\frac{158410849N}{4478976}-\frac{1951205N^{2}}{20155392}+\frac{162107N^{3}}{161243136}+\frac{2881576087Q}{3359232}-\frac{213374875NQ}{6718464}-\frac{720205N^{2}Q}{746496}$ (110)
$-\frac{1213N^{3}Q}{2239488}-\frac{237912325Q^{2}}{559872}+\frac{2904071NQ^{2}}{746496}+\frac{2387851N^{2}Q^{2}}{4478976}+\frac{11N^{3}Q^{2}}{1492992}+\frac{146737Q^{3}}{2592}+\frac{2015NQ^{3}}{13824}-\frac{209N^{2}Q^{3}}{82944}$
$+\frac{673Q^{4}}{144}-\frac{1915NQ^{4}}{4608}-\frac{5Q^{5}}{6}+\left(\frac{11111}{32}+\frac{992585N}{15552}+\frac{47023N^{2}}{31104}+\frac{1397N^{3}}{124416}-\frac{14576275Q}{15552}-\frac{6227NQ}{54}-\frac{5197N^{2}Q}{7776}\right$.
$\left.-\frac{23N^{3}Q}{3456}+\frac{2176717Q^{2}}{2592}+\frac{674003NQ^{2}}{10368}-\frac{659N^{2}Q^{2}}{10368}-\frac{7N^{3}Q^{2}}{6912}-\frac{39937Q^{3}}{144}-\frac{6193NQ^{3}}{576}-\frac{35N^{2}Q^{3}}{384}+\frac{215Q^{4}}{8}\right$.
$\left.-\frac{5NQ^{4}}{24}\right)\zeta_{3}+\left(\frac{1703}{5184}-\frac{30931N}{62208}-\frac{203N^{2}}{69120}-\frac{N^{3}}{51840}-\frac{12353Q}{31104}+\frac{4771NQ}{5760}+\frac{233N^{2}Q}{31104}+\frac{N^{3}Q}{69120}+\frac{1427Q^{2}}{25920}\right$.
$\left.-\frac{511NQ^{2}}{1280}-\frac{41N^{2}Q^{2}}{10368}+\frac{7Q^{3}}{240}+\frac{73NQ^{3}}{1152}+\frac{N^{2}Q^{3}}{2304}\right)\pi^{4}+\left(-\frac{1144559}{432}-\frac{321965N}{1728}-\frac{67N^{2}}{96}+\frac{599057Q}{108}\right$.
$\left.+\frac{534791NQ}{1728}+\frac{847N^{2}Q}{864}-\frac{1665455Q^{2}}{432}-\frac{93157NQ^{2}}{576}-\frac{37N^{2}Q^{2}}{72}+\frac{76435Q^{3}}{72}+\frac{235NQ^{3}}{9}-105Q^{4}+\frac{5NQ^{4}}{6}\right)\zeta_{5}$
$+\left(-\frac{155}{6804}+\frac{2735N}{27216}-\frac{25N^{2}}{54432}+\frac{55Q}{1512}-\frac{18325NQ}{108864}+\frac{25N^{2}Q}{54432}-\frac{5Q^{2}}{432}+\frac{2995NQ^{2}}{36288}+\frac{5Q^{3}}{2268}-\frac{5NQ^{3}}{378}\right)\pi^{6}$
$+\left(-\frac{3037}{9}+\frac{2369N}{72}-\frac{19N^{2}}{144}+\frac{6817Q}{12}-\frac{16729NQ}{288}+\frac{19N^{2}Q}{144}-\frac{6445Q^{2}}{24}+\frac{2951NQ^{2}}{96}+\frac{112Q^{3}}{3}-5NQ^{3}\right)\zeta_{3}^{~2}$
$+\left(\frac{78659}{24}-\frac{203N}{12}-\frac{299551Q}{48}+\frac{259NQ}{16}+\frac{181027Q^{2}}{48}+\frac{35NQ^{2}}{48}-\frac{21077Q^{3}}{24}+70Q^{4}\right)\zeta_{7}~,$
$\delta^{5}_{10}$ $=$ $\frac{3873526685}{15116544}-\frac{1261849117N}{60466176}-\frac{1111975N^{2}}{30233088}+\frac{2132959N^{3}}{241864704}-\frac{43N^{4}}{2519424}-\frac{4120586761Q}{10077696}+\frac{217226153NQ}{6718464}$ (111)
$+\frac{24155359N^{2}Q}{26873856}-\frac{1574323N^{3}Q}{161243136}+\frac{35N^{4}Q}{1990656}+\frac{12775867Q^{2}}{69984}-\frac{16632703NQ^{2}}{1119744}-\frac{239203N^{2}Q^{2}}{559872}+\frac{41N^{3}Q^{2}}{31104}$
$-\frac{3373Q^{3}}{144}+\frac{3283NQ^{3}}{1152}-\frac{133N^{2}Q^{3}}{4608}+\frac{7Q^{4}}{64}-\frac{469NQ^{4}}{1536}-\frac{Q^{5}}{2}+\left(\frac{2181835}{11664}-\frac{33815N}{1458}-\frac{246587N^{2}}{46656}-\frac{3619N^{3}}{186624}\right$.
$\left.-\frac{N^{4}}{11664}-\frac{2534795Q}{15552}+\frac{502549NQ}{5184}+\frac{119575N^{2}Q}{20736}+\frac{37N^{3}Q}{4608}+\frac{7N^{4}Q}{124416}-\frac{9701Q^{2}}{81}-\frac{390451NQ^{2}}{5184}-\frac{11453N^{2}Q^{2}}{7776}\right$.
$\left.+\frac{31N^{3}Q^{2}}{7776}+\frac{50953Q^{3}}{432}+\frac{4409NQ^{3}}{432}+\frac{143N^{2}Q^{3}}{576}-\frac{47Q^{4}}{2}-\frac{19NQ^{4}}{48}\right)\zeta_{3}+\left(-\frac{10277}{23328}+\frac{17249N}{58320}+\frac{667N^{2}}{933120}\right$.
$\left.-\frac{23N^{3}}{233280}+\frac{N^{4}}{933120}+\frac{27649Q}{31104}-\frac{153709NQ}{311040}-\frac{1493N^{2}Q}{311040}+\frac{97N^{3}Q}{622080}-\frac{N^{4}Q}{1244160}-\frac{199Q^{2}}{360}+\frac{13021NQ^{2}}{51840}\right$.
$\left.+\frac{137N^{2}Q^{2}}{38880}-\frac{17N^{3}Q^{2}}{311040}+\frac{77Q^{3}}{720}-\frac{137NQ^{3}}{3456}-\frac{13N^{2}Q^{3}}{17280}\right)\pi^{4}+\left(\frac{185473}{144}+\frac{88807N}{864}+\frac{28717N^{2}}{5184}+\frac{N^{3}}{216}\right$.
$\left.-\frac{378875Q}{144}-\frac{345655NQ}{1728}-\frac{3679N^{2}Q}{576}-\frac{N^{3}Q}{144}+\frac{387313Q^{2}}{216}+\frac{99619NQ^{2}}{864}+\frac{475N^{2}Q^{2}}{288}+\frac{N^{3}Q^{2}}{432}-\frac{3915Q^{3}}{8}\right$.
$\left.-\frac{2825NQ^{3}}{144}-\frac{5N^{2}Q^{3}}{32}+\frac{125Q^{4}}{3}-\frac{NQ^{4}}{3}\right)\zeta_{5}+\left(\frac{445}{3402}-\frac{145N}{2268}-\frac{3155Q}{13608}+\frac{6025NQ}{54432}+\frac{10Q^{2}}{81}-\frac{3055NQ^{2}}{54432}\right$.
$\left.-\frac{5N^{2}Q^{2}}{27216}-\frac{65Q^{3}}{3402}+\frac{25NQ^{3}}{3024}+\frac{5N^{2}Q^{3}}{54432}\right)\pi^{6}+\left(\frac{2342}{9}-\frac{209N}{24}+\frac{23N^{2}}{36}-\frac{4145Q}{9}+\frac{3271NQ}{144}-N^{2}Q+240Q^{2}\right$.
$\left.-\frac{2287NQ^{2}}{144}+\frac{31N^{2}Q^{2}}{72}-\frac{344Q^{3}}{9}+\frac{49NQ^{3}}{24}-\frac{13N^{2}Q^{3}}{144}\right)\zeta_{3}^{~2}+\left(-\frac{216181}{96}-\frac{567N}{32}+\frac{131411Q}{32}-\frac{105NQ}{4}\right$.
$\left.-2296Q^{2}+\frac{301NQ^{2}}{8}+\frac{3563Q^{3}}{8}-\frac{119NQ^{3}}{48}-\frac{49Q^{4}}{3}+\frac{7NQ^{4}}{6}\right)\zeta_{7}~.$
Appendix B Solutions at the Wilson-Fisher fixed point
The beta functions can be solved order by order in $\epsilon$, which leads to the solutions at fixed point as,
$g^{\ast}$ $=$ $\sqrt{\frac{6\epsilon}{N}}\bigg{\{}1+\frac{22}{N}+\frac{726}{N^{2}}-\frac{326180}{N^{3}}-\frac{349658330}{N^{4}}+\bigg{[}-\frac{155}{6N}-\frac{1705}{N^{2}}+\frac{912545}{N^{3}}+\frac{3590574890}{3N^{4}}\bigg{]}\epsilon+\bigg{[}\frac{1777}{144N}$ (112)
$+\frac{1}{N^{2}}\bigg{(}\frac{29093}{36}-1170\zeta_{3}\bigg{)}+\frac{1}{N^{3}}\bigg{(}-\frac{106755739}{72}-228060\zeta_{3}\bigg{)}+\frac{92543220\zeta_{3}-1920046163}{N^{4}}\bigg{]}\epsilon^{2}+\bigg{[}-\frac{217}{2592N}$
$+\frac{1}{N^{2}}\bigg{(}\frac{709151}{864}+3126\zeta_{3}-\frac{13\pi^{4}}{2}+360\zeta_{5}\bigg{)}+\frac{1}{N^{3}}\bigg{(}\frac{779869165}{432}+2229915\zeta_{3}-1267\pi^{4}-2438640\zeta_{5}\bigg{)}$
$+\frac{1}{N^{4}}\bigg{(}\frac{2437481884309}{1296}+278186790\zeta_{3}+864409\pi^{4}-1469512080\zeta_{5}\bigg{)}\bigg{]}\epsilon^{3}+\bigg{[}\frac{1}{N}\bigg{(}\frac{6973}{41472}-\frac{155\zeta_{3}}{96}\bigg{)}$
$+\frac{1}{N^{2}}\bigg{(}-\frac{49050023}{124416}-\frac{40871\zeta_{3}}{24}+\frac{1563\pi^{4}}{80}-4320\zeta_{5}+\frac{5\pi^{6}}{14}+81\zeta_{3}^{2}\bigg{)}+\frac{1}{N^{3}}\bigg{(}-\frac{17059272503}{15552}-\frac{77297317\zeta_{3}}{48}$
$+\frac{667967\pi^{4}}{48}+\frac{23014935\zeta_{5}}{4}-\frac{16935\pi^{6}}{7}-366930\zeta_{3}^{2}-\frac{2941029\zeta_{7}}{2}\bigg{)}+\frac{1}{N^{4}}\bigg{(}-\frac{68647291122307}{62208}+\frac{40450011887\zeta_{3}}{12}$
$-\frac{1345559\pi^{4}}{24}+\frac{13461052935\zeta_{5}}{2}-\frac{9381745\pi^{6}}{7}-37063008\zeta_{3}^{2}-6633588717\zeta_{7}\bigg{)}\bigg{]}\epsilon^{4}\bigg{\}}~,$
and
$h^{\ast}$ $=$ $6\sqrt{\frac{6\epsilon}{N}}\bigg{\{}1+\frac{162}{N}+\frac{68766}{N^{2}}+\frac{41224420}{N^{3}}+\frac{28762554870}{N^{4}}+\bigg{[}-\frac{215}{2N}-\frac{86335}{N^{2}}-\frac{75722265}{N^{3}}-\frac{69633402510}{N^{4}}\bigg{]}\epsilon+\bigg{[}\frac{927}{16N}$ (113)
$+\frac{1}{N^{2}}\bigg{(}\frac{270911}{6}-26190\zeta_{3}\bigg{)}+\frac{1}{N^{3}}\bigg{(}\frac{1156981601}{24}-26334180\zeta_{3}\bigg{)}+\frac{1}{N^{4}}\bigg{(}\frac{116908872997}{2}-23638808340\zeta_{3}\bigg{)}\bigg{]}\epsilon^{2}$
$+\bigg{[}-\frac{3461}{864N}+\frac{1}{N^{2}}\bigg{(}\frac{25209239}{864}+115860\zeta_{3}-\frac{291\pi^{4}}{2}-70200\zeta_{5}\bigg{)}+\frac{1}{N^{3}}\bigg{(}\frac{18166643735}{432}+198108185\zeta_{3}-175491\pi^{4}$
$-169538400\zeta_{5}\bigg{)}+\frac{1}{N^{4}}\bigg{(}\frac{5465613593485}{144}+235067598534\zeta_{3}-172334673\pi^{4}-211183698960\zeta_{5}\bigg{)}\bigg{]}\epsilon^{3}$
$+\bigg{[}\frac{1}{N}\bigg{(}\frac{5945}{13824}-\frac{215\zeta_{3}}{32}\bigg{)}+\frac{1}{N^{2}}\bigg{(}-\frac{1204206197}{41472}-\frac{973091\zeta_{3}}{8}+\frac{5793\pi^{4}}{8}+104895\zeta_{5}-\frac{975\pi^{6}}{14}-15795\zeta_{3}^{2}\bigg{)}+$
$\frac{1}{N^{3}}\bigg{(}-\frac{35896723781}{648}-\frac{1912729075\zeta_{3}}{16}+\frac{10821883\pi^{4}}{8}+\frac{2017087425\zeta_{5}}{4}-\frac{1245950\pi^{6}}{7}-23961852\zeta_{3}^{2}-\frac{637290423\zeta_{7}}{2}\bigg{)}$
$+\frac{1}{N^{4}}\bigg{(}-\frac{179864809668917}{2304}-\frac{552249422829\zeta_{3}}{4}+\frac{34006612103\pi^{4}}{20}+\frac{1850885090325\zeta_{5}}{2}-\frac{1654305365\pi^{6}}{7}$
$-15580792368\zeta_{3}^{2}-725681165454\zeta_{7}\bigg{)}\bigg{]}\epsilon^{4}\bigg{\}}~.$
Appendix C Large $N$ expansion at the Wilson-Fisher fixed point
Here we present the results of the $1/N^{4}$ and $1/N^{5}$ orders in the large $N$ expansion of $\gamma_{Q}$ at Wilson-Fisher fixed point. We have,
$\Delta_{Q{\tiny\mbox{FP}}}^{4}$ $=$ $\bigg{[}-2481664+1861248Q\bigg{]}\epsilon+\bigg{[}-\frac{21069760}{3}+8828264Q-3043440Q^{2}\bigg{]}\epsilon^{2}+\bigg{[}\frac{259232542}{9}-861840\zeta_{3}+\bigg{(}-30411798$ (114)
$-1662768\zeta_{3}\bigg{)}Q+\bigg{(}8984340+1478304\zeta_{3}\bigg{)}Q^{2}-495720Q^{3}\bigg{]}\epsilon^{3}+\bigg{[}-\frac{25851511}{4}+44545394\zeta_{3}-7182\pi^{4}-40335840\zeta_{5}$
$+\bigg{(}\frac{124128583}{24}-31054098\zeta_{3}-\frac{69282\pi^{4}}{5}+38233080\zeta_{5}\bigg{)}Q+\bigg{(}-4410989-2172420\zeta_{3}+\frac{61596\pi^{4}}{5}-4110480\zeta_{5}\bigg{)}Q^{2}$
$+\bigg{(}1024488+388854\zeta_{3}+186840\zeta_{5}\bigg{)}Q^{3}-\frac{213597Q^{4}}{4}\bigg{]}\epsilon^{4}+\bigg{[}-\frac{29834168837}{1296}-\frac{498959309\zeta_{3}}{6}+\frac{22272697\pi^{4}}{60}$
$+110252862\zeta_{5}-9004500\zeta_{3}^{2}-\frac{373480\pi^{6}}{7}-47146617\zeta_{7}+\bigg{(}\frac{23354482969}{864}+48023731\zeta_{3}-\frac{5175683\pi^{4}}{20}-89443629\zeta_{5}$
$+\frac{354010\pi^{6}}{7}+10983168\zeta_{3}^{2}+42115059\zeta_{7}\bigg{)}Q+\bigg{(}-\frac{27985893}{4}+8241600\zeta_{3}-\frac{36207\pi^{4}}{2}-1692738\zeta_{5}-\frac{38060\pi^{6}}{7}$
$-2645568\zeta_{3}^{2}+326592\zeta_{7}\bigg{)}Q^{2}+\bigg{(}\frac{1270377}{4}-3338073\zeta_{3}+\frac{64809\pi^{4}}{20}+1459530\zeta_{5}+\frac{1730\pi^{6}}{7}+11124\zeta_{3}^{2}-19278\zeta_{7}\bigg{)}Q^{3}$
$+\bigg{(}\frac{252477}{16}+92178\zeta_{3}+36288\zeta_{5}+9072\zeta_{7}\bigg{)}Q^{4}\bigg{]}\epsilon^{5}~,$
$\Delta_{Q{\tiny\mbox{FP}}}^{5}$ $=$ $\bigg{[}-2852566016+2139424512Q\bigg{]}\epsilon+\bigg{[}\frac{9747487616}{3}-424181248Q-1627608960Q^{2}\bigg{]}\epsilon^{2}+\bigg{[}\frac{77119431952}{9}$ (115)
$+1600458624\zeta_{3}+\bigg{(}-\frac{36719656688}{3}-2589857280\zeta_{3}\bigg{)}Q+\bigg{(}5365027104+905601024\zeta_{3}\bigg{)}Q^{2}-223974720Q^{3}\bigg{]}\epsilon^{3}$
$+\bigg{[}-\frac{18452489398}{3}+16928531152\zeta_{3}+\frac{66685776\pi^{4}}{5}-25646146560\zeta_{5}+\bigg{(}\frac{62653214582}{9}-12498672264\zeta_{3}$
$-21582144\pi^{4}+24847024320\zeta_{5}\bigg{)}Q+\bigg{(}-4161747404-280446528\zeta_{3}+\frac{37733376\pi^{4}}{5}-3130401600\zeta_{5}\bigg{)}Q^{2}$
$+\bigg{(}640256832+94633056\zeta_{3}+135475200\zeta_{5}\bigg{)}Q^{3}-28653588Q^{4}\bigg{]}\epsilon^{4}+\bigg{[}-\frac{12059984664415}{648}-\frac{98689756858\zeta_{3}}{3}$
$+\frac{2116066394\pi^{4}}{15}+141568450614\zeta_{5}-\frac{237464320\pi^{6}}{7}-4542718032\zeta_{3}^{2}-92577398445\zeta_{7}+\bigg{(}\frac{2294461732357}{108}$
$+21292069758\zeta_{3}-\frac{520778011\pi^{4}}{5}-123620853240\zeta_{5}+\frac{230065040\pi^{6}}{7}+7314253776\zeta_{3}^{2}+79133304537\zeta_{7}\bigg{)}Q$
$+\bigg{(}-\frac{27277501615}{6}-237707128\zeta_{3}-\frac{11685272\pi^{4}}{5}+8399505492\zeta_{5}-\frac{28985200\pi^{6}}{7}-2829860928\zeta_{3}^{2}$
$-1796643828\zeta_{7}\bigg{)}Q^{2}+\bigg{(}-215265924-1832415606\zeta_{3}+\frac{3943044\pi^{4}}{5}+1094776020\zeta_{5}+179200\pi^{6}+151863552\zeta_{3}^{2}$
$-230521788\zeta_{7}\bigg{)}Q^{3}+\bigg{(}56025189+32423328\zeta_{3}+10158048\zeta_{5}+10360224\zeta_{7}\bigg{)}Q^{4}-2457216Q^{5}\bigg{]}\epsilon^{5}~,$

### v4

@@
 Abstract
-Recently, an Operator Product Expansion method has been proposed for renormalization in quantum field theories.
-This method extracts UV divergences from two-point propagator-type integrals, and it is believed to be applicable to generic QFTs.
-In the present work, we apply the OPE method to the renormalization of scalar theory with cubic interactions in six-dimensions.
-We have successfully computed the 5-loop corrections to the anomalous dimensions of the $\phi^{Q}$ operator, as well as the large $N$ expansion of scaling dimensions at the Wilson-Fisher fixed point to $1/N^{5}$ order.
-This computation sets a new loop-order record for the anomalous dimension of the $\phi^{Q}$ operator in cubic scalar theory, and further demonstrates the efficiency and capacity of the OPE method in renormalization.
+In this work, we compute the anomalous dimensions of the $\phi^{Q}$ operator in six-dimensional cubic scalar theory.
+The renormalization analysis is carried out within the framework of the Operator Product Expansion method, while the ultraviolet divergences of Feynman integrals are evaluated using the graphical function method.
+Inspired by the intrinsic connection between Wilson coefficients and anomalous dimensions, an algorithm was proposed recently, which provides a practical and systematic framework for calculating the anomalous dimensions of masses, fields, and composite operators, with broad potential applicability to generic quantum field theories.
+Notably, the HyperlogProcedures package, developed based on the graphical function method, enables the computation of two-point propagator-type integrals, derived herein for capturing ultraviolet divergences, to very high loop orders.
+With these advanced techniques, we have successfully computed the anomalous dimensions of the $\phi^{Q}$ operator up to five loops.
+Furthermore, we present a large $N$ expansion of the scaling dimensions at the Wilson-Fisher fixed point, extended to the $1/N^{5}$ order.
+This computation sets a new loop-order record for the anomalous dimension of the $\phi^{Q}$ operator in cubic scalar theory, while further validating the efficiency and versatility of the proposed algorithm in renormalization analyses.
 Keywords:
@@
 1 Introduction
-During the development of the renormalization group method, scalar field theories have served as an ideal playground for testing multi-loop computation techniques and providing insights into more complicated field theories such as non-Abelian gauge theories.
-Yet the renormalization of scalar field theories is itself of great importance, as explored in Wilson’s theory of critical phenomena Wilson:1971dc , that in certain statistical systems, the critical exponents near second-order phase transitions are related to the anomalous dimensions of the field or composite operators constructed from fields PELISSETTO2002549 ; Moshe:2003xn .
-Two notable scalar field theories are the six-dimensional cubic scalar field theory and the four-dimensional quartic scalar theory, both of which are renormalizable theories.
-These two theories belong to the same $O(N)$ universality class, and it has been conjectured that the $O(N)$ quartic scalar theory at its UV fixed point is equivalent to the $O(N)$ cubic scalar theory at its IR fixed point Fei:2014yja ; Fei:2014xta , namely they approach to the same critical theory in the fixed points.
-Verifying this conjecture with high-loop precision for the scaling dimensions of fields or composite operators represents an intriguing problem, but it demands perturbative computation of renormalization functions to higher-loop orders.
-Because of the technical advancements, recently there has been a growing interest in pushing renormalization computations to the next loop order.
-It has been a long history for renormalization computations of quartic and cubic scalar theories, dating back to the early era studies of critical exponents up to $\epsilon^{2}$ orders for generalized Ising and Baxter models Wilson:1971dc , as well as the generalized Heisenberg model Wilson:1971vs .
-The 3-loop computations of the beta functions, field renormalization and mass renormalization of the quartic scalar theory were completed just one year after the 1-loop and 2-loop results BREZIN1973227 ; Brezin:1973igb , with 4-loop results following five years later Kazakov:1979ik .
-The full 5-loop renormalization of the quartic scalar theory required over a decade of collaborative efforts from numerous researchers Chetyrkin:1981jq ; Gorishnii:1983gp ; Kazakov:1983dyk ; Gorishnii:1983jr ; Kleinert:1991rg ; Adzhemyan:2013jra .
-Throughout this development, techniques such as the Gegenbauer polynomial $x$-space, IR-rearrangement Vladimirov:1979zm ; Chetyrkin:1980pr ; Caswell:1981ek , $R$-operation ($R^{\ast}$-operation) Chetyrkin:1982nn ; CHETYRKIN1984419 ; Larin:2002sc , as well as the modern Feynman diagram generating and Feynman integral reduction algorithms were refined and played the major roles.
-Early attempts were also made to push renormalization computations to higher-loop order, for instance the primitive divergences of the 7-loop beta function Broadhurst:1995km , or extend them to composite operators, for instance the 4-loop anomalous dimension of operators with derivatives Derkachov:1997pf .
-However, the complete 6-loop computation of the quartic scalar theory became feasible only approximately 15 years after the 5-loop result, firstly with the 6-loop field renormalization result Batkovich:2016jus , followed by the beta function and mass renormalization for the $O(1)$ model Kompaniets:2016hct , and subsequently for the $O(N)$ model Kompaniets:2017yct .
-A major breakthrough came with the development of the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which enabled the successful computation of the 7-loop beta function, mass renormalization, and field renormalization Schnetz:2016fhy , and even the 8-loop field renormalization function Schnetz:2025opm .
+In the development of the renormalization group (RG) method, scalar field theories have emerged as an ideal testing ground for validating multi-loop computation techniques while offering valuable insights into more intricate field theories, such as non-Abelian gauge theories.
+Nevertheless, the renormalization of scalar field theories is inherently significant in its own right.
+As elucidated in Wilson’s theory of critical phenomena Wilson:1971dc , critical exponents near second-order phase transitions in certain statistical systems are intimately linked to the anomalous dimensions of fields or composite operators PELISSETTO2002549 ; Moshe:2003xn .
+Two prominent examples of renormalizable scalar field theories are the six-dimensional cubic scalar field theory and the four-dimensional quartic scalar field theory.
+Both theories belong to the same $O(N)$ universality class, and a key conjecture posits that the $O(N)$ quartic scalar theory at its ultraviolet (UV) fixed point is equivalent to the $O(N)$ cubic scalar theory at its infrared (IR) fixed point Fei:2014yja ; Fei:2014xta .
+Namely, they approach to the same critical theory at their respective fixed points.
+Verifying this conjecture with high-loop precision for the scaling dimensions of fields or composite operators represents an intriguing problem, yet it requires perturbative computations of renormalization functions to higher-loop orders.
+Thanks to recent technical advancements, there has been a growing surge of interest in pushing renormalization computations to the next loop order.
+The renormalization computations of quartic and cubic scalar theories possess a long-standing history, tracing back to early-era studies of critical exponents up to $\epsilon^{2}$ orders for generalized Ising and Baxter models Wilson:1971dc , as well as the generalized Heisenberg model Wilson:1971vs .
+The 3-loop computations of the beta functions, field renormalization and mass renormalization for the quartic scalar theory were completed just one year later BREZIN1973227 ; Brezin:1973igb , with 4-loop results following five years thereafter Kazakov:1979ik .
+The full 5-loop renormalization of the quartic scalar theory, however, demanded over a decade of collaborative efforts from numerous researchers Chetyrkin:1981jq ; Gorishnii:1983gp ; Kazakov:1983dyk ; Gorishnii:1983jr ; Kleinert:1991rg ; Adzhemyan:2013jra .
+Throughout this development, techniques such as the IR-rearrangement Vladimirov:1979zm ; Chetyrkin:1980pr ; Caswell:1981ek , the $R$-operation ($R^{\ast}$-operation) Chetyrkin:1982nn ; CHETYRKIN1984419 ; Larin:2002sc , and modern Feynman diagram generating and Feynman integral reduction algorithms were refined and played major roles.
+Early attempts were also made to push renormalization computations to higher-loop orders, for instance the primitive divergences of the 7-loop beta function Broadhurst:1995km .
+Or extend them to composite operators, for instance the 4-loop anomalous dimension of operators with derivatives Derkachov:1997pf .
+Nevertheless, the complete 6-loop computation of the quartic scalar theory only became feasible approximately 15 years after the 5-loop result.
+Firstly with the 6-loop field renormalization result Batkovich:2016jus , followed by the beta function and mass renormalization for the $O(1)$ model Kompaniets:2016hct , and subsequently for the $O(N)$ model Kompaniets:2017yct .
+A major breakthrough emerged with the development of the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which enabled the successful computation of the 7-loop beta function, mass renormalization, and field renormalization Schnetz:2016fhy , as well as the 8-loop field renormalization function Schnetz:2025opm .
 The history of renormalization computations for the cubic scalar theory has closely paralleled that of the quartic scalar theory, which is not surprising since both have benefited from the same technical advancements.
-However, due to greater complexity, progress in the cubic scalar theory has consistently lagged by 1 or 2 loops compared to its quartic counterpart at similar stages.
+However, due to its greater complexity, progress in the cubic scalar theory has consistently lagged by 1 or 2 loops compared to its quartic counterpart at similar stages.
 Early 1-loop and 2-loop analysis of the cubic scalar theory revealed its asymptotic freedom Macfarlane:1974vp ; Ma:1975vn .
 Over the subsequent 40 years, 3-loop deAlcantaraBonfim:1980pe ; deAlcantaraBonfim:1981sy ; Fei:2014xta and 4-loop Gracey:2015tta renormalization computations were gradually completed, with the 5-loop result finished only very recently Kompaniets:2021hwg ; Borinsky:2021jdb .
-The frontier record once again came from the graphical function method, by which the 6-loop beta function, field renormalization and mass renormalization functions of the cubic scalar theory have been successfully acquired this very year Schnetz:2025wtu .
-Besides the computations of the beta functions and anomalous dimensions of fields and mass, the renormalization of composite operators constitutes an equal important yet more challenging problem.
-Phenomenologically, this is relevant to specific critical exponents in phase transition problems.
+Once again, the frontier record stems from the graphical function method, through which the 6-loop beta function, field renormalization and mass renormalization functions of the cubic scalar theory have been successfully acquired this year Schnetz:2025wtu .
+Besides the beta functions and the anomalous dimensions of fields and mass, the renormalization of composite operators constitutes an equally important yet more challenging problem.
+Phenomenologically, this is relevant to specific critical exponents in phase transition phenomena.
 One class of composite operators of interest is the fixed-charge symmetric traceless operator $\phi^{Q}$.
-In statistical physics, the anomalous dimensions of $\phi^{Q}$ for certain values of $Q$ are relevant to phase transitions in diverse systems PELISSETTO2002549 , such as liquid crystals, graphite-intercalation compounds, polymers, etc.
-DePrato:2003yd ; Calabrese:2004ca .
-Interest in computing the scaling dimensions of the $\phi^{Q}$ operator in cubic and quartic scalar theories grows from 2019, when the semi-classical method is introduced to compute the scaling dimension of $\phi^{Q}$ for generic $Q$ at the large $Q$ limit.
-This leads to all-loop results at leading and sub-leading orders in the $Q$ expansion for the quartic scalar theory Arias-Tamargo:2019xld ; Badel:2019oxl ; Antipin:2020abu ; Giombi:2020enj and the cubic scalar theory Arias-Tamargo:2020fow ; Antipin:2021jiw .
-Progress in the semi-classical approach stimulates the growing attention to the perturbative method approach towards renormalization computation of the $\phi^{Q}$ operator, aiming to validate the semi-classical results and also to obtain results for all orders of $Q$.
-For the quartic scalar theory, the scaling dimension of the $\phi^{Q}$ operator has been computed perturbatively at 4-loop Jack:2021ypd and 5-loop order Jin:2022nqq .
-While the 6-loop result is also available Bednyakov:2022guj , in which the data of $Q=1,\ldots,5$ was computed via perturbative Feynman diagram method and the data of $Q=6,7$ was imported from the semi-classical method.
-Concurrently, for the cubic scalar theory, perturbative computations have been carried out at 3-loop Jack:2021ziq and 4-loop order Huang:2024hsn .
-High-loop order renormalization computations would never be possible without the continuous evolution of both traditional and modern techniques.
-One of the important techniques to mention is the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which takes the advantage of computing scalar integrals in coordinate space.
-It pushes renormalization computations to 8-loops for the four-dimensional quartic scalar theory and 6-loops for the six-dimensional cubic scalar theory, and for the first time, with some tricks it also pushes the renormalization computations to 7-loops in the Gross-Neveu model involving spinning particles Huang:2025ree .
-The graphical function method is designed to evaluate scalar integrals with no more than four external points, which is certainly not the case for integrals of correlation functions involving $\phi^{Q}$ operators.
-To address this, we have proposed an Operator Product Expansion (OPE) method for renormalization computations by analyzing the UV divergence structures of correlation functions in the large momentum expansion limit Huang:2024hsn .
-This OPE method has proven highly useful and efficient for determining anomalous dimensions Huang:2024hsn , and this fact has also been shown in the condensate calculations Marino:2024uco ; Liu:2025bqq .
-In the OPE framework, the momenta of the two operators are taken to the infinite limit.
-This can be considered pictorially as two hard fields propagating in a soft field background Huang:2016bmv ; Huang:2022qnx , reducing the integral evaluation to a single-scale problem characterized by the large momentum parameter.
-By identifying the contributing terms in the OPE, the essential UV divergences required for renormalization computations are found to be determined by two-point hard sub-graphs.
-Concurrently, the Wilson coefficients in the OPE are proportional to a combination of $Z$-factors and the contributions of these two-point hard sub-graphs.
-By requiring the UV finiteness of the Wilson coefficients, the $Z$-factors can be determined once the UV divergences of the hard graphs are known.
-This approach thus reduces renormalization computations to the evaluation of two-point propagator-type integrals.
-Note that the generality of OPE across any quantum field theories enables diverse applications of this OPE method, to theories with scalars or particles with spins, and also to the situations involving operator mixing and operators with derivatives, with Lorentz indices, etc.
-In this work, we apply the OPE method developed in Huang:2024hsn to compute the 5-loop scaling dimension of the $\phi^{Q}$ operators in the six-dimensional $O(N)$ cubic scalar theory, thereby extending beyond the 4-loop result reported in that paper.
+In statistical physics, the anomalous dimensions of $\phi^{Q}$ for certain values of $Q$ are relevant to phase transitions in diverse systems PELISSETTO2002549 , such as liquid crystals, graphite-intercalation compounds and polymers DePrato:2003yd ; Calabrese:2004ca .
+Interest in computing the scaling dimensions of the $\phi^{Q}$ operator in cubic and quartic scalar theories has grown since 2019, when the semi-classical method is introduced to compute the scaling dimension of $\phi^{Q}$ at the large $Q$ limit.
+This advancement leads to all-loop results at leading and sub-leading orders in the $Q$ expansion for the quartic scalar theory Arias-Tamargo:2019xld ; Badel:2019oxl ; Antipin:2020abu ; Giombi:2020enj and the cubic scalar theory Arias-Tamargo:2020fow ; Antipin:2021jiw .
+Progress in the semi-classical approach has stimulated increasing attention to the perturbative method approach towards renormalization computation of the $\phi^{Q}$ operator, aiming to validate the semi-classical results and also to obtain results for complete $Q$ dependence.
+For the quartic scalar theory, the scaling dimension of the $\phi^{Q}$ operator has been computed perturbatively at 4-loop Jack:2021ypd and 5-loop orders Jin:2022nqq .
+While the 6-loop result is also available Bednyakov:2022guj , where the data of $Q=1,\ldots,5$ was computed via perturbative Feynman diagram method and the data of $Q=6,7$ was adopted from the semi-classical method.
+Concurrently, for the cubic scalar theory, perturbative computations have been carried out at 3-loop Jack:2021ziq and 4-loop orders Huang:2024hsn .
+High-loop order renormalization computations would never be feasible without the continuous evolution of both traditional and modern techniques.
+One notable approach is the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which takes the advantage of computing scalar integrals in coordinate space.
+To date, it has advanced renormalization computations to 8-loop order in the four-dimensional quartic scalar theory and 6-loop order in the six-dimensional cubic scalar theory.
+Furthermore, with some tricks, it has for the first time pushed renormalization computations to 7-loop order in the Gross-Neveu model Huang:2025ree .
+However, the graphical function method is inherently limited to scalar integrals with no more than four external points.
+When investigating the anomalous dimensions of higher-length composite operators, for instance the $\phi^{Q}$ operator, the number of external legs can be very large, rendering the graphical function method inapplicable directly.
+To address this challenge, a novel renormalization algorithm based on the Operator Product Expansion (OPE) has been proposed Huang:2024hsn .
+As established in the literature, the renormalization group equation (RGE) flow of Wilson coefficients is governed by the anomalous dimensions of operators, e.g., as described in Collins_1984 .
+This principle enables the extraction of anomalous dimensions from Wilson coefficients and has been occasionally employed to compute the anomalous dimensions of specific operators.
+For instance, the OPE of two scalar fields was used to determine the anomalous dimension of the $\phi^{2}$ operator Collins_1984 , and similarly, the OPE of two stress-energy tensors was applied to compute the anomalous dimension of the Konishi operator in $\mathcal{N}=4$ super-Yang-Mills theory Eden:2012fe .
+In practice, however, standard renormalization computations typically rely on traditional techniques such as the $R^{\ast}$-operation method and the massive vacuum bubble method.
+By considering the OPE of a composite operator with a fundamental field, the proposed algorithm Huang:2024hsn extends the applicability of the OPE framework to general composite operators in a systematic manner, making it suitable for implementation in modern computational tools.
+In this algorithm, the UV finiteness of the corresponding OPE coefficients links the anomalous dimensions of higher-dimensional composite operators to those of lower-dimensional operators, thereby establishing recursive relations for anomalous dimensions.
+Utilizing these recursion relations, the anomalous dimensions of complicated operators can be non-trivially derived from those of simple operators.
+Additionally, the UV divergences required for renormalization computation are treated globally without the need to subtract sub-divergences, and the latter is often computationally cumbersome in other renormalization techniques such as the $R^{\ast}$-operation method.
+The algorithm also provides a systematic approach to deform integrals of correlation functions involving fields and/or operators into two-point propagator-type integrals, which is particularly advantageous when dealing with systems with numerous external legs.
+We have explicitly validated the OPE based algorithm in the cubic scalar theory and Gross-Neveu-Yukawa theory Huang:2024hsn , confirming its highly versatile and efficient.
+The utility of determining anomalous dimensions via OPE has also been tested in condensate calculations Marino:2024uco ; Liu:2025bqq .
+Notably, the generality of OPE across quantum field theories allows the proposed algorithm to be applied to diverse scenarios, ranging from theories involving scalars or particles with spins to situations that include operator mixing, as well as operators carrying derivatives and Lorentz indices.
+In this study, we apply the OPE based algorithm proposed in Huang:2024hsn , combined with the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , to calculate the 5-loop scaling dimension of the $\phi^{Q}$ operators in the six-dimensional $O(N)$ cubic scalar theory, thereby extending beyond the 4-loop result.
 This paper is structured as follows.
-In §2, we provide a concise review of the cubic scalar theory, its renormalization functions, and the OPE method for renormalization computation.
-Treatments for organizing and simplifying Feynman graphs of the correlation functions involving the $\phi^{Q}$ operator are also presented.
-The §3 is devoted to a detailed demonstration of the OPE method using 2-loop examples, showing how the field renormalization function, beta function and the anomalous dimensions of the $\phi^{Q}$ operator can be computed efficiently.
-In §4, general discussion for the computation of 5-loop anomalous dimension of $\phi^{Q}$ operator has been presented, and various 5-loop results can be found in this section and the Appendix.
+In §2, we provide a concise review of the cubic scalar theory, its renormalization functions, and the aforementioned renormalization algorithm.
+Additionally, we present approaches for organizing and simplifying Feynman graphs of correlation functions involving the $\phi^{Q}$ operator.
+§3 is dedicated to a detailed demonstration of the OPE based algorithm via 2-loop examples, illustrating how the field renormalization function, beta function, and anomalous dimensions of the $\phi^{Q}$ operator can be computed efficiently.
+In §4, we present a general discussion on the computation of the 5-loop anomalous dimension of the $\phi^{Q}$ operator, with various 5-loop results provided in this section and the Appendix.
 In §5, conclusion and discussion are presented.
 2 Review of the theory and the method
-In this section we review the six-dimensional $O(N)$ cubic scalar theory, with emphasis on its renormalization computations and the techniques for organizing and simplifying Feynman diagrams.
-To make this paper self-contained, we also provide a brief review of the OPE method for renormalization computations Huang:2024hsn , especially for its application in the context of cubic scalar theory.
+In this section, we review the six-dimensional $O(N)$ cubic scalar theory, with a focus on its renormalization computations and the techniques for organizing and simplifying Feynman diagrams.
+To make this paper self-contained, we also present a brief overview of the OPE based renormalization algorithm Huang:2024hsn , particular in the context of its application to the cubic scalar theory.
 2.1 The cubic scalar theory and its renormalization
 The theory under consideration is the six-dimensional $O(N)$ cubic scalar theory, defined by the Lagrangian,
-$\mathcal{L}=\frac{1}{2}(\partial\phi_{i})^{2}+\frac{1}{2}(\partial\sigma)^{2}-\frac{g}{2!}\sigma\phi_{i}^{2}-\frac{h}{3!}\sigma^{3}~,$ (1)
-where $\sigma$ and $\phi_{i},i=1,\ldots,N$ denote two types of real scalar fields, with the $\phi_{i}$ fields transforming as an $O(N)$ vector.
-This theory is perturbatively renormalizable in six-dimensional space-time.
+$\mathcal{L}=\frac{1}{2}(\partial\phi_{i})^{2}+\frac{1}{2}(\partial\sigma)^{2}+\frac{g}{2!}\sigma\phi_{i}^{2}+\frac{h}{3!}\sigma^{3}~,$ (1)
+where $\sigma$ and $\phi_{i},i=1,\ldots,N$ denote two types of real scalar fields, with the $\phi_{i}$ transforming as an $O(N)$ vector.
+This theory is perturbatively renormalizable in six-dimensional spacetime.
 The primary focus of this paper is the renormalization analysis of the symmetric traceless $\phi^{Q}$ operator, defined as,
@@
 where $T_{i_{1}\cdots i_{Q}}$ is a symmetric traceless tensor and $Q$ is a fixed charge number.
-We will work in the $\overline{\mbox{MS}}$ regularization scheme tHooft:1973mfk in $d=(6-\epsilon)$-dimension.
-The classical scaling dimension of $\phi^{Q}$ is Q(d/21)Q(d/2-1), while the total scaling dimension is
-$\Delta_{Q}=Q\left(\frac{d}{2}-1\right)+\gamma_{Q}+Q\gamma_{\phi}~.$ (3)
-The difference between the total scaling dimension and the classical scaling dimension is defined as the anomalous dimension of the $\phi^{Q}$ operator.
-The gamma function $\gamma_{Q}$ of $\phi^{Q}$ operator can be computed perturbatively by Feynman diagrams of inserting the $\phi^{Q}$ operator in the 1PI Green function with $Q$ external legs.
-However, explicit evaluation of these correlation functions is complicated by the cumbersome $O(N)$ tensor structures inherent to the operator, which cause the number of terms in Feynman rules to grow exponentially with increasing $Q$.
-To address this problem, it is convenient to introduce an auxiliary $O(N)$ zero-norm complex vector $q$ satisfying $^{2}=0, qq$$\cdot\bar{q}=1$, and define the tensor $T$ as Antipin:2020abu ; Jin:2022nqq
+We work in the $\overline{\mbox{MS}}$ regularization scheme tHooft:1973mfk in $d=(6-\epsilon)$-dimension.
+The classical scaling dimension of $\phi^{Q}$ is Q(d/21)Q(d/2-1), while the total scaling dimension reads
+$\Delta_{Q}=Q\left(\frac{d}{2}-1\right)+\gamma_{\varphi^{Q}}~.$ (3)
+The difference between the total and classical scaling dimensions defines the anomalous dimension $\gamma_{\phi^{Q}}$ of the $\phi^{Q}$ operator.
+It can be computed perturbatively via Feynman diagrams involving the insertion of the $\phi^{Q}$ operator into 1PI Green functions with $Q$ external legs.
+However, explicit evaluation of these correlation functions is complicated by the cumbersome $O(N)$ tensor structures inherent to the operator, which cause the number of terms in the Feynman rules to grow exponentially with increasing $Q$.
+To address this challenge, it is convenient to introduce an auxiliary $O(N)$ zero-norm complex vector $q$ satisfying $^{2}=0, qq$$\cdot\bar{q}=1$, and define the tensor $T$ as Antipin:2020abu ; Jin:2022nqq
 $T_{i_{1}\ldots i_{Q}}:=\frac{1}{Q!}q_{i_{1}}q_{i_{2}}\cdots q_{i_{Q}}~.$ (4)
-This allows us to redefine the $\phi^{Q}$ operator in terms of a $\varphi^{Q}$ operator,
+This allows us to reexpress the $\phi^{Q}$ operator in terms of a $\varphi^{Q}$ operator,
 $\phi^{Q}:=\frac{1}{Q!}\varphi^{Q}~,$ (5)
-where the complex field $\varphi$ is a redefinition of the scalar field $\varphi:=q\cdot\phi$, with its complex conjugate field $\bar{\varphi}:=\bar{q}\cdot\phi$.
-This definition effectively strips $O(N)$ indices from the correlation functions, while the anomalous dimension of $\phi^{Q}$ operator is equivalent to that of $\varphi^{Q}$ operator.
+where the complex field $\varphi$ is defined as $\varphi:=q\cdot\phi$, with its complex conjugate $\bar{\varphi}:=\bar{q}\cdot\phi$.
+This redefinition effectively eliminates $O(N)$ indices from the correlation functions, and the anomalous dimension of $\phi^{Q}$ operator is equivalent to that of $\varphi^{Q}$ operator.
 The latter can now be computed by $\varphi^{Q}$-$Q\varphi$ correlation functions, which are free of $O(N)$ index structures.
-In the $\varphi$-field representation, the $\sigma\phi^{2}$ vertices along the line of the operator and the external legs are reformulated as $\sigma\varphi\bar{\varphi}$ vertices, while other $\sigma\phi^{2}$ vertices appearing in loop structures of $\varphi$ fields contribute factors of $N$ after summing over color indices.
-Due to charge conservation, scalar charge flows along $\varphi$ fields in the corresponding Feynman diagrams.
-For purpose of renormalization computation, following standard procedures, we introduce renormalized fields as,
+In the $\varphi$-field representation, the $\sigma\phi^{2}$ vertices along the operator and external legs are reformulated as $\sigma\varphi\bar{\varphi}$ vertices, while other $\sigma\phi^{2}$ vertices appearing in loop structures of $\varphi$ fields contribute factors of $N$ after summing over $O(N)$ indices.
+Owing to charge conservation, scalar charge flows along the $\varphi$ fields in the corresponding Feynman diagrams.
+For the purpose of renormalization computation, we follow standard textbook procedures and introduce renormalized fields as
 $\varphi_{0}=Z^{\frac{1}{2}}_{\varphi}\varphi~~~,~~~\sigma_{0}=Z^{\frac{1}{2}}_{\sigma}\sigma~,$ (6)
-and renormalized coupling constants as,
-$g_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{g}g~~~,~~~h_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{h}h~~~\mbox{with}~~~\widetilde{Z}_{g}=Z_{g}Z^{-\frac{1}{2}}_{\sigma}Z^{-1}_{\varphi}~~~,~~~\widetilde{Z}_{h}=Z_{h}Z^{-\frac{3}{2}}_{\sigma}~,$ (7)
-where $g_{0},h_{0}$ denote bare coupling constants and $\varphi_{0},\sigma_{0}$ denote bare fields.
-We also define the renormalized operator $\varphi^{Q}_{R}=Z_{\varphi^{Q}}\varphi_{0}^{Q}$.
-Since $\varphi_{0}^{Q}=Z_{\varphi}^{\frac{Q}{2}}\varphi^{Q}$, we get
+along with renormalized coupling constants as
+$g_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{g}g~~~,~~~h_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{h}h~~~\mbox{with}~~~\widetilde{Z}_{g}=Z_{g}Z^{-\frac{1}{2}}_{\sigma}Z^{-1}_{\varphi}~~~,~~~\widetilde{Z}_{h}=Z_{h}Z^{-\frac{3}{2}}_{\sigma}~.$ (7)
+Here, $g_{0},h_{0}$ denote bare coupling constants and $\varphi_{0},\sigma_{0}$ denote bare fields.
+We also define the renormalized operator $\varphi^{Q}_{R}=Z_{\varphi^{Q}}\varphi_{0}^{Q}$, where the anomalous dimension $\gamma_{\varphi^{Q}}$ is derived from $Z_{\varphi^{Q}}$.
+Since $\varphi_{0}^{Q}=Z_{\varphi}^{\frac{Q}{2}}\varphi^{Q}$, we obtain
 $\varphi_{R}^{Q}=\widetilde{Z}_{\varphi^{Q}}\varphi^{Q}~~~\mbox{with}~~~\widetilde{Z}_{\varphi^{Q}}=Z_{\varphi^{Q}}Z_{\varphi}^{\frac{Q}{2}}~.$ (8)
-Then $\widetilde{Z}_{\varphi^{Q}}$ can be assigned as the Feynman rule of the $\varphi_{R}^{Q}$-$Q\varphi$ vertex, and the anomalous dimension of the $\varphi^{Q}$ operator can be determined by the $Z_{\varphi^{Q}}$-factor.
-Provided that the $Z_{\varphi}$, $Z_{\sigma}$, $Z_{g}$ and $Z_{h}$ have been determined, we can compute beta functions via the standard field theory method.
-The beta functions $\beta(g)=\frac{\partial g}{\partial\ln\mu}$ and $\beta(h)=\frac{\partial h}{\partial\ln\mu}$ for the renormalized couplings can be derived from $\widetilde{Z}_{h},\widetilde{Z}_{g}$ by solving the following matrix,
+The factor $\widetilde{Z}_{\varphi^{Q}}$ can then be assigned as the Feynman rule for the $\varphi_{R}^{Q}$-$Q\varphi$ vertex.
+Once $Z_{\varphi}$, $Z_{\sigma}$, $Z_{g}$ and $Z_{h}$ are determined, beta functions can be computed using standard field theory methods.
+These beta functions follow from $\frac{\partial g}{\partial\ln\mu}$ and $\frac{\partial h}{\partial\ln\mu}$, which are derived by solving matrix equation,
 $\begin{pmatrix}\frac{1}{g}+\frac{\partial\ln\widetilde{Z}_{g}}{\partial g}&\frac{\partial\ln\widetilde{Z}_{g}}{\partial h}\\ \frac{\partial\ln\widetilde{Z}_{h}}{\partial g}&\frac{1}{h}+\frac{\partial\ln\widetilde{Z}_{h}}{\partial h}\\ \end{pmatrix}\begin{pmatrix}\frac{\partial g}{\partial\ln\mu}\\ \frac{\partial h}{\partial\ln\mu}\\ \end{pmatrix}=\begin{pmatrix}-\frac{\epsilon}{2}\\ -\frac{\epsilon}{2}\\ \end{pmatrix}~$.
 (9)
-Then the field anomalous dimensions of $\varphi$ and $\sigma$ can be computed by definition as,
+The field anomalous dimensions of $\varphi$ and $\sigma$ are then computed by definition as
 $\gamma_{\varphi}=\frac{1}{2}\left(\frac{\partial\ln Z_{\varphi}}{\partial g}\beta(g)+\frac{\partial\ln Z_{\varphi}}{\partial h}\beta(h)\right)~~~,~~~\gamma_{\sigma}=\frac{1}{2}\left(\frac{\partial\ln Z_{\sigma}}{\partial g}\beta(g)+\frac{\partial\ln Z_{\sigma}}{\partial h}\beta(h)\right)~.$ (10)
-Furthermore, once $Z_{\varphi^{Q}}$ has been determined, the anomalous dimension of the $\varphi^{Q}$ operator can be computed by
-$\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}=-\frac{\partial\ln Z_{\varphi^{Q}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}Z_{\varphi}^{-\frac{Q}{2}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}+Q\left(\frac{1}{2}\frac{\partial\ln Z_{\varphi}}{\partial\ln\mu}\right)~,$ (11)
-where the first term computes $\gamma_{Q}:=\gamma_{\varphi^{Q}}$ and the second term computes $Q\gamma_{\varphi}$.
-With the knowledge of beta functions and the anomalous dimensions, in order to study the large $N$ expansion of the anomalous dimensions around the Wilson-Fisher fixed point, one need to solve the fixed points from beta functions,
+Furthermore, once $Z_{\varphi^{Q}}$ is determined, the anomalous dimension of the $\varphi^{Q}$ operator is given by
+$\gamma_{\varphi^{Q}}:=-\frac{\partial\ln Z_{\varphi^{Q}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}Z_{\varphi}^{-\frac{Q}{2}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}+Q\left(\frac{1}{2}\frac{\partial\ln Z_{\varphi}}{\partial\ln\mu}\right)~.$ (11)
+We denote the first term as $\gamma_{Q}$, and the second term evaluates to $Q\gamma_{\varphi}$, such that $\gamma_{\varphi^{Q}}=\gamma_{Q}+Q\gamma_{\varphi}$.
+To study the large $N$ expansion of the anomalous dimensions around the Wilson-Fisher fixed point, we first solve for the fixed points of the beta functions,
 $\beta(g^{\ast})=0~~~,~~~\beta(h^{\ast})=0~.$ (12)
-After substituting the solutions $g^{\ast},h^{\ast}$ into the anomalous dimensions and expanding the results in the series of $1/N$, one produces the large $N$ expansion of the cubic scalar theory at the IR fixed point.
-The renormalization computation problems require the determination of the $Z$-factors. In general, a $Z$-factor of up to $L$-loop contributions takes the schematic form as
+Substituting the solutions $g^{\ast},h^{\ast}$ into the anomalous dimensions and expanding the results in a series of $1/N$ yields the large $N$ expansion of the cubic scalar theory at the IR fixed point.
+Renormalization computations require determining the $Z$-factors. In general, a $Z$-factor including up to $L$-loop contributions takes the schematic form
 $Z_{{\tiny\texttt{X}}}=1+\sum_{n_{\ell}=1}^{L}\sum_{n_{\epsilon}=1}^{n_{\ell}}\frac{z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}}{\epsilon^{n_{\epsilon}}}~,$ (13)
-in which the symbol X can represent the fields $\varphi,\sigma$, the couplings $h,g$ or the operator $\varphi^{Q}$ of arbitrary value $Q$.
-The $Z_{{\tiny\texttt{X}}}$-factor can be obtained by determining the coefficients $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ up to the appropriate loop orders.
-All such $Z$-factors can be conveniently computed using two-point propagator-type integrals, following the spirit of the OPE method to renormalization computations.
-We will review the OPE method in the next subsection.
-2.2 The OPE method for the renormalization
-In general quantum field theory, the product of two time-ordered operators admits an asymptotic expansion in the limit where their coordinates coincide, known as the Operator Product Expansion.
-The expansion basis is a set of composite operators of the theory, ordered according to their dimensions, while the expansion coefficients are functions of the coordinate difference, which are in general singular.
-When considering the OPE in momentum space by Fourier transformation with respect to the coordinate difference, the expansion is performed in the limit of large momentum $p\to\infty$, schematically expressed as,
+where X represents the fields $\varphi,\sigma$, the couplings $h,g$ or the operator $\varphi^{Q}$.
+The $Z_{{\tiny\texttt{X}}}$-factor is obtained by determining the coefficients $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ up to the relevant loop orders.
+All such $Z$-factors can be conveniently computed using two-point propagator-type integrals, following the spirit of the OPE based renormalization algorithm.
+2.2 The OPE based algorithm for renormalization
+The OPE is a classical technique for investigating physical quantities in the asymptotic expansion limit, and it is textbook knowledge that the anomalous dimension of an operator is linked to the Wilson coefficients of the OPE Collins_1984 .
+In principle, this connection enables the computation of anomalous dimensions directly from Wilson coefficients.
+In Huang:2024hsn , we proposed an algorithm that explicitly realize this connection by relating renormalization $Z$-factors to the UV finiteness of Wilson coefficients.
+This algorithm was validated in practical renormalization calculations for the cubic scalar theory and Gross-Neveu-Yukawa theory, confirming its versatility and efficiency.
+Unlike other renormalization computational techniques, such as the $R^{\ast}$-operation method, this algorithm treats UV divergences globally in renormalization procedures, eliminating the need to subtract sub-divergences arising from certain Feynman diagrams.
+It also provides a systematic approach to extracting the essential UV divergences from two-point integrals, applicable not only to beta functions and the anomalous dimensions of fundamental fields, but also to the anomalous dimensions of composite operators, including those involving correlation functions with a large number of external legs.
+To describe this algorithm, we begin with a general OPE.
+The product of two time-ordered operators is asymptotically expanded into a basis of composite operators, where the expansion basis is ordered by the operators’ dimensions.
+In momentum space, the expansion is performed in the large momentum limit $p\to\infty$, schematically expressed as
 $\widehat{O}_{\tiny\mbox{II}}^{\tiny\mbox{ren}}(k-p)\widehat{O}_{\tiny\mbox{I}}^{\tiny\mbox{ren}}(p)~\sim~C_{i}^{\tiny\mbox{ren}}(g_{i};p)\widehat{O}_{i}^{\tiny\mbox{ren}}(k)+\cdots$ (14)
-where $k$ is a finite momentum, and the momenta of the two operators tend to the opposite directions of infinite limits [^1] Pictorially, we can consider the two operators with large momentum as two hard fields propagating in the background of soft fields.
-This idea has also been realized in the study of the boundary operators of BCFW recursion relations Huang:2022qnx .
+where $k$ is a finite momentum, and the momenta of the two operators approach infinity in opposite directions [^1] Pictorially, we can regard the two large momentum operators as hard fields propagating in a background of soft fields.
+This idea has also been implemented in the study of boundary operators in BCFW recursion relations Huang:2016bmv ; Huang:2022qnx .
 By shifting two external momenta $p_{1},p_{2}$ with a complex auxiliary momentum $zq$ as $p_{1}\to p_{1}+zq$, $p_{2}\to p_{2}-zq$, at the large $z$ limit we get two hard fields propagating in the background of soft fields..
-The superscript $^{\tiny\mbox{ren}}$ in the operator $\widehat{O}$ represents a renormalized operator in dimensional regularization scheme.
-Alternatively, we may consider the OPE of bare operators, schematically written as,
+The superscript $^{\tiny\mbox{ren}}$ denotes a renormalized operator within the dimensional regularization scheme.
+Alternatively, we may consider the OPE of bare operators, schematically written as
 $\widehat{O}_{\tiny\mbox{II}}^{\tiny\mbox{bare}}(k-p)\widehat{O}_{\tiny\mbox{I}}^{\tiny\mbox{bare}}(p)~\sim~C_{i}^{\tiny\mbox{bare}}(g_{i0};p)\widehat{O}_{i}^{\tiny\mbox{bare}}(k)+\cdots$ (15)
@@
 $\widehat{O}^{{\tiny\mbox{ren}}}=Z_{\widehat{O}}\widehat{O}^{{\tiny\mbox{bare}}}~,$ (16)
-from eqn.(14) and eqn.(15) we get the relation,
+combing eqn.(14) and eqn.(15) yields the relation
 $C_{i}^{\tiny\mbox{ren}}(g_{i};p)=\frac{Z_{\widehat{O}_{\tiny\mbox{I}}}Z_{\widehat{O}_{\tiny\mbox{II}}}}{Z_{\widehat{O}_{i}}}C_{i}^{\tiny\mbox{bare}}(g_{i0};p)~.$ (17)
-Special attention should be paid to the OPE of fundamental fields, since the $Z$-factor of a fundamental field $\varphi$ or $\sigma$ does not correspond directly to $Z_{\varphi}$ or $Z_{\sigma}$.
-Instead, it is defined through $\varphi=Z_{\varphi}^{-\frac{1}{2}}\varphi_{0}$ or $\sigma=Z_{\sigma}^{-\frac{1}{2}}\sigma_{0}$.
-Furthermore, one usually computes the amputated correlation functions, and an extra factor $Z_{\varphi}$ or $Z_{\sigma}$ should be included for each fundamental field in the OPE of amputated correlation functions.
-With these two considerations, a fundamental field $\varphi$ or $\sigma$ contributes a $Z^{\frac{1}{2}}_{\varphi}$ or $Z^{\frac{1}{2}}_{\sigma}$ factor to the numerator of the Wilson coefficient relation (17) in the context of amputated OPE.
-However, the $Z$-factor in the denominator corresponds to an operator but not an external leg, so it still contributes to the $Z^{-\frac{1}{2}}_{\varphi}$ or $Z^{-\frac{1}{2}}_{\sigma}$ factor.
-Due to the fact that the Wilson coefficients of renormalized operators $C_{i}^{\tiny\mbox{ren}}(g_{i};p)$ are UV finite, once the Wilson coefficients of bare operators $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed, the ratios of $Z$-factors can be determined using UV finiteness conditions generated from the cancellation of $\epsilon$ poles.
-It could be possible that some operators in the basis have the same dimension, leading to the operator mixing situation.
-In this case, the relation (16) is generalized to the vectors of mixed operators related by matrices of $Z$-factors.
-The $Z$-matrices can be determined by the UV finiteness conditions from all the OPE coefficients corresponding to the mixed operators.
-Importantly, the above discussion is based on general operators, which means that it is applicable to either simple operators, or operators with Lorentz indices, with fermion and gauge fields, etc.
-The Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ of operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ can be computed by the large momentum expansion of bare correlation functions.
+This relation explicitly realize the connection between anomalous dimensions and Wilson coefficients.
+By carefully selecting the basis operator $\widehat{O}_{i}$, we can compute the desired $Z$-factors using simpler correlation functions.
+Special attention must be paid to the OPE of fundamental fields.
+For a fundamental field $\phi$, its associated $Z$-factor does not directly correspond to $Z_{\phi}$.
+Instead, the field renormalization is defined as $\phi=Z_{\phi}^{-\frac{1}{2}}\phi_{0}$.
+Furthermore, amputated correlation functions are typically computed, and an additional factor of $Z_{\phi}$ must be included for each fundamental field in the OPE of amputated correlation functions.
+Incorporating these two considerations, a fundamental field $\phi$ contributes a factor of $Z^{\frac{1}{2}}_{\phi}$ to the numerator of the Wilson coefficient relation (17) in the context of amputated OPEs.
+Since the Wilson coefficients of renormalized operators, $C_{i}^{\tiny\mbox{ren}}(g_{i};p)$, are UV finite, once the bare Wilson coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed, the ratios of $Z$-factors can be determined by enforcing UV finiteness conditions, specifically, the cancellation of $\epsilon$ poles.
+It is possible for basis operators to have identical dimensions, leading to operator mixing.
+In such cases, the relation (16) generalizes to vectors of mixed operators related by matrices of $Z$-factors.
+These $Z$-matrices are determined by imposing UV finiteness on all OPE coefficients corresponding to the mixed operators.
+Notably, the above discussion applies to generic scalar and tensor operators within the framework of general quantum field theories.
+The bare Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ for the bare operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ is computed via the large momentum expansion of bare correlation functions.
 For an operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ composed of $n$ fields, the corresponding Wilson coefficient is given by
@@
 where $k=-\sum_{i=1}^{n}k_{i}$.
-Pictorially, this corresponding to splitting the Feynman graph of the original correlation function into hard and soft sub-graphs by cutting $n$ internal propagators and/or external legs, with the Wilson coefficient receiving contributions from all possible $n$-cut hard sub-graphs.
-If the operator contains fields with derivatives, for instance $\partial_{\mu}\varphi_{0}(k)$, we can compute the coefficient by acting $i\frac{\partial}{\partial k^{\mu}}$ to the bare correlation function.
-For hard sub-graphs, since all the $k_{i},i=1,\ldots,n$ have been set to zero and there remains only two operators with hard momenta, we effectively get two-point graphs.
-Hence the Wilson coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed by the two-point propagator-type integrals.
-This greatly simplifies the computation, especially when dealing with correlation functions with large number of external legs.
-The above discussion outlines the core idea of computing renormalization functions via OPE method.
-To determine beta functions and anomalous dimensions, we first need the $Z$-factors, which are extracted from UV finiteness conditions generated from (17), i.e.,
+Pictorially, this corresponds to splitting the Feynman graph of the original correlation function into hard and soft sub-graphs by cutting $n$ internal propagators and/or external legs.
+The Wilson coefficient receives contributions from all possible $n$-cut hard sub-graphs.
+For operators involving derivative fields, for instance $\partial_{\mu}\varphi_{0}(k)$, the coefficient is computed by acting with $i\frac{\partial}{\partial k^{\mu}}$ on the bare correlation function.
+For hard sub-graphs, since all $k_{i},i=1,\ldots,n$ are set to zero, leaving only two operators with hard momenta, we effectively reduce the problem to two-point graphs.
+Thus, the bare Wilson coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are evaluated using two-point propagator-type integrals.
+This significantly simplifies computations, particularly when dealing with correlation functions with a large number of external legs.
+The above discussion outlines the core idea of the OPE based algorithm for computing renormalization functions.
+To determine beta functions and anomalous dimensions, we first extract $Z$-factors from UV finiteness conditions implied by eqn.(17),
 $\frac{Z_{\widehat{O}_{\tiny\mbox{I}}}Z_{\widehat{O}_{\tiny\mbox{II}}}}{Z_{\widehat{O}_{i}}}C_{i}^{\tiny\mbox{bare}}(g_{i0};p)=\mbox{UV~finite}~.$ (19)
-To establish these UV finiteness conditions, we first need to compute the coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ from the corresponding two-point integrals.
-Then using the general ansatz (13) for $Z$-factors, we can expand the complete expression as series expansion of $\epsilon$, and generate equations by setting coefficients of all $\epsilon$ poles to zero.
-The $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ coefficients of the $Z_{\tiny\texttt{X}}$-factor can be obtained by solving these equations, and consequently the $Z_{\tiny\texttt{X}}$-factor can be successfully determined to the desired loop order, provided the two-point integrals can be computed to that order.
-Now let us apply the OPE method to the renormalization computation of the cubic scalar theory, working in coordinate space.
-As in the standard field theory computation, the $Z_{\varphi}$ and $Z_{\sigma}$-factors are computed from two-point correlation functions.
-This computation can be naturally framed within the OPE method.
-Consider the OPE,
+To establish these UV finiteness conditions, firstly we compute the bare coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ from the corresponding two-point integrals.
+Then using the general ansatz (13) for $Z$-factors, we expand the full expression as a series in $\epsilon$, and generate equations by setting the coefficients of all $\epsilon$ poles to zero.
+Solving these equations yields the coefficients $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ of the $Z_{\tiny\texttt{X}}$-factor, thereby determining $Z_{\tiny\texttt{X}}$-factor to the desired loop order, provided the two-point integrals can be computed to that order.
+We now apply this algorithm to the cubic scalar theory of interest, working in coordinate space.
+As standard, the $Z_{\varphi}$ and $Z_{\sigma}$-factors are computed from two-point correlation functions.
+This computation naturally fits within this algorithmic framework.
+Consider the OPEs,
 $\varphi\bar{\varphi}~\sim~c_{\mathbf{1}}~\mathbf{1}+\cdots~~~,~~~\sigma\sigma~\sim~c^{\prime}_{\mathbf{1}}~\mathbf{1}+\cdots~,$ (20)
-where $\mathbf{1}$ is the unit operator of zero dimension, and $c$ generically represents Wilson coefficient, which is in general different between the two relations.
-Since they are the OPE of two fundamental fields, the large momentum expansion of them produces,
+where $\mathbf{1}$ denotes the zero-dimensional unit operator, and $c,c^{\prime}$ are distinct Wilson coefficients for the two expansions.
+For the OPE of two fundamental fields, the large momentum expansion yields
 $\langle\varphi(x_{1})\bar{\varphi}(x_{2})\rangle~\cong~Z_{\varphi}G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}))~~~,~~~\langle\sigma(x_{1})\sigma(x_{2})\rangle~\cong~Z_{\sigma}G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}))~,$ (21)
-where we use notation $G(g_{i0};\varphi(x_{1}),\varphi(x_{2}),\varphi(\mbox{\textst{$x$}}_{3}),\ldots)$ to represent the bare Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$.
-It is computed by the two-point integrals, derived from the bare correlation function with $x$ denoting the corresponding external leg removed, as in eqn.(18).
-Since they are already two-point functions, no additional legs need removal, so $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}))$ are computed directly from Feynman graphs of two-point correlation functions.
-Then from UV finiteness conditions of (21), the $Z$-factors of fields $\varphi$ and $\sigma$ can be determined.
-The $Z$-factors of coupling constants $g,h$ can be computed from the OPE,
+where we use the notation $G(g_{i0};\varphi(x_{1}),\varphi(x_{2}),\varphi(\mbox{\textst{$x$}}_{3}),\ldots)$ to represent the bare Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$.
+As defined in eqn.(18), this coefficient is computed via two-point integrals derived from bare correlation functions, with $x$ indicating the removal of the corresponding external leg.
+Since they are already two-point functions, no additional legs need ro be removal.
+Thus $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}))$ are evaluated directly from the Feynman graphs of the two-point correlation functions.
+The $Z$-factors $Z_{\varphi}$ and $Z_{\sigma}$ are then determined by imposing UV finiteness conditions of (21).
+The $Z$-factors for the coupling constants $g$ and $h$ are computed from the OPEs,
 $\varphi\bar{\varphi}~\sim~c_{\sigma}~\sigma+\cdots~~~,~~~\sigma\sigma~\sim~c^{\prime}_{\sigma}~\sigma+\cdots~.$ (22)
-In the large momentum expansion, it produces $Z$-factor ratios $Z^{\frac{1}{2}}_{\varphi}Z^{\frac{1}{2}}_{\varphi}/Z^{-\frac{1}{2}}_{\sigma}$ and $Z^{\frac{1}{2}}_{\sigma}Z^{\frac{1}{2}}_{\sigma}/Z^{-\frac{1}{2}}_{\sigma}$ respectively.
-So we obtain,
+In the large momentum limit, these expansions yield $Z$-factor ratios $Z^{\frac{1}{2}}_{\varphi}Z^{\frac{1}{2}}_{\varphi}/Z^{-\frac{1}{2}}_{\sigma}$ and $Z^{\frac{1}{2}}_{\sigma}Z^{\frac{1}{2}}_{\sigma}/Z^{-\frac{1}{2}}_{\sigma}$, respectively.
+Here, the soft $\sigma$ field is treated as an operator and dose not require amputation.
+This leads to
 $\langle\varphi(x_{1})\bar{\varphi}(x_{2})\sigma(x_{3})\rangle~\cong~Z_{\sigma}^{\frac{1}{2}}Z_{\varphi}G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~,$ (23)
 $\langle\sigma(x_{1})\sigma(x_{2})\sigma(x_{3})\rangle~\cong~Z^{\frac{3}{2}}_{\sigma}G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~.$ (24)
-Further using the relation (7) we get [^2] In the previous paper Huang:2024hsn , we have taken the convention that the $g_{0}$ or $h_{0}$ in the denominator is absorbed into the bare correlation function for three-point functions.
-This leads to slightly different expressions in that paper, where $G$ represents the bare correlation function divided by $g_{0}$ or $h_{0}$ respectively.
+Using the coupling renormalization relations (7), we further obtain [^2] In the previous paper Huang:2024hsn , we adopted the convention that $g_{0}$ or $h_{0}$ in the denominator is absorbed into the bare correlation function for three-point functions.
+This leads to slightly different expressions in that paper, where $G$ represents the bare correlation function divided by $g_{0}$ or $h_{0}$, respectively.
 $\langle\varphi(x_{1})\bar{\varphi}(x_{2})\sigma(x_{3})\rangle~\cong~\frac{Z_{g}g}{g_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~,$ (25)
 $\langle\sigma(x_{1})\sigma(x_{2})\sigma(x_{3})\rangle~\cong~\frac{Z_{h}h}{h_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~.$ (26)
-The $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ are directly computed from the two-point integrals of corresponding three-point functions by removing one leg of $\sigma$ field, and the $Z_{g},Z_{h}$ can be determined from UV finiteness conditions of (25), (26).
-With results of $Z_{\varphi},Z_{\sigma}$ and $Z_{g},Z_{h}$, the beta functions and anomalous dimensions of fields can be readily computed by definitions (9), (10).
-The $Z$-factor of the $\varphi^{Q}$ operator can be computed from the OPE,
-$\varphi^{Q}\bar{\varphi}~\sim~c_{i}~\varphi^{Q-1}~.$ (27)
-After large momentum expansion it produces the ratio of $Z$-factors as,
+The coefficients $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ are computed directly from two-point integrals of the corresponding three-point functions, with one $\sigma$ field leg removed.
+The factors $Z_{g}$ and $Z_{h}$ are then determined by enforcing UV finiteness on (25) and (26).
+With $Z_{\varphi},Z_{\sigma}$ and $Z_{g},Z_{h}$ in hand, the beta functions and field anomalous dimensions are readily computed via definitions (9) and (10).
+Finally, the $Z$-factor of the $\varphi^{Q}$ operator is computed from the OPE,
+$\varphi^{Q}\bar{\varphi}~\sim~c_{Q}~\varphi^{Q-1}~.$ (27)
+After large momentum expansion of this OPE yields the $Z$-factors ratio
 $\frac{Z_{\varphi^{Q}}Z^{\frac{1}{2}}_{\varphi}}{Z_{\varphi^{Q-1}}}=\frac{Z_{\varphi^{Q}}Z^{\frac{Q}{2}}_{\varphi}}{Z_{\varphi^{Q-1}}Z^{\frac{Q-1}{2}}_{\varphi}}=\frac{\widetilde{Z}_{\varphi^{Q}}}{\widetilde{Z}_{\varphi^{Q-1}}}~.$ (28)
-Thus, the $Z_{\varphi^{Q}}$-factor can be computed recursively using the UV finiteness conditions,
+Thus, $Z_{\varphi^{Q}}$ can be computed recursively by imposing UV finiteness condition,
 $\frac{\widetilde{Z}_{\varphi^{Q}}}{\widetilde{Z}_{\varphi^{Q-1}}}G_{Q}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})=\mbox{UV~finite}~,$ (29)
-where $G_{Q}$ is computed from Feynman graphs of the $\varphi^{Q}$-$Q\varphi$ correlation function by removing (Q1)(Q-1) external legs.
-They are two-point integrals with large momentum localized at the $\varphi^{Q}$ operator and one remaining external leg.
-Using the results of $Z_{\varphi^{Q}}$ and the previously computed beta functions, the anomalous dimension of the $\varphi^{Q}$ operator can be computed by definition (11).
+where $G_{Q}$ is computed from Feynman graphs of the $\varphi^{Q}$-$Q\varphi$ correlation function with (Q1)(Q-1) external legs removed.
+These integrals are two-point propagator-type integrals, with large momentum localized at the $\varphi^{Q}$ operator and one remaining external leg.
+Using the resulting $Z_{\varphi^{Q}}$ and the previously computed beta functions, the anomalous dimension of the $\varphi^{Q}$ operator os obtained via eqn.(11).
+As a side remark, the transformation of general integrals into two-point integrals is a common technique for evaluating UV divergences across various scenarios.
+Examples include the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , the IR-rearrangement technique Vladimirov:1979zm ; Chetyrkin:1980pr ; Caswell:1981ek , and works that do not rely on the OPE approach Fei:2014xta ; Jack:2021ziq .
+In all these cases, two-point integrals serve as a convenient computational tool for extracting UV divergences from Feynman integrals.
+However, with the aid of the OPE, two-point integrals are encoded as Wilson coefficients, from which anomalous dimensions can be systematically derived.
 2.3 From Feynman diagrams to two-point integrals
-The UV finiteness conditions discussed in the previous subsection require computing two-point integrals derived from two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions.
-In this subsection, we review some techniques which will help to simplify the problem of generating Feynman diagrams and transfer them to two-point graphs for subsequent evaluation of two-point integrals.
+The UV finiteness conditions discussed in the preceding subsection require computing two-point integrals derived from two-point, three-point, and $\varphi^{Q}$-$Q\varphi$ correlation functions.
+In this subsection, we review key techniques that facilitate simplifying the generation of Feynman diagrams and transforming them into two-point graphs for the subsequent evaluation of two-point integrals.
 The $\Phi$-field representation
-The cubic scalar theory we consider involves two types of fields $\varphi,\sigma$, and two types of cubic interactions.
-The complex field $\varphi$ carries a scalar charge that flows along the line of $\varphi,\bar{\varphi}$ fields, which can be represented by an arrow indicating the direction of charge flow.
-When generating Feynman diagrams, we need to account for all possible field configurations in the internal lines, which becomes a rather difficult task when the number of internal lines is large.
-It has also been observed that for a given external field configuration, different internal field configurations only contribute to varying combinations of couplings, while the underlying integrals remain identical.
-Thus there is no need to generate all Feynman diagrams and redundantly compute the same integrals.
-To clarify the separation between the structure of couplings and the integrals, we will adopt the $\Phi$-field representation.
-In this framework, we combine the $\varphi_{i},i=1,\ldots,N$ fields and the $\sigma$ field into a single $(N+1)$ component scalar field as,
+The cubic scalar theory under consideration involves two types of fields $\varphi,\sigma$ and two classes of cubic interactions.
+The complex field $\varphi$ carries a scalar charge that flows along $\varphi$ and $\bar{\varphi}$ lines.
+This charge flow can be represented by an arrow indicating its direction.
+Generating Feynman diagrams requires accounting for all possible internal field configurations, a task that becomes increasing cumbersome as the number of internal lines grows.
+Notably, for a fixed external field configuration, different internal field configurations only affect the combinations of couplings, while the underlying scalar integrals remain identical.
+Thus, there is no need to generate all Feynman diagrams explicitly or redundantly compute identical integrals.
+To decouple the coupling structure from the integral calculations, we adopt the $\Phi$-field representation.
+In this framework, we unify the $\varphi_{i},i=1,\ldots,N$ fields and the $\sigma$ field into a single $(N+1)$ component scalar field
 $\Phi_{I}=(\Phi_{0},\Phi_{1},\ldots,\Phi_{N}):=(\sigma,\varphi_{1},\ldots,\varphi_{N})~,$ (30)
-and the cubic interactions can be universally encoded as
+with the cubic interactions universally encoded as
 $\frac{1}{6}\lambda_{IJK}\Phi_{I}\Phi_{J}\Phi_{K}~~~\mbox{where}~~~\lambda_{000}=h~~\mbox{and}~~\lambda_{0ij}=g\delta_{ij}~~\forall i,j\neq 0~,$ (31)
-with all other components of the symmetric coupling $\lambda_{IJK}$ vanishing.
-In the $\Phi$-field representation, we essentially generate the topologies of Feynman diagrams without specifying the explicit configuration of component fields.
-In this case, we have separated the information of the scalar integral part from the factor of couplings.
-The Feynman scalar integral can be derived from a given topology by assigning propagators according to Feynman rules, while the factors of couplings of all Feynman diagrams sharing the same topology are encoded in the combination of $\lambda_{IJK}$ couplings, which we define as the coupling factor.
+and all other components of the symmetric coupling tensor $\lambda_{IJK}$ vanish.
+In the $\Phi$-field representation, we generate Feynman diagram topologies without specifying the explicit component fields of $\Phi_{I}$.
+This effectively separates the scalar integral information from the coupling-dependent factors.
+The Feynman scalar integral for a given topology is derived by assigning propagators according to standard Feynman rules, while the coupling factors for all diagrams sharing that topology are encapsulated in combinations of the $\lambda_{IJK}$ couplings.
+We refer to this combined term as the coupling factor.
 The symmetry factor and the coupling factor
-When writing the Feynman integral of correlation functions, we first generate all corresponding topologies with private Mathematica code.
-For each topology, we get the scalar integral from the combination of propagators, multiplied by the standard symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ from the total symmetry of the graph and external legs.
-Each topology effectively represents a collection of numerous Feynman diagrams that share the same external state configuration but differ in their internal state configurations.
-By summing over the indices of the $\lambda_{IJK}$, we get the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of the contributing Feynman graphs.
-To clarify the concept of the coupling factor, let us take the two-point two-loop topologies $\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ depicted in Fig.(2) as an example.
+When constructing the Feynman integrals for correlation functions, we first generate all corresponding topologies using custom Mathematica code.
+For each topology, the scalar integral is derived from the product of propagators, multiplied by the standard symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$, accounting for the total symmetry of the graph and external legs.
+Each topology effectively represents a set of numerous Feynman diagrams that share the same external state configuration but differ in their internal state configurations.
+By summing over the indices of the $\lambda_{IJK}$, we obtain the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ for the contributing Feynman graphs.
+To illustrate the concept of the coupling factor, we consider the two-point two-loop topology shown below,
+For this topology, we analyze two external field configurations: (1) two external $\sigma$ field represented by dashed lines, (2) two external $\varphi_{i}$ and $\bar{\varphi}_{j}$ fields represented by arrowed lines, as depicted in the LHS of Fig.(1).
+Each vertex is assigned a $\lambda_{IJK}$ tensor, and summing over the internal indices yields the respective coupling factors
+$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\sigma\sigma})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{0I_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}0}=Ng^{4}+2Ng^{3}h+h^{4}~,$ (32)
+$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\varphi_{i}\bar{\varphi}_{j}})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{iI_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}j}=(2g^{3}h+2g^{4})\delta_{ij}~.$ (33)
+Alternatively, we can draw all contributing Feynman diagrams with their internal field configurations, as shown in the RHS of Fig.(1).
+It can be verified that summing over the contributions of all four Feynman diagrams for each external field configuration reproduces the coupling factors derived above.
+Since generating topologies is significantly more efficient than generating full Feynman diagrams, and evaluating coupling factors involves only algebraic operations, the $\Phi$-field representation substantially enhances computational efficiency.
 $Ng^{4}$$Ng^{3}h$$Ng^{3}h$$h^{4}$$g^{3}h$$g^{3}h$$g^{4}$$g^{4}$
@@
 The internal $\Phi$ fields represent all possible internal state configurations in component fields.
-For this topology, we can consider the external field configuration of two external $\sigma$ field represented by the dashed lines, or the field configuration of two external $\varphi,\bar{\varphi}$ fields represented by the arrowed lines, as depicted in the LHS of Fig.(1).
-Each vertex is assigned a $\lambda_{IJK}$, and after summing over the internal indices, we get the respective coupling factors as,
-$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\sigma\sigma})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{0I_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}0}=Ng^{4}+2Ng^{3}h+h^{4}~,$ (32)
-$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\varphi\varphi})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{iI_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}i}=2g^{3}h+2g^{4}~.$ (33)
-Alternatively, we can draw all contributing Feynman diagrams with the explicit internal field configurations, as shown in the RHS of Fig.(1).
-It can be verified that summing over all four Feynman diagrams of each external field configuration indeed produces the coupling factors evaluated above.
-Since generating topologies is significantly more efficient than generating complete Feynman diagrams, and the evaluation of coupling factors involves only algebraic operations, the $\Phi$-field representation substantially enhances the computational capability.
 The Feynman graphs of $\varphi^{Q}$-$Q\varphi$ correlation functions
-In the UV finiteness conditions, the function $G$ is computed schematically as
+In the UV finiteness conditions, the function $G$ is schematically computed as
 $G=\sum\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{{\tiny\mbox{topo}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{{\tiny\mbox{topo}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{{\tiny\mbox{topo}}})~,$ (34)
-where the summation runs over all possible topologies, and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{{\tiny\mbox{topo}}})$ is computed from the two-point integrals generated by these topologies.
-Remark that the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of a topology depends on the explicit external field configuration, but the symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}$ are irrelevant to the external fields, they are computed from the topologies.
+where the summation runs over all possible topologies, and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{{\tiny\mbox{topo}}})$ is evaluated from the two-point integrals generated by these topologies.
+It is worth noting that the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of a topology depends on the explicit external field configuration, whereas the symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ and UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}$ are topology-dependent quantities that are independent of external fields.
 For two-point correlation functions, $G$ is directly evaluated using all 1PI two-point topologies.
-For three-point correlation functions, we start from all 1PI three-point topologies and compute the independent two-point graphs by removing one external leg from these topologies in all possible ways.
-For the $\varphi^{Q}$-$Q\varphi$ correlation function, we start from all 1PI $\varphi^{Q}\to Q\varphi$ topologies, and $G_{Q}$ can be computed from all independent two-point graphs generated by removing (Q1)(Q-1) external legs from these topologies in all possible ways.
-By investigating the structure of 1PI $\varphi^{Q}\to Q\varphi$ topologies, it has been proven that only a part of these topologies contributes to the two-point integrals Huang:2024hsn .
-These contributing topologies are called irreducible graphs, characterized by the property that when the operator vertex is deleted, all remaining external legs stay connected in a single graph (except those that reduce to a single line).
-The other topologies are called reducible graphs, with the property that when deleting the operator vertex causes the external legs to distribute into two or more separate connected components.
-Hence we only need to generate the irreducible graphs of the 1PI $\varphi^{Q}\to Q\varphi$ topologies, and compute $G_{Q}$ from their corresponding two-point graphs after all possible removals of (Q1)(Q-1) legs.
-The summation of these graphs gives,
+For three-point correlation functions, we start from all 1PI three-point topologies and construct independent two-point graphs by removing one external leg from these topologies in every possible manner.
+For the $\varphi^{Q}$-$Q\varphi$ correlation function, we begin with all 1PI $\varphi^{Q}\to Q\varphi$ topologies.
+$G_{Q}$ is then computed from all independent two-point graphs generated by removing (Q1)(Q-1) external legs from these topologies in all possible ways.
+By analyzing the structure of 1PI $\varphi^{Q}\to Q\varphi$ topologies, it has been proven that only a subset of these topologies contributes to the two-point integrals Huang:2024hsn .
+These contributing topologies are termed irreducible graphs, defined by the property that deleting the operator vertex leaves all remaining external legs connected within a single graph (with the exception of configurations that reduce to a single line).
+The remaining topologies are classified as reducible graphs, which exhibit the property that deleting the operator vertex partitions the external legs into two or more disjoint connected components.
+Thus, we only need to generate irreducible graphs from the set of 1PI $\varphi^{Q}\to Q\varphi$ topologies, and compute $G_{Q}$ from their corresponding two-point graphs after all possible removals of (Q1)(Q-1) legs.
+The summation of these graphs yields
 $G_{Q}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})=\mathbb{U}_{\footnotesize\mbox{v}}\left(\frac{1}{x_{Q}^{2\lambda}}\right)+\sum_{n=2}^{Q}C_{Q-1}^{n-1}\mathtt{F}_{n}(\mbox{\textst{$x$}}_{\sigma_{1}},\cdots,\mbox{\textst{$x$}}_{\sigma_{n-1}},x_{Q})~.$ (35)
-The first term comes from the tree-level contribution and evaluates to 1, while
+The first term corresponds to the tree-level contribution and evaluates to 1, while
 $\mathtt{F}_{n}(\mbox{\textst{$x$}}_{\sigma_{1}},\cdots,\mbox{\textst{$x$}}_{\sigma_{n-1}},x_{Q})=(n-1)!\sum_{n_{\ell}=n-1}^{L}G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})~.$ (36)
-The $G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}$ receives contributions from all $n_{\ell}$-loop irreducible topologies.
-As in the standard procedure, it can be evaluated using the symmetry factor, coupling factor and UV divergence as,
+$G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}$ receives contributions from all $n_{\ell}$-loop irreducible topologies.
+Following standard procedures, it can be evaluated using the symmetry factor, coupling factor and UV divergence as
 $G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})=\sum_{{\tiny\mbox{irreducible}}\atop{\tiny\mbox{topologies}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{n_{\ell}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\cdots\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{n_{\ell}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\cdots\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{n_{\ell}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\cdots\mbox{\textst{{\scriptsize$\varphi$}}}})~,$ (37)
-where the summation is over all non-isomorphic graphs generated by removing (Q1)(Q-1) external legs of the irreducible 1PI topologies of the $\varphi^{Q}$-$Q\varphi$ correlation function.
-Let us summarize this section.
-After generating the independent topologies of $\varphi^{Q}$-$Q\varphi$ correlation functions, one computes the symmetry factor and coupling factor of each topology, and the UV divergences of two-point integrals in the dimensional regularization scheme to an appropriate $\epsilon$ order.
-Then the $G$ and $G_{Q}$ can be readily computed by summing over all two-point graphs.
-The $Z$-factors could be determined from UV finiteness conditions, and eventually the anomalous dimensions can be computed from these $Z$-factors.
+where the summation is performed over all non-isomorphic graphs generated by removing (Q1)(Q-1) external legs from the irreducible 1PI topologies of the $\varphi^{Q}$-$Q\varphi$ correlation function.
+Let us summarize this section as follows.
+After generating the independent topologies of $\varphi^{Q}$-$Q\varphi$ correlation functions, we compute the symmetry factor and coupling factor of each topology, along with the UV divergences of two-point integrals in the dimensional regularization scheme up to the desired $\epsilon$ order.
+The quantities $G$ and $G_{Q}$ can then be readily obtained by summing over all two-point graphs.
+The $Z$-factors are determined from the UV finiteness conditions, and the anomalous dimensions are ultimately computed from these $Z$-factors.
 3 Two-loop demonstration
-For illustrative purposes, in this section we will present detailed computations of the two-loop anomalous dimensions of the $\varphi^{Q}$ operator.
-This requires computing the field renormalization $Z_{\varphi}$ and $Z_{\sigma}$-factors, the coupling renormalization $Z_{h}$ and $Z_{g}$-factors and the $\varphi^{Q}$ operator renormalization $Z_{\varphi^{Q}}$-factor.
-These factors are determined from the corresponding two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions respectively.
-All results related to Feynman diagrams are computed by private Mathematica code, while the UV divergences of two-point integrals are computed by HyperlogProcedures Maple package HP .
-The strategy of computation is formalized as a sequential algorithm, following the computations of the symmetry factors of topologies, the coupling factors of Feynman diagrams, the UV divergence of two-point integrals and finally the determination of $Z$-factors from UV finiteness conditions.
+For illustrative purposes, this section presents detailed computations of the two-loop anomalous dimensions of the $\varphi^{Q}$ operator.
+Beta functions and anomalous dimensions are derived from $Z$-factors following standard quantum field theory definitions, where the $Z$-factors themselves are determined via the OPE based algorithm.
+This algorithm provides a non-trivial implementation of the connection between OPE Wilson coefficients and anomalous dimensions, enabling the systematic determination of $Z$-factors through summation over two-point integrals.
+The HyperlogProcedures package HP developed from graphical function method assures the computation of two-point integral to very high loop order.
+The computational work-flow is formalized as a sequential algorithm, proceeding through four key steps: (1) computing the symmetry factors of topologies, (2) evaluating the coupling factors of Feynman diagrams, (3) determining the UV divergences of two-point integrals, and (4) extracting $Z$-factors from UV finiteness conditions.
 3.1 Two-point correlation functions and the field renormalization
-Let us begin with the field renormalization computation.
-The $Z$-factors $Z_{\varphi}$ and $Z_{\sigma}$ can be computed from the two-point correlation functions $\langle\varphi\bar{\varphi}\rangle$ and $\langle\sigma\sigma\rangle$ respectively as in eqn.(21).
-The one-loop and two-loop topologies in $\Phi$-field representation are sketched in Fig.(2).
+We begin with the computation of field renormalization.
+As specified in eqn.(21), the $Z$-factors $Z_{\varphi}$ and $Z_{\sigma}$ are derived from the two-point correlation functions $\langle\varphi\bar{\varphi}\rangle$ and $\langle\sigma\sigma\rangle$, respectively.
+Notably, while the derivations coincide with those of traditional renormalization methods, we frame it explicitly within the OPE-based algorithm for consistency.
+The one-loop and two-loop topologies in the $\Phi$-field representation are illustrate in Fig.(2).
 $\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$
 Figure 2: The 1-loop and 2-loop 1PI topologies of two-point functions in $\Phi$-field representation.
-Each topology represents a collection of diagrams with distinct allowable internal field configurations of $\sigma,\varphi_{i}$, and these collections are characterized by coupling factors.
-The symmetry factors and the coupling factors of these topologies are listed in the table below,
+Each topology represents a set of diagrams with distinct valid internal configurations of the $\sigma$ and $\varphi_{i}$ fields, where these sets are characterized by their respective coupling factors.
+The symmetry factors and coupling factors for these topologies are listed in the table below,
 topology $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma_{\Phi\Phi})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\sigma\sigma})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\varphi\bar{\varphi}})$
-$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi}$ 1/2 $h_{0}^{2}+Ng_{0}^{2}$ $2g_{0}^{2}$
-$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi}$ 1/2 $N(2g_{0}^{4}+g_{0}^{2}h_{0}^{2})+h_{0}^{4}$ $Ng_{0}^{4}+2g_{0}^{4}+g_{0}^{2}h_{0}^{2}$
-$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ 1/2 $N(g_{0}^{4}+2g_{0}^{3}h_{0})+h_{0}^{4}$ $2g_{0}^{4}+2g_{0}^{3}h_{0}$
-There is no need to cut any external legs of these topologies, and their corresponding UV divergences can be directly computed by TwoPoint command in HyperlogProcedures.
-Feynman integrals in the $\overline{\mbox{MS}}$ dimensional regularization scheme can be evaluated as power series of parameter $\epsilon$ with the results(an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral has not been explicitly shown),
+$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi}$ 1/2 $h^{2}+Ng^{2}$ $2g^{2}$
+$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi}$ 1/2 $N(2g^{4}+g^{2}h^{2})+h^{4}$ $Ng^{4}+2g^{4}+g^{2}h^{2}$
+$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ 1/2 $N(g^{4}+2g^{3}h)+h^{4}$ $2g^{4}+2g^{3}h$
+No external legs need to be cut from these topologies, and their UV divergences are computed directly using the TwoPoint command in the HyperlogProcedures package.
+Feynman integrals in the $\overline{\mbox{MS}}$ dimensional regularization scheme are evaluated as power series in $\epsilon$, with results presented below (an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral is implicitly included and not explicitly shown),
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi})=-\frac{1}{3\epsilon}-\frac{4}{9}+\left(\frac{\pi^{2}}{72}-\frac{13}{27}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (38)
@@
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi})=-\frac{1}{3\epsilon^{2}}-\frac{1}{\epsilon}+\left(\frac{\pi^{2}}{36}-\frac{15}{8}\right)+\left(-\frac{7\zeta_{3}}{36}+\frac{\pi^{2}}{12}-\frac{769}{288}\right)\epsilon+\mathcal{O}(\epsilon^{2})~.$ (40)
-Then the bare correlation functions are readily obtained as,
+The bare correlation functions are then obtained as
 $G(g_{0},h_{0};\Phi,\Phi)=1-\sum_{L={\tiny\mbox{L1}},\atop{\tiny\mbox{L2-I}},{\tiny\mbox{L2-II}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{L}_{\Phi\Phi})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{L}_{\Phi\Phi})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{L}_{\Phi\Phi})~,$ (41)
-where the field $\Phi$ takes the corresponding component $\sigma$ or $\varphi$ in $\mathbb{C}_{\footnotesize\mbox{f}}$ factor.
-Attention should be paid to the $\pm$ signs of individual terms in the summation.
-The bare correlation function is computed by summing contributions from all tree-level and loop Feynman diagrams.
-Note that the TwoPoint command only computes the scalar integrals, omitting all $i$ factors from, for instance, the vertices or propagators.
-When summing these contributions, we must recover the factor $i$ for each loop order.
-Each vertex carries a factor of i-i.
-Each propagator contains a factor of $i$ in its numerator, but after Wick rotation to Euclidean space, the denominator acquires an additional factor of 1-1, resulting in an overall factor of i-i for each propagator.
-Each loop integration gains a factor of $i$ from the Wick rotation $\ell_{0}\to i\ell_{E0}$.
-Supposing a loop graph has $n_{v}$ vertices, $n_{e}$ propagators, $n_{\ell}$ loops and $n_{p}$ external legs, then the total factor would be $(-i)^{n_{v}}(-i)^{n_{e}}i^{n_{\ell}}$.
-From graph theory, we have the relations $n_{v}+n_{\ell}-n_{e}=1$ and $3n_{v}=2n_{e}+n_{p}$, and after calculation we get the final factor $(-1)^{n_{p}}i$ for each loop order.
-For two-point functions, $n_{p}=2$, so each loop order has a factor $i$, whereas the tree-level contrition has a factor i-i.
-If we factor out an overall i-i, the two-point correlation function can be evaluated as the tree-level contribution minus the sum of loop contributions.
-The renormalized two-point correlation functions eqn.(21) should be UV finite, so we get the UV finiteness conditions
-$Z_{\varphi}G(g_{0},h_{0};\varphi(x_{0}),\bar{\varphi}(x_{1}))=\mbox{UV~finite}~~~,~~~Z_{\sigma}G(g_{0},h_{0};\sigma(x_{0}),\sigma(x_{1}))=\mbox{UV~finite}~,$ (42)
-which means they should be free of $\epsilon$ poles.
-In order to get the equations of UV finiteness conditions, we rewrite the bare couplings $g_{0},h_{0}$ in the bare correlation functions to $g,h$ by relations eqn.(7), then expand the full expression at the limit $g,h\to 0$ followed by the limit $\epsilon\to 0$.
-Note that in the rewriting of couplings with eqn.(7), each $g$ or $h$ is accompanied by a factor $\mu^{\frac{\epsilon}{2}}$, so from the coupling factors we know that at $L$-loop order there will be a $\mu^{L\epsilon}$ factor.
-Combined with the $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor, at $L$-loop order there will be an overall $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor.
-Expansion of the factor around $\epsilon\to 0$ produces powers of $\ln\frac{\mu^{2}}{p^{2}}$ terms.
-Since each $\ln\frac{\mu^{2}}{p^{2}}$ order is independent, the UV finiteness conditions also require the coefficient of each $\ln\frac{\mu^{2}}{p^{2}}$ order is free of $\epsilon$ poles.
-Hence the higher Log order terms will produce additional equations.
-Inspecting the UV divergence structures of $L$-loop integrals, the leading divergence is $1/\epsilon^{L}$}, thus at the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ order there will in general be $L$ equations for the cancellation of $L$ poles.
-At the $\ln^{k}\frac{\mu^{2}}{p^{2}}$ order, the leading divergence is $1/\epsilon^{L-k}$, so there will be (Lk)(L-k) equations.
-We can either use the equations from the leading Log terms [^3] The leading Log term denotes terms with $\ln^{0}(\mu^{2}/p^{2})$ dependence, and the sub-Log term denotes terms with $\ln(\mu^{2}/p^{2})$ dependence, etc., or include those equations from higher Log terms.
-If we only use equations of the leading Log terms, then all equations of $Z_{\varphi},Z_{\sigma},Z_{g},Z_{h}$ should be solved simultaneously to completely determine the $Z$-factors.
-However, if we also use the equations of sub-Log terms, then each $Z$-factor can be determined by solving its own equations of UV finiteness conditions.
-This will be demonstrated in the following computations.
-Now let us take a look at the UV finiteness conditions. For $Z_{\varphi}$-factor, at one-loop, after the above prescribed expansions, we get
+where $\Phi$ corresponds to either $\sigma$ or $\varphi$ in the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$.
+Care must be taken with the sign of each term in the summation.
+Bare correlation functions include contributions from all tree-level and loop Feynman diagrams, but the TwoPoint command only computes the scalar integrals, omitting $i$-factors from vertices, propagators, or other possible sources.
+These $i$-factors must be recovered using Feynman rules derived for the Euclidean-space Lagrangian.
+Graphical analysis shows the overall factor is $(-1)^{n_{p}}i$, where $n_{p}$ is the number of external legs.
+For two-point integrals ($n_{p}=2$), each loop order carries a factor $i$, while the tree-level contrition has a factor i-i.
+Factoring out an overall i-i, the two-point correlation function is expressed as the tree-level contribution minus the sum of loop contributions.
+The renormalized two-point correlation functions eqn.(21) must be UV finite, leading to the UV finiteness conditions
+$Z_{\varphi}G(g_{0},h_{0};\varphi(x_{0}),\bar{\varphi}(x_{1}))=\mbox{UV~finite}~~~,~~~Z_{\sigma}G(g_{0},h_{0};\sigma(x_{0}),\sigma(x_{1}))=\mbox{UV~finite}~.$ (42)
+To derive these conditions explicitly, firstly we rewrite the bare couplings $g_{0},h_{0}$ in the bare correlation functions in terms of renormalized couplings $g,h$ using eqn.(7), then expand the full expression in the double limit $g,h\to 0$, followed by $\epsilon\to 0$.
+Note that eqn.(7) implies each $g$ or $h$ is accompanied by a factor $\mu^{\frac{\epsilon}{2}}$, so $L$-loop contributions carry a $\mu^{L\epsilon}$ factor.
+Combined with the implicit $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor, $L$-loop terms include an overall $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor.
+Expanding this factor around $\epsilon\to 0$ generates terms proportional to powers of $\ln\frac{\mu^{2}}{p^{2}}$.
+Since each power of $\ln\frac{\mu^{2}}{p^{2}}$ is independent, UV finiteness requires the coefficient of every $\ln\frac{\mu^{2}}{p^{2}}$ term to be free of $\epsilon$ poles, introducing additional equations beyond those from the leading terms.
+Inspecting the UV divergence structures of $L$-loop integrals, the leading divergence is $1/\epsilon^{L}$}, so the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ order generally yields $L$ equations for canceling $L$ poles.
+For $\ln^{k}\frac{\mu^{2}}{p^{2}}$ terms, the leading divergence is $1/\epsilon^{L-k}$, producing (Lk)(L-k) equations.
+We may either use only equations from leading Log terms [^3] Leading Log terms refer to those with $\ln^{0}(\mu^{2}/p^{2})$ dependence, sub-Log terms to $\ln(\mu^{2}/p^{2})$ dependence, and so on., or include those from higher Log orders.
+If only leading Log equations are used, $Z_{\varphi},Z_{\sigma},Z_{g},Z_{h}$ must be solved simultaneously to fully determine the $Z$-factors.
+In contrast, incorporating sub-Log equations allows each $Z$-factor to be determined independently via it own UV finiteness conditions.
+We now analyze the UV finiteness conditions for the $Z_{\varphi}$-factor. At one-loop order, following the expansions outlined above, we obtain
 $\left(\frac{1}{\epsilon}\left(z_{\phi}^{11}+\frac{g^{2}}{3}\right)+\left(1+\frac{7g^{2}}{9}\right)+\mathcal{O}(\epsilon)\right)+\left(\frac{g^{2}}{3}+\frac{7g^{2}}{9}\epsilon+\mathcal{O}(\epsilon^{2})\right)\ln\frac{\mu^{2}}{p^{2}}+\cdots$ (43)
-The coefficients of sub-Log terms are already UV finite, and they will not provide any additional equations.
-In order for the leading Log term to be finite, the $1/\epsilon$ pole should cancel, and we get
+The coefficients of the sub-Log terms are already UV finite and do not yield additional equations. For the leading Log term to be finite, the $1/\epsilon$ pole must cancel, leading to the condition,
 $z_{\varphi}^{11}=-\frac{g^{2}}{3}~.$ (44)
-At two-loop order, we need to solve $z^{21}_{\varphi},z^{22}_{\varphi}$. Again after expansions, the vanishing of $1/\epsilon$ and $1/\epsilon^{2}$ poles at leading Log term produces two equations,
+At two-loop order, we solve for the coefficients $z^{21}_{\varphi},z^{22}_{\varphi}$.
+Expanding the expression, the cancelation of $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term gives two equations,
 $0=z_{\varphi}^{21}+z_{\varphi}^{11}+\frac{7}{3}\left(\frac{2}{3}g^{2}z_{g}^{11}-\frac{1}{3}g^{2}z_{\varphi}^{11}-\frac{1}{3}g^{2}z_{\sigma}^{11}\right)+\left(\frac{g^{2}}{3}-\frac{67g^{4}N}{432}+\frac{293g^{4}}{216}+\frac{5g^{3}h}{3}-\frac{67g^{2}h^{2}}{432}\right)~,$
 $0=z_{\varphi}^{22}+\left(\frac{2}{3}g^{2}z_{g}^{11}-\frac{1}{3}g^{2}z_{\varphi}^{11}-\frac{1}{3}g^{2}z_{\sigma}^{11}\right)+\left(\frac{5g^{4}}{18}+\frac{g^{3}h}{3}-\frac{g^{4}N}{36}-\frac{g^{2}h^{2}}{36}\right)~,$
-and the vanishing of $1/\epsilon$ pole at sub-Log term produces one equation,
+while the cancelation of the $1/\epsilon$ pole in the sub-Log term provides a third equation,
 $0=\left(\frac{2}{3}g^{2}z_{g}^{11}-\frac{1}{3}g^{2}z_{\varphi}^{11}-\frac{1}{3}g^{2}z_{\sigma}^{11}\right)+\left(\frac{5g^{4}}{9}+\frac{2g^{3}h}{3}-\frac{g^{4}N}{18}-\frac{g^{2}h^{2}}{18}\right)~.$
-In the traditional approach, the renormalization computation should be performed order by order.
-This is similar to the situation that solving the $Z$-factors using only equations of leading Log terms.
-For instance, by using two equations of leading Log term shown above, the determination of $z_{\varphi}^{21},z_{\varphi}^{22}$ will also rely on the results of all one-loop $Z$-factors, including $Z_{g},Z_{h},Z_{\varphi},Z_{\sigma}$, which must be determined a priori.
-Alternatively, if considering the UV finiteness conditions of all Log terms of a correlation function, the generated equations are sufficient to completely solve the corresponding $Z$-factor.
-This means that the computation does not rely on other $Z$-factors of lower-loop orders.
-For instance, the equation from sub-Log term shown above can be solved to obtain a combination of one-loop $Z$-factors.
-This one-loop $Z$-factor combination is exactly the one appearing in the equations of leading Log term, and after substituting it back to those two equations, we can completely determine the two-loop $Z_{\varphi}$-factor.
-Furthermore, as a by-product, the equations of sub-Log terms provide constraints to other $Z$-factors at lower-loop orders, which can be served as consistent checks for lower-loop $Z$-factors computed by other correlation functions.
-Consequently, we determine the two-loop $Z_{\varphi}$-factors as,
+Computations can be performed iteratively by loop order, solving for $Z$-factors using only leading Log term equations.
+For instance, using the two leading Log equations above, determining $z_{\varphi}^{21}$ and $z_{\varphi}^{22}$ requires prior knowledge of one-loop $Z$-factors, including $Z_{g},Z_{h},Z_{\varphi},Z_{\sigma}$.
+Alternatively, incorporating UV finiteness conditions from all Log terms of a correlation function generates a sufficient set of equations to solve for the corresponding $Z$-factor independently, without relying on lower-loop $Z$-factors from other channels.
+For instance, the sub-Log term equation can be solved to obtain a combination of one-loop $Z$-factors.
+This combination directly appears in the leading Log equations.
+Substituting this result back into the two leading Log equations fully determines the two-loop $Z_{\varphi}$-factor.
+As a byproduct, sub-Log term equations also provide consistency checks for lower-loop $Z$-factors computed from other correlation functions.
+Following this approach, we determine the two-loop $Z_{\varphi}$-factor coefficients as
 $z_{\varphi}^{21}=\frac{11g^{4}N}{432}-\frac{13g^{4}}{216}-\frac{g^{3}h}{9}+\frac{11g^{2}h^{2}}{432}~~~,~~~z_{\varphi}^{22}=-\frac{g^{4}N}{36}+\frac{5g^{4}}{18}+\frac{g^{3}h}{3}-\frac{g^{2}h^{2}}{36}~.$ (45)
-The determination of $Z_{\sigma}$-factor follows exactly the same procedures, and the only difference is the coupling factors.
-After the same computations, eventually we determine the two-loop $Z_{\sigma}$-factor as
+The $Z_{\sigma}$-factor is determined via an identical procedure, with the only difference being the coupling factors. After analogous computations, the two-loop $Z_{\sigma}$-factor coefficients are
 $z_{\sigma}^{11}=-\frac{g^{2}N}{6}-\frac{h^{2}}{6}~,$ (46)
 $z_{\sigma}^{21}=\frac{11g^{2}h^{2}N}{432}-\frac{g^{4}N}{216}-\frac{g^{3}hN}{9}-\frac{13h^{4}}{432}~~~,~~~z_{\sigma}^{22}=\frac{g^{4}N}{9}+\frac{g^{3}hN}{3}-\frac{g^{2}h^{2}N}{36}+\frac{5h^{4}}{36}~.$ (47)
-The higher-loop $Z_{\varphi},Z_{\sigma}$-factors can be determined following the same procedures.
-However there would be much more Feynman graphs, and the evaluation of two-point integrals would become considerably more difficult.
+Higher-loop $Z_{\varphi}$ and $Z_{\sigma}$-factor can be determined using the same iterative procedure.
+However, this becomes increasingly computationally demanding due to the exponential growth in the number of Feynman diagrams and the heightened complexity of evaluating two-point integrals at higher loops.
 3.2 Three-point correlation functions and the beta functions
-The determination of $Z_{h},Z_{g}$-factors is quite similar as that of $Z_{\varphi},Z_{\sigma}$-factors.
-We start from OPE and arrive at the UV finiteness conditions as eqn.(25), (26).
-Once the large momentum expansion of the bare correlation functions are computed, we can determine $Z_{g}$ and $Z_{h}$ from the UV finiteness conditions.
-Remind that the UV divergence is computed by all possible three-point topologies with one leg removed.
-Following the computation strategy, the one-loop and two-loop topologies in $\Phi$-field representation are sketched in Fig.(3), as well as their independent cut graphs.
 $\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$
@@
 They are generated by the non-isomorphic computation of all possible removals of one external leg.
-The $Z_{g}$-factor is determined by $\langle\varphi\bar{\varphi}\sigma\rangle$ with a $\sigma$ field removed, and the $Z_{h}$-factor is determined by $\langle\sigma\sigma\sigma\rangle$ with a $\sigma$ field removed.
-The symmetry factors and coupling factors of cut graphs are listed in Table.(1).
 topology $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\sigma\sigma\mbox{\textst{{\scriptsize$\Phi$}}}})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\varphi\bar{\varphi}\mbox{\textst{{\scriptsize$\Phi$}}}})$
@@
 Table 1: The symmetry factors and coupling factors of cut graphs for 2-loop three-point topologies.
-The UV divergence is computed by TwoPoint command for two-point integral of cut graphs, with the results [^4] Again an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral has not been explicitly shown,
+We now apply the proposed algorithm to determine the $Z_{h}$ and $Z_{g}$-factors.
+Starting from the corresponding OPEs, we derive the UV finiteness conditions given in eqn.(25) and (26).
+These equations non-trivially relate the respective $Z$-factor ratios to the sum of two-point integrals, which are generated by removing one external leg from three-point topologies in all possible ways.
+Following the algorithmic strategy, the one-loop and two-loop topologies in the $\Phi$-field representation, along with their independent cut graphs, are illustrated in Fig.(3).
+The $Z_{g}$-factor is determined from $\langle\varphi\bar{\varphi}\sigma\rangle$ with one $\sigma$ field removed, while the $Z_{h}$-factor is determined from $\langle\sigma\sigma\sigma\rangle$ with one $\sigma$ field removed.
+The symmetry factors and coupling factors of the cut graphs are listed in Table.(1).
+[^4] We note that the symmetry factor defined herein is given by (n1)!(n-1)!
+divided the internal symmetry of the diagram and the permutation symmetry of external legs, where $n$ is the number of external legs.
+This differs from the conventional definition, which only accounts for internal symmetry, as used in Fei:2014xta .
+For example, using data from Table.(1), the conventional symmetry factors for two-loop topologies are $S(\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}}=\frac{1}{2}$, $S(\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}}=1$, $S(\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=\frac{1}{2}$.
+The ratios of our symmetry factors to the conventional ones are $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})/S(\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=2$ and $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})/S(\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\tiny$\Phi$}}}})=2$.
+These differences can be absorbed into the coupling factors if the conventional symmetry factor definition is adopted, effectively encoding the summation of external legs into structure factors.
+UV divergences for the two-point integrals of the cut graphs are computed using TwoPoint command in HyperlogProcedures package.
+The one-loop results [^5] As before, an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral is implicit included and not explicitly shown.
+are
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{\epsilon}+3+\left(\frac{23}{4}-\frac{\pi^{2}}{16}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (48)
+while the two-loop results are
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{67}{72\epsilon}-\left(\frac{2671}{864}-\frac{5\pi^{2}}{288}\right)-\left(\frac{80911}{10368}-\frac{335\pi^{2}}{3456}-\frac{5\zeta_{3}}{144}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (49)
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{67}{72\epsilon}-\left(\frac{2743}{864}-\frac{5\pi^{2}}{288}\right)-\left(\frac{86167}{10368}-\frac{335\pi^{2}}{3456}-\frac{5\zeta_{3}}{144}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (50)
-and
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{21}{8\epsilon}+\left(\frac{275}{32}-\frac{5\pi^{2}}{96}\right)+\left(\frac{2785}{128}-\frac{35\pi^{2}}{128}-\frac{5\zeta_{3}}{48}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (51)
@@
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{2\epsilon}+\left(\frac{83}{24}-\zeta_{3}\right)+\left(\frac{3983}{288}-\frac{5\pi^{2}}{96}-\frac{16\zeta_{3}}{3}-\frac{\pi^{4}}{120}\right)\epsilon+\mathcal{O}(\epsilon^{2})~.$ (53)
-Special attention should be paid to the tree-level contribution, which is evaluated by the two-point integral of a three-point vertex with one leg removed.
-Evaluated in the coordinate space, the result is not 1, but
+Special care is required for the tree-level contribution, which is evaluated from the two-point integral of a three-point vertex with one leg removed.
+In coordinate space, this result is not unity but,
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L0}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=1+\epsilon+\left(\frac{3}{4}-\frac{\pi^{2}}{48}\right)\epsilon^{2}+\left(\frac{1}{2}-\frac{\pi^{2}}{48}-\frac{\zeta_{3}}{24}\right)\epsilon^{3}+\left(\frac{5}{16}-\frac{\pi^{2}}{64}-\frac{\zeta_{3}}{24}+\frac{\pi^{4}}{23040}\right)\epsilon^{4}+\mathcal{O}(\epsilon^{5})~.$ (54)
-Appropriate $\epsilon$ order terms should be included for the $L$-loop renormalization computation.
-Then the bare correlation functions for two-loop computation are readily obtained as [^5] For three-point function, $n_{p}=3$.
-As discussed in the two-point case, each loop order has a factor i-i, while at the tree-level the factor is also i-i.
-If factoring out an overall factor i-i, the three-point correlation function can be evaluated by tree contribution plus all loop contributions.,
+For $L$-loop renormalization computations, the appropriate orders of $\epsilon$ must be included.
+The bare correlation functions for two-loop computations are then obtained as [^6] For three-point function, $n_{p}=3$, so the overall $i$-factors for tree and loop contributions have the same sign.
 $G(g_{0},h_{0};\Phi,\Phi,\mbox{\textst{$\Phi$}})=g_{\Phi}(1+\epsilon)+\sum_{L={\tiny\mbox{L1}},{\tiny\mbox{L2-I}}_{1},{\tiny\mbox{L2-I}}_{2},\atop{\tiny\mbox{L2-II}}_{1},{\tiny\mbox{L2-II}}_{2},{\tiny\mbox{L2-III}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{L}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{L}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{L}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})~,$ (55)
-where the field $\Phi$ takes the corresponding component $\sigma$ or $\varphi$ in $\mathbb{C}_{\footnotesize\mbox{f}}$ factors, and $g_{\Phi}=g$ for $\langle\varphi\bar{\varphi}\sigma\rangle$ while $g_{\Phi}=h$ for $\langle\sigma\sigma\sigma\rangle$.
-The UV finiteness conditions of eqn.(25) and eqn.(26) can be realized by,
+where $\Phi$ corresponds to $\sigma$ or $\varphi$ in the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$, and $g_{\Phi}=g$ for $\langle\varphi\bar{\varphi}\sigma\rangle$ while $g_{\Phi}=h$ for $\langle\sigma\sigma\sigma\rangle$.
+The UV finiteness conditions of eqn.(25) and (26) are explicitly,
 $\frac{Z_{g}g}{g_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\varphi,\bar{\varphi},\mbox{\textst{$\sigma$}})=\mbox{UV finite}~~,~~\frac{Z_{h}h}{h_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\sigma,\sigma,\mbox{\textst{$\sigma$}})=\mbox{UV~finite}~.$ (56)
-After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ by relations eqn.(7), then expanding the full expression at the limit $g,h\to 0$ followed by the limit $\epsilon\to 0$, we get equations for $Z$-factors by the vanishing of $\epsilon$ poles in each $\ln\frac{\mu^{2}}{p^{2}}$ term [^6] The bare coupling in the denominator will introduce a $\mu^{\frac{\epsilon}{2}}$, which cancels the $\mu^{\frac{\epsilon}{2}}$ factor in numerator.
-In the bare correlation function, at $L$-loop order there will be a $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor from integration, and a $\mu^{\frac{\epsilon}{2}(2L+1)}$ factor from the couplings.
-If we take out an overall factor $\mu^{\frac{\epsilon}{2}}$ which will not affect the $\epsilon$ expansion, then each $L$-loop is assigned with a $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor contributing to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion at $\epsilon\to 0$ limit..
-For $Z_{g}$-factor, at one-loop there is only one equation from the vanishing of $1/\epsilon$ pole of the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term, which yields,
+After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ using relations eqn.(7), and expanding the full expression in the double limit $g,h\to 0$, followed by $\epsilon\to 0$, we obtain equations for $Z$-factors by enforcing the cancelation of $\epsilon$ poles in each $\ln\frac{\mu^{2}}{p^{2}}$ term [^7] The bare coupling in the denominator introduces a $\mu^{\frac{\epsilon}{2}}$, which cancels the $\mu^{\frac{\epsilon}{2}}$ factor in the numerator.
+In the bare correlation function, $L$-loop contribution include a $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor from integration and a $\mu^{\frac{\epsilon}{2}(2L+1)}$ factor from couplings.
+Factoring our an overall $\mu^{\frac{\epsilon}{2}}$ (which does not affect the $\epsilon$ expansion), each $L$-loop term carries a $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor that contributes to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion as $\epsilon\to 0$..
+For the $Z_{g}$-factor, the one-loop equation yields,
 $g^{3}+g^{2}h+gz_{g}^{11}=0~~~\to~~~z_{g}^{11}=-g^{2}-gh~.$ (57)
-At two-loop order, there will be two equations from the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term,
+At two-loop order, the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term provides two equations,
 $0=z_{g}^{22}+z_{g}^{21}+(9g^{2}+6gh+1)z_{g}^{11}+3g\Big(hz_{h}^{11}-(2g+h)z_{\phi}^{11}-(g+2h)z_{\sigma}^{11}\Big)$
@@
 $~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(\frac{5g^{4}N}{12}+\frac{7g^{4}}{6}-\frac{g^{3}hN}{6}+\frac{7g^{3}h}{3}+\frac{11g^{2}h^{2}}{12}+\frac{gh^{3}}{3}\right)~,$ (59)
-and one equation from the sub-Log term,
+while the sub-Log term provides a third equation,
 $0=(3g+2h)z_{g}^{11}+\Big(hz_{h}^{11}-(2g+h)z_{\phi}^{11}-(g+2h)z_{\sigma}^{11}\Big)$
 $~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(\frac{5g^{3}N}{6}+\frac{7g^{3}}{3}-\frac{g^{2}hN}{3}+\frac{14g^{2}h}{3}+\frac{11gh^{2}}{6}+\frac{2h^{3}}{3}\right)~.$ (60)
-Again we confirm that the one-loop $Z$-factors can be solved as a whole by the equation of sub-Log term. Consequently, from the equations of leading Log terms we get
+As before, the sub-Log term equation allows solving for the combination of one-loop $Z$-factors.
+Substituting this result into the leading Log equations fully determines the two-loop $Z_{g}$-factor coefficients,
 $z_{g}^{21}=\frac{7g^{3}hN}{72}-\frac{11g^{4}N}{144}-\frac{49g^{4}}{72}-\frac{19g^{3}h}{36}-\frac{101g^{2}h^{2}}{144}-\frac{gh^{3}}{36}~~,~~z_{g}^{22}=\frac{5g^{4}N}{12}+\frac{7g^{4}}{6}-\frac{g^{3}hN}{6}+\frac{7g^{3}h}{3}+\frac{11g^{2}h^{2}}{12}+\frac{gh^{3}}{3}~.$
-Similarly, solving the equations of UV finiteness conditions for $Z_{h}$-factors, we determine the two-loop $Z_{h}$-factor. We get $z_{h}^{11}=-(g^{3}N+h^{3})/h$, and
+Following analogous steps for the $Z_{h}$-factors, we solve its UV finiteness conditions to obtain the two-loop results.
+The one-loop coefficient is $z_{h}^{11}=-(g^{3}N+h^{3})/h$, and the two-loop coefficients are
 $z_{h}^{21}=\frac{7g^{2}h^{2}N}{48}-\frac{g^{5}N}{12h}-\frac{9g^{4}N}{8}-\frac{3g^{3}hN}{8}-\frac{23h^{4}}{48}~~,~~z_{h}^{22}=\frac{g^{5}N}{h}+\frac{3g^{4}N}{2}+\frac{3g^{3}hN}{2}-\frac{g^{2}h^{2}N}{4}+\frac{5h^{4}}{4}~.$
-The higher-loop $Z_{h},Z_{g}$-factors can be similarly determined by including more graphs and more orders of $\epsilon$ series of two-point integrals.
-With knowledge of $Z_{h},Z_{g},Z_{\phi}$ and $Z_{\sigma}$, we can compute beta functions by solving equations (9). After expanding the solutions at the limits $g,h\to 0$, we get two-loop results as,
+Higher-loop $Z_{h}$ and $Z_{g}$-factors can be determined similarly by including additional topologies and higher-order $\epsilon$ terms in the two-point integrals.
+With $Z_{h},Z_{g},Z_{\phi}$ and $Z_{\sigma}$ determined, the beta functions are computed by solving eqn.(9). Expanding the solutions in the limit $g,h\to 0$, the two-loop results are
 $\beta(g)$ $=$ $-\frac{\epsilon}{2}g+\left(\frac{g^{3}N}{12}-\frac{2g^{3}}{3}-g^{2}h+\frac{gh^{2}}{12}\right)$ (61)
@@
 $\beta(h)$ $=$ $-\frac{\epsilon}{2}h+\left(\frac{g^{2}hN}{4}-g^{3}N-\frac{3h^{3}}{4}\right)+\left(\frac{31g^{2}h^{3}N}{144}-\frac{g^{5}N}{6}-\frac{161g^{4}hN}{72}-\frac{5g^{3}h^{2}N}{12}-\frac{125h^{5}}{144}\right)~.$~~ (62)
-The 2-loop anomalous dimension of field $\varphi$ or $\sigma$ is then given by definition (10) as,
+The two-loop anomalous dimension of the $\varphi$ and $\sigma$ fields are then obtained via eqn.(10),
 $\gamma_{\varphi}=\frac{g^{2}}{6}+\frac{13g^{4}}{216}+\frac{g^{3}h}{9}-\frac{11g^{4}N}{432}-\frac{11g^{2}h^{2}}{432}~~~,~~~\gamma_{\sigma}=\frac{h^{2}}{12}+\frac{g^{2}N}{12}+\frac{g^{4}N}{216}+\frac{g^{3}hN}{9}-\frac{11g^{2}h^{2}N}{432}+\frac{13h^{4}}{432}~.$ (63)
 3.3 Anomalous dimensions of the $\varphi^{2}$ and $\varphi^{3}$ operators
-The $\widetilde{Z}_{\varphi^{Q}}$-factors for the $\varphi^{Q}$ operator are determined by the UV finiteness conditions (29), and the corresponding two-point integrals are evaluated by (35).
-When $Q=2$, all contributing topologies are irreducible topologies.
-The tree-level, 1-loop and 2-loop topologies of $\varphi^{2}$-$2\varphi$ correlation function have been sketched in Fig.(4), together with their corresponding cut graphs.
+The $\widetilde{Z}_{\varphi^{Q}}$-factors for the $\varphi^{Q}$ operator are non-trivially related to two-point integrals via UV finiteness conditions in eqn.(29), where the two-point integrals are generated by removing (Q1)(Q-1) external legs from $\varphi^{Q}\to Q\varphi$ correlation function graphs.
+These integrals are evaluated using eqn.(35).
+For $Q=2$, all contributing topologies are irreducible.
+The tree-level, one-loop and two-loop topologies of the $\varphi^{2}$-$2\varphi$ correlation function, along with their corresponding cut graphs, are illustrated in Fig.(4).
 $\Gamma^{\tiny\mbox{L0}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L0}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-III}_{1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}i}$$\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{2}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-III}_{2}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$
@@
 The UV divergences are computed from these cut graphs.
-Note that due to the symmetry of graphs, a topology could generate more than one non-isomorphic cut graphs, similar as we have mentioned in the three-point function computation.
-For the cut graphs related to the same irreducible topology, the symmetry factors as well as the coupling factors are the same as those of the topology.
-Now Let us compute the $Z$-factor of $\varphi^{2}$, from eqn.(35) we get the explicit expression for $\varphi^{2}$ as
+Due to graph symmetry, a single topology may generate multiple non-isomorphic cut graphs (as observed in three-point function computations), and cut graphs derived from the same irreducible topology share identical symmetry and coupling factors.
+For $Q=2$, eqn.(35) yields the explicit expression for $\varphi^{2}$,
 $G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})=1+\sum_{j=1}^{L}G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})~.$ (64)
-There is only one 1-loop irreducible topology $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi{\varphi}}$, which generates one cut graph $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}}$.
-With results,
-$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g^{2}_{0}~~~,~~~\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{\epsilon}+2+\epsilon\left(\frac{5}{2}-\frac{\pi^{2}}{24}\right)+\mathcal{O}(\epsilon^{2})~,$ (65)
-we get the required UV contribution of all 1-loop irreducible topologies for $\varphi^{2}\to 2\varphi$ as
+There is only one 1-loop irreducible topology $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi{\varphi}}$, which generates the cut graph $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}}$ with the following inputs,
+$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g^{2}_{0}~~~,~~~\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{\epsilon}+2+\epsilon\left(\frac{5}{2}-\frac{\pi^{2}}{24}\right)+\mathcal{O}(\epsilon^{2})~.$ (65)
+The one-loop irreducible UV contribution for $\varphi^{2}\to 2\varphi$ is
 $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{1-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})=\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})~.$ (66)
-Note that in (65), an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ with $L=1$ is not explicitly shown in the result of UV divergence, while there will also be a factor $\mu^{\epsilon L}$ from the couplings $\mathbb{C}_{\footnotesize\mbox{f}}$ at $L$-loop.
-Hence for (66), if recovering the $(1/p^{2})^{\frac{\epsilon}{2}}$ factor and turning $g_{0},h_{0}\to g,h$, there will be an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{\frac{\epsilon}{2}}$ factor, which contributes to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion.
-Generally, for $G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}$ there will be an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{j\frac{\epsilon}{2}}$ factor.
-For presentation purpose, the factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ will always be treated as implicit in the expressions of UV divergences.
-The 1-loop $\widetilde{Z}_{\varphi^{2}}$ can be determined from the UV finiteness of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, where we have used the fact that the 1PI contribution of $\widetilde{Z}_{\varphi^{1}}=1$.
-After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ by relations eqn.(7), then expanding the full expression at the limits $g,h\to 0$ followed by the limit $\epsilon\to 0$, we get the condition for the vanishing of $1/\epsilon$ pole in the leading Log term as,
-$0=z_{\varphi^{2}}^{11}+g^{2}~~~\to~~~z_{\varphi^{2}}^{11}=-g^{2}~,$ (67)
-and the sub-Log term provides no more conditions.
-Similarly, in order to determine the 2-loop $\widetilde{Z}_{\varphi^{2}}$, the contribution of $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ should be included.
-There are 5 independent irreducible topologies, which generates 7 cut graphs.
-The symmetry factors and coupling factors can be simply computed as,
+Note that the $L=1$ loop order factor $(1/p^{2})^{\frac{\epsilon}{2}}$ is implicitly included in the UV divergence result, and $L$-loop coupling factors introduce a $\mu^{\epsilon L}$ factor.
+Converting bare couplings $g_{0},h_{0}\to g,h$ introduces an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{\frac{\epsilon}{2}}$ factor, which contributes to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion.
+Generally, $G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}$ carries an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{j\frac{\epsilon}{2}}$ factor, and we always treat the $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor implicit in all UV divergence expressions.
+The one-loop $\widetilde{Z}_{\varphi^{2}}$ is determined by enforcing UV finiteness of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, where we have used the fact that the 1PI contribution of $\widetilde{Z}_{\varphi^{1}}=1$.
+Rewriting bare couplings via eqn.(7) and expanding in the limit $g,h\to 0$ followed by $\epsilon\to 0$, the $1/\epsilon$ pole cancelation in the leading Log term gives,
+$0=z_{\varphi^{2}}^{11}+g^{2}~~~\to~~~z_{\varphi^{2}}^{11}=-g^{2}~.$ (67)
+No additional conditions arise from sub-Log terms.
+For the two-loop $\widetilde{Z}_{\varphi^{2}}$-factor, $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ is included.
+There are 5 independent irreducible topologies generating 7 cut graphs, with symmetry factors
 $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~,$ (68)
-and
+and coupling factors
 $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}N+g_{0}^{2}h_{0}^{2}~~,~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}+g_{0}^{3}h_{0}~~,~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=2g_{0}^{4}~~,~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}~.$ (69)
-The UV divergences are again computed by TwoPoint command in HyperlogProcedures, which are
+UV divergences are again computed by TwoPoint command in HyperlogProcedures, which are
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{55}{72\epsilon}+\left(\frac{\pi^{2}}{72}-\frac{1759}{864}\right)+\epsilon\left(\frac{\zeta_{3}}{36}+\frac{55\pi^{2}}{864}-\frac{42487}{10368}\right)+\mathcal{O}(\epsilon^{2})~,$ (70)
@@
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon}+\left(\frac{71}{24}-\zeta_{3}\right)+\epsilon\left(-\frac{\pi^{4}}{120}-\frac{13\zeta_{3}}{3}-\frac{\pi^{2}}{24}+\frac{2807}{288}\right)+\mathcal{O}(\epsilon^{2})~.$ (76)
-Then the required UV contribution of all 2-loop irreducible topologies for $\varphi^{2}\to 2\varphi$ is
+The two-loop irreducible UV contribution for $\varphi^{2}\to 2\varphi$ is
 $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})=\sum_{L={\tiny\mbox{L2-I}},{\tiny\mbox{L2-II}}_{1},{\tiny\mbox{L2-II}}_{2},\atop{\tiny\mbox{L2-III}}_{1},{\tiny\mbox{L2-III}}_{2},{\tiny\mbox{L2-IX}},{\tiny\mbox{L2-X}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})~.$ (77)
-After the series expansion of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, we get the UV finiteness conditions for the vanishing of $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term as,
+Expanding ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, the UV finiteness conditions, from canceling $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term, are
 $0=z_{\varphi^{2}}^{21}+\left(2g^{2}+1\right)z_{\varphi^{2}}^{11}+\left(g^{2}-\frac{7g^{4}N}{144}+\frac{193g^{4}}{72}+\frac{g^{3}h}{4}-\frac{7g^{2}h^{2}}{144}\right)~,$ (78)
 $0=z_{\varphi^{2}}^{22}+g^{2}z_{\varphi^{2}}^{11}+\left(\frac{g^{4}N}{12}-\frac{g^{4}}{6}-g^{3}h+\frac{g^{2}h^{2}}{12}\right)~.$ (79)
-The $z_{\varphi^{2}}^{11}$ has already been determined in the 1-loop computation, thus by solving above two equations we get,
+Using $z_{\varphi^{2}}^{11}$, solving these equations yields
 $z_{\varphi^{2}}^{21}=\frac{7g^{4}N}{144}-\frac{49g^{4}}{72}-\frac{g^{3}h}{4}+\frac{7g^{2}h^{2}}{144}~~~,~~~z_{\varphi^{2}}^{22}=-\frac{g^{4}N}{12}+\frac{7g^{4}}{6}+g^{3}h-\frac{g^{2}h^{2}}{12}~.$ (80)
-Note that the vanishing of $1/\epsilon$ pole in the sub-Log term provides an additional equation $0=g^{2}z_{\varphi^{2}}^{11}+g^{4}$.
-Hence even if we did not compute $z_{\varphi^{2}}^{11}$ at 1-loop, equations of leading Log and sub-Log terms are sufficient to solve $z_{\varphi^{2}}^{11},z_{\varphi^{2}}^{21},z_{\varphi^{2}}^{22}$, and completely determine the 2-loop $\widetilde{Z}_{\varphi^{2}}$-factor.
-The determination of $\widetilde{Z}_{\varphi^{3}}$-factor follows exactly the same procedure, where now we should use the UV finiteness conditions of $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$.
-The topologies of $\varphi^{3}$-$3\varphi$ correlation function are sketched in Fig.(5).
+Notably, the sub-Log term’s $1/\epsilon$ pole cancelation provides an additional equation $0=g^{2}z_{\varphi^{2}}^{11}+g^{4}$, enabling full determination of $z_{\varphi^{2}}^{11},z_{\varphi^{2}}^{21},z_{\varphi^{2}}^{22}$ even without prior one-loop computations.
+The $\widetilde{Z}_{\varphi^{3}}$-factor is determined via UV finiteness of $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$.
+The $\varphi^{3}$-$3\varphi$ correlation function topologies, as shown in Fig.(5),
 $\Gamma^{\tiny\mbox{L0}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-XII}_{1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-XII}_{2}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-XII}_{3}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$
@@
 A gray arrow points to the cut graphs of each new irreducible topology, which are generated by the non-isomorphic computation of all possible removals of two external legs.
-Notice that the loop topologies in the first row are equivalent to the $\varphi^{2}\to 2\varphi$ irreducible topologies we have already computed.
-There are only 2 new irreducible topologies $\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi{\varphi}\varphi}$, $\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi{\varphi}\varphi}$ at 2-loops for $\varphi^{3}\to 3\varphi$.
-Again from eqn.(35) we get the explicit expression for $\varphi^{3}$ as
+include irreducible topologies from $\varphi^{2}\to 2\varphi$, plus two new two-loop irreducible topologies $\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi{\varphi}\varphi}$ and $\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi{\varphi}\varphi}$.
+From eqn.(35), we get
 $G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})=1+C_{2}^{1}\sum_{j=1}^{L}G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})+(2!)C_{2}^{2}\sum_{j=2}^{L}G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{3})~.$ (81)
-Since $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{1-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ and $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ have already been computed in $\varphi^{2}$ case as (66) and (77), in order to compute 2-loop $\widetilde{Z}_{\varphi^{3}}$-factor, we only need to compute the additional contributions from these 2 new irreducible topologies of $\varphi^{3}\to 3\varphi$.
-They generate in total 4 cut graphs as shown in Fig.(5).
-Their symmetry factors and coupling factors are given by
+Since $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{1-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ and $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ have already been computed in $\varphi^{2}$ case as (66) and (77), only the new two-loop irreducible topologies require computation.
+These generate 4 cut graphs, with symmetry factors and coupling factors as,
 $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~~~,~~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{3}h_{0}~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}~.$ (82)
-The UV divergences computed by TwoPoint command are,
+UV divergences are,
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon}+\frac{17}{8}+\epsilon\left(\frac{171}{32}-\frac{\pi^{2}}{24}\right)+\mathcal{O}(\epsilon^{2})~,$ (83)
@@
 $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XII}_{3}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{171}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(-\frac{\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{1313}{128}\right)+\mathcal{O}(\epsilon^{2})~.$ (86)
-Then the required UV contribution of all 2-loop irreducible topologies for $\varphi^{3}\to 3\varphi$ is
+Then the required UV contribution of 2-loop irreducible topologies for $\varphi^{3}\to 3\varphi$ is
 $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{3})=\sum_{L={\tiny\mbox{L2-XI}},{\tiny\mbox{L2-XII}}_{1}\atop{\tiny\mbox{L2-XII}}_{2},{\tiny\mbox{L2-XII}}_{3}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})~.$ (87)
-With the previously computed $\widetilde{Z}_{\varphi^{2}}$-factor, we can take the series expansion for $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$, and get the following equations from the vanishing of $1/\epsilon$, $1/\epsilon^{2}$ poles in the leading Log term and $1/\epsilon$ pole in the sub-Log term,
+Expanding $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$, the UV finiteness conditions, from canceling $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term, and $1/\epsilon$ pole in the sub-Log term, give three equations,
 $0=z_{\varphi^{3}}^{21}+(4g^{2}+1)z_{\varphi^{3}}^{11}+\left(3g^{2}-\frac{7g^{4}N}{48}+\frac{355g^{4}}{24}+\frac{5g^{3}h}{4}-\frac{7g^{2}h^{2}}{48}\right)~,$ (88)
 $0=z_{\varphi^{3}}^{22}+3g^{2}z_{\varphi^{3}}^{11}+\left(\frac{g^{4}N}{4}+\frac{5g^{4}}{2}-3g^{3}h+\frac{g^{2}h^{2}}{4}\right)~~~,~~~0=2g^{2}z_{\varphi^{3}}^{11}+6g^{4}~.$ (89)
-By solving these 3 equations we completely determine the 2-loop $\widetilde{Z}_{\varphi^{3}}$-factor as
+Solving these yields,
 $z_{\varphi^{3}}^{11}=-3g^{2}~~,~~z_{\varphi^{3}}^{21}=\frac{7g^{4}N}{48}-\frac{67g^{4}}{24}-\frac{5g^{3}h}{4}+\frac{7g^{2}h^{2}}{48}~~,~~z_{\varphi^{3}}^{22}=-\frac{g^{4}N}{4}+\frac{13g^{4}}{2}+3g^{3}h-\frac{g^{2}h^{2}}{4}~.$ (90)
-With results (67), (80) and (90), the 2-loop $\widetilde{Z}_{\varphi^{2}}$, $\widetilde{Z}_{\varphi^{3}}$-factors are finally determined as,
+With results (67), (80) and (90), the two-loop $\widetilde{Z}_{\varphi^{2}}$, $\widetilde{Z}_{\varphi^{3}}$-factors are finally determined as,
 $\widetilde{Z}_{\varphi^{2}}=1+\frac{z_{\varphi^{2}}^{11}}{\epsilon}+\frac{z_{\varphi^{2}}^{21}}{\epsilon}+\frac{z_{\varphi^{2}}^{22}}{\epsilon^{2}}~~~,~~~\widetilde{Z}_{\varphi^{3}}=1+\frac{z_{\varphi^{3}}^{11}}{\epsilon}+\frac{z_{\varphi^{3}}^{21}}{\epsilon}+\frac{z_{\varphi^{3}}^{22}}{\epsilon^{2}}~.$ (91)
-The gamma functions of $\varphi^{Q}$ can be computed by $\gamma_{\varphi^{Q}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}$, hence we get
-$\gamma_{\varphi^{2}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial h}\beta(h)=-g^{2}+\left(\frac{7g^{4}N}{72}-\frac{49g^{4}}{36}-\frac{g^{3}h}{2}+\frac{7g^{2}h^{2}}{72}\right)~,$ (92)
-$\gamma_{\varphi^{3}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial h}\beta(h)=-3g^{2}+\left(\frac{7g^{4}N}{24}-\frac{67g^{4}}{12}-\frac{5g^{3}h}{2}+\frac{7g^{2}h^{2}}{24}\right)~,$ (93)
-where $\beta(g),\beta(h)$ have been provided in (61), (62). By the ansatz that the 2-loop $\gamma_{\varphi^{Q}}$ takes the form,
-$\gamma_{\varphi^{Q}}=c_{1}Q(Q-1)(Q-2)+c_{2}Q(Q-1)~,$ (94)
-and from results $\gamma_{\varphi^{2}},\gamma_{\varphi^{3}}$ we determine
+The $\gamma_{Q}:=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}$ are computed using $\beta(g)$ and $\beta(h)$ from eqns.(61) and (62),
+$\gamma_{2}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial h}\beta(h)=-g^{2}+\left(\frac{7g^{4}N}{72}-\frac{49g^{4}}{36}-\frac{g^{3}h}{2}+\frac{7g^{2}h^{2}}{72}\right)~,$ (92)
+$\gamma_{3}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial h}\beta(h)=-3g^{2}+\left(\frac{7g^{4}N}{24}-\frac{67g^{4}}{12}-\frac{5g^{3}h}{2}+\frac{7g^{2}h^{2}}{24}\right)~.$ (93)
+Assuming $\gamma_{Q}$ follows the ansatz $\gamma_{Q}=c_{1}Q(Q-1)(Q-2)+c_{2}Q(Q-1)$, we determine
 $c_{1}=-\frac{g^{4}}{4}-\frac{g^{3}h}{6}~~~,~~~c_{2}=-\frac{g^{2}}{2}+\frac{7g^{4}N}{144}-\frac{49g^{4}}{72}-\frac{g^{3}h}{4}+\frac{7g^{2}h^{2}}{144}~.$ (94)
-Then the scaling dimension of the $\varphi^{Q}$ operator can be calculated by its definition as,
-$\Delta_{Q}=\left(\frac{d}{2}-1\right)Q+\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}~,$ (96)
-where the anomalous dimension of the field $\varphi$ has been computed in (63).
+Then the scaling dimension of the $\varphi^{Q}$ operator is
+$\Delta_{Q}=\left(\frac{d}{2}-1\right)Q+\gamma_{Q}+Q\gamma_{\varphi}~,$ (95)
+where $\gamma_{\varphi}$ is the field anomalous dimension from eqn.(63).
 4 Five-loop anomalous dimensions
-The procedures described in previous section for 2-loop computations can be straightforwardly generalized to higher-loop calculations following exactly the same manner, with the primary difference being the need to consider a larger number of graphs.
-The $Z_{\varphi},Z_{\sigma}$-factors are determined from their corresponding two-point topologies in the manner analogous to (41).
-Similarly, the $Z_{h},Z_{g}$-factors are determined from the two-point integrals of cut graphs, generated by removing one external leg from the respective three-point topologies, following the approach outlined in (55).
-The beta functions are then computed from $\widetilde{Z}_{g},\widetilde{Z}_{h}$ using the definition (9).
-The field anomalous dimensions $\gamma_{\varphi}$ and $\gamma_{\sigma}$ can be computed from the $Z_{\varphi},Z_{\sigma}$-factors combined with the beta functions, following the definition (10).
-The $\widetilde{Z}_{\varphi^{Q}}$ factor is determined from the two-point integrals of cut graphs generated by removing (Q1)(Q-1) external legs from irreducible 1PI $\varphi^{Q}\to Q\varphi$ topologies, in a manner similar to that described in eqn.(35).
-The anomalous dimension of the $\varphi^{Q}$ operator is then given by $\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}$, where $\gamma_{\varphi^{Q}}$ is computed from $\widetilde{Z}_{\varphi^{Q}}$ and the beta functions by its definition.
-Determining the 5-loop $\gamma_{\varphi^{Q}}$ for general $Q$ requires the results of $L$-loop $\gamma_{\varphi^{Q_{L}}}$ with $Q_{L}=2,\ldots,L+1$ up to 5-loops.
-Since UV divergences are computed from scalar integrals, we do not need to consider the specific field configurations of Feynman diagrams, but only their topologies.
-All topologies for two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions, as well as their cut graphs are generated by private Mathematica codes.
-The number of 1PI topologies for two-point and three-point functions, and irreducible correlation functions is listed in Table.(2).
+The two-loop computation procedures outlined in the preceding section can be straightforwardly generalized to higher-loop calculations following an identical framework.
+The primary distinction lies in the need to account for a substantially larger number of Feynman diagrams, as topological complexity increases with loop order.
+The field renormalization factors $Z_{\varphi}$ and $Z_{\sigma}$ are determined from their respective two-point topologies, following the approach analogous to (41).
+Similarly, the $Z_{h}$ and $Z_{g}$-factors are derived from two-point integrals of cut graphs, generated by removing one external leg from the respective three-point topologies, following the approach outlined in (55).
+For the composite operator renormalization factors $\widetilde{Z}_{\varphi^{Q}}$, these are determined from two-point integrals of cut graphs produced by removing (Q1)(Q-1) external legs from irreducible 1PI $\varphi^{Q}\to Q\varphi$ topologies, in a manner parallel to that described in eqn.(35).
+The anomalous dimension $\gamma_{\varphi^{Q}}$ of the $\varphi^{Q}$ operator is then given by $\gamma_{Q}+Q\gamma_{\varphi}$, where $\gamma_{Q}$ is computed from $\widetilde{Z}_{\varphi^{Q}}$ and the beta functions.
+Determining the 5-loop $\gamma_{\varphi^{Q}}$ for general $Q$ requires results of $L$-loop $\gamma_{{Q_{L}}}$ with $Q_{L}=2,\ldots,L+1$ up to 5-loops.
+Notably, since UV divergences are evaluated from scalar Feynman integrals, the specific internal field configurations of the diagrams need not be considered, only their topological structures are relevant.
+All topologies for two-point, three-point and $\varphi^{Q}$-$Q\varphi$ correlation functions, along with their corresponding cut graphs, are generated using custom Mathematica codes.
+The number of 1PI topologies for two-point, three-point functions and the irreducible correlation functions is listed in Table.(2).
 $Q=2$ $Q=3$ $Q=4$ $Q=5$ $Q=6$ Total 2-point 3-point
@@
 These are the graphs to be computed by TwoPoint command for UV divergences.
-The number grows rapidly as the number of loops increases.
-Cut graphs are generated through non-isomorphic computations of all possible external leg removals, and each of these cut graphs should be computed to determine their UV divergences.
-The number of cut graphs is listed in Table.(2).
-It can be seen that the number of cut graphs to be computed is about 3 to 4 times larger than the number of the original topologies.
-The UV divergences are computed via TwoPoint command in HyperlogProcedures Maple package.
-The computational capability and time consumption of these evaluations are sensitive to both the loop order and the order of the dimensional regulator $\epsilon$.
-To apply the UV finiteness conditions for $L_{\tiny\mbox{max}}$-loop computation, an $L$-loop divergence in dimensional regularization should be computed up to the $\epsilon^{L_{\tiny\mbox{max}}-1-L}$ order.
-Increasing $L_{\tiny\mbox{max}}$ by one order requires UV divergence computations with an additional order in $\epsilon$ for all loop integrals, which significantly increases time consumption.
-For 5-loop computations, an $L$-loop divergence should be computed to the $\epsilon^{4-L}$ order, and the current version of HyperlogProcedures is capable of evaluating all required two-point integrals.
-Time consumption varies depending on the computer’s CPU and memory usage.
-On a normal desktop with a CPU base frequency 3.20GHz and 64GB memory, the approximate time required to evaluate all loop integrals is listed in Table.(3).
+The number grows rapidly with increasing loop order.
+Cut graphs are generated via non-isomorphic enumeration of all possible external leg removals, and each cut graph must be individually evaluated to extract its UV divergence.
+As shown in the Table.(2), the number of cut graphs to be computed is approximately 3–4 times larger than the number of original topologies.
+UV divergences are calculated using TwoPoint command in the HyperlogProcedures Maple package.
+The computational feasibility and time cost of these evaluations are highly sensitive to both the loop order and the truncation order of the dimensional regulator $\epsilon$.
+For an $L_{\tiny\mbox{max}}$-loop renormalization computation, each $L$-loop divergence in dimensional regularization must be computed up to the $\epsilon^{L_{\tiny\mbox{max}}-1-L}$ order.
+Increasing $L_{\tiny\mbox{max}}$ by one loop requires extending the $\epsilon$ expansion of all loop integrals by an additional order, which substantially increases computational time.
+For 5-loop computations, each $L$-loop divergence must be evaluated to the $\epsilon^{4-L}$ order, and the current version of HyperlogProcedures is capable of evaluating all required two-point integrals.
+Time consumption varies depending on the computer’s CPU performance and memory capacity.
+On a standard desktop workstation equipped with a 3.20GHz base frequency CPU and 64GB of RAM, the approximate time required to evaluate all necessary loop integrals is summarized in Table.(3).
 $Q=2$ $Q=3$ $Q=4$ $Q=5$ $Q=6$ Total 2-point 3-point
@@
 The working time less than 0.01 hour is simply denoted by 0.
-Time consumption also increases rapidly as the loop order rises.
+Time consumption also scales rapidly with increasing loop order.
 However, for 5-loop computations, it remains within a reasonably acceptable range.
-In contrast, the symmetry factors and coupling factors for these topologies are generated almost instantaneously compared to the time required for evaluating UV divergences.
-We have computed $Z_{\varphi}$, $Z_{\sigma}$, $Z_{h}$, $Z_{g}$ and all $\widetilde{Z}_{\varphi^{Q_{L}}},Q_{L}=2,\ldots,L+1$ up to 5-loops.
-From these results, we have derived the beta functions $\beta(g),\beta(h)$, the field anomalous dimensions $\gamma_{\varphi}$, $\gamma_{\sigma}$, and the scaling dimensions of the $\varphi^{Q}$} operator up to 5-loops.
-Results of lower than 5-loops have been compared with existing literature with agreements Fei:2014xta ; Gracey:2015tta ; Kompaniets:2021hwg .
-The 5-loop correction to the anomalous dimension $\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}$ is presented in Appendix §A.
-The scaling dimension of the $\varphi^{Q}$ operator at the Wilson-Fisher fixed point can be computed as
-$\Delta_{Q\tiny\mbox{FP}}=\left(2-\frac{\epsilon}{2}\right)Q+\left(\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}\right)\Big{|}_{g=g^{\ast},h=h^{\ast}}~,$ (97)
+In contrast, the generation of symmetry factors and coupling factors for these topologies is nearly instantaneous, with negligible computational overhead compared to the time-intensive evaluation of UV divergences.
+We have computed the field renormalization factors $Z_{\varphi}$, $Z_{\sigma}$, coupling renormalization factors $Z_{h}$, $Z_{g}$, and all composite operator renormalization factors $\widetilde{Z}_{\varphi^{Q_{L}}},Q_{L}=2,\ldots,L+1$ up to 5-loops.
+From these results, we have derived the beta functions $\beta(g),\beta(h)$, field anomalous dimensions $\gamma_{\varphi}$, $\gamma_{\sigma}$, and scaling dimensions of the $\varphi^{Q}$} operator, all up to 5-loops.
+These results are provided in ancillary files to the arXiv, and we have cross-checked them against existing literature with excellent agreements.
+For example, the up to 3-loop pieces of $\gamma_{\varphi},\gamma_{\sigma},\beta_{g},\beta_{h}$ are identical to those in Fei:2014xta , after unifying possible overall factors $(4\pi)$ from different conventions.
+The up to 4-loop pieces of results are identical to those in the $O(N)$ sector in Gracey:2015tta , after unifying possible overall factors $2$ from different conventions, and the fixed point solutions of beta functions up to $\epsilon^{4}$ order are identical to those in Kompaniets:2021hwg after compensating an overall $\sqrt{2}$ factor due to different convention.
+The 5-loop correction to the anomalous dimension $\gamma_{Q}+Q\gamma_{\varphi}$ is presented in Appendix §A.
+The scaling dimension of the $\varphi^{Q}$ operator at the Wilson-Fisher fixed point is computed as
+$\Delta_{Q\tiny\mbox{FP}}=\left(2-\frac{\epsilon}{2}\right)Q+\left(\gamma_{Q}+Q\gamma_{\varphi}\right)\Big|_{g=g^{\ast},h=h^{\ast}}~,$ (96)
 where $g^{\ast},h^{\ast}$ are the fixed-point solutions of the beta functions.
-Our 5-loop fixed-point solutions, provided in Appendix §B, are in perfect agreement with the known results in the literature Kompaniets:2021hwg .
-By substituting these fixed-point solutions into the anomalous dimensions, we obtain the $\Delta_{Q}$ in the large $N$ limit.
+Our 5-loop fixed point solutions, provided in Appendix §B, are in perfect agreement with the literature Kompaniets:2021hwg .
+By substituting these fixed-point solutions into the anomalous dimensions yields $\Delta_{Q}$ in the large $N$ limit.
 The leading, sub-leading and sub-sub-leading order contributions of the large $N$ expansion are as follows,
@@
 $~~~~~~~~~~~~~~~~~~~~~~+\bigg(3348\zeta_{3}^{2}+\frac{549\zeta_{3}}{2}+3663\zeta_{5}-\frac{10\pi^{6}}{7}-\frac{3093\pi^{4}}{20}-\frac{190189}{24}\bigg)Q^{2}$
-$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\bigg{(}\frac{7551}{8}-702\zeta_{3}^{2}-2322\zeta_{3}-1215\zeta_{5}+\frac{5\pi^{6}}{7}+\frac{72\pi^{4}}{5}\bigg{)}Q^{3}\bigg{]}\epsilon^{5}~,$
-and expressions $\Delta_{Q{\tiny\mbox{FP}}}^{4},\Delta_{Q{\tiny\mbox{FP}}}^{5}$ are presented in Appendix §C.
-As a confirmation of these results, the 3-loop contribution up to $\epsilon^{3}$ order is consistent with literature findings Vasiliev:1981yc ; Vasiliev:1981dg ; Vasiliev:1982dc ; Jack:2021ziq  [^7] The value $a_{t}^{(3)}$ in paper Jack:2021ziq that computes $\Delta_{Q}^{(3)}$ is incorrect.
+$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\bigg(\frac{7551}{8}-702\zeta_{3}^{2}-2322\zeta_{3}-1215\zeta_{5}+\frac{5\pi^{6}}{7}+\frac{72\pi^{4}}{5}\bigg)Q^{3}\bigg]\epsilon^{5}~.$
+Expressions $\Delta_{Q{\tiny\mbox{FP}}}^{4}$ and $\Delta_{Q{\tiny\mbox{FP}}}^{5}$ are presented in Appendix §C.
+These results are further validated by the following consistency checks.
+The 3-loop contribution up to $\epsilon^{3}$ order is consistent with literature findings Vasiliev:1981yc ; Vasiliev:1981dg ; Vasiliev:1982dc ; Jack:2021ziq  [^8] The value $a_{t}^{(3)}$ in paper Jack:2021ziq that computes $\Delta_{Q}^{(3)}$ is incorrect.
 After using the correct value $a_{t}^{(3)}=\frac{g^{6}}{3}+g^{5}h$, the $\Delta_{Q}^{(3)}$ and the $1/N^{3}$ term in their paper are consistent with the 3-loop order part of our results..
@@
 The 4-loop contribution up to $1/N^{2}$ order has been previously provided in Huang:2024hsn , while the 5-loop contributions up to $1/N^{5}$ order presented here are new.
-Our results have been provided in ancillary files to the arXiv submission of this article.
+All the above results have been provided in ancillary files to the arXiv submission of this article.
 5 Conclusion
-In this paper, we apply the OPE method of renormalization proposed in Huang:2024hsn to evaluate the 5-loop scaling dimensions of the $\phi^{Q}$ operator in the six-dimensional cubic scalar theory.
-The result is derived from over 40 000 two-point propagator-type integrals computed using the graphical function method, combined with non-trivial summation among these integrals.
-We have successfully obtained the 5-loop correction to the anomalous dimension of the $\phi^{Q}$ operator, as well as the large $N$ expansion of the scaling dimension at the Wilson-Fisher fixed point up to the $1/N^{5}$ order.
-It stands for a new record for the scaling dimensions of $\phi^{Q}$ operator in cubic scalar theory.
-The result shows consistency with lower-loop perturbative calculations reported in other literature and aligns with the corresponding terms derived via semi-classical methods.
-Furthermore, the non-trivial cancellation of $\epsilon$ poles in the final result provides a strong consistency check validating its correctness.
-This work further demonstrates the efficiency and power of the OPE method in renormalization computations.
-Efforts to apply the OPE method to address the challenges posed by more complex problems are currently underway.
-A promising project aims to surpass the 6-loop record in quartic scalar theory by computing the 7-loop scaling dimensions of the $\phi^{Q}$ operator.
-Such research would not only advance the precision of renormalization computations but also drive the evolution of modern techniques for multi-loop integral evaluations.
-Also as previously noted, the OPE method is highly general and extendable beyond scalar theories.
-It is interesting to investigate the renormalization computations by OPE method in other situations, such as the Gross-Neveu-Yukawa theory with fermions, or scalar-QED theory with gauge fields, etc.
-In these theories, Feynman integrals are no longer purely scalar integrals, but involve non-trivial numerators.
-While further developments may enable the computation of such non-trivial master integrals via graphical function method adapted to numerator structures, the most viable current approach relies on the computational algorithms Smirnov:2008iw ; Maierhofer:2017gsa ; Wu:2023upw ; Guan:2024byi based on the IBP method Tkachov:1981wb ; Chetyrkin:1981qh ; Laporta:2000dsw .
-Another potential challenge would be applying the OPE method to the anomalous dimensions of more intricate operators, such as those with higher derivatives, Lorentz indices, or in scenarios with operator mixing Derkachov:1997ch ; Cao:2021cdt .
-While the application of the OPE method remains straightforward in these cases Huang:2024hsn , practical computations demand more sophisticated techniques.
-It is interesting to explore the diverse applications of the OPE method in renormalization, and further enhance its capabilities.
+In this paper, we employ the OPE based renormalization algorithm proposed in Huang:2024hsn to compute the 5-loop scaling dimensions of the $\phi^{Q}$ operator in the six-dimensional cubic scalar theory.
+This algorithm represents a non-trivial implementation of the connection between Wilson coefficients of OPE and anomalous dimensions, enabling the determination of renormalization $Z$-factors by treating UV divergences globally.
+A key advantage is its avoidance of cumbersome sub-divergence subtraction.
+Instead, it systematically extracts the essential UV divergences for renormalization directly from two-point integrals.
+Advancements in modern computational techniques, particularly the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP adopted herein, have rendered feasible the evaluation of these two-point integrals even beyond five loops.
+The transformative impact of the graphical function method, combined with the efficiency and robustness of the proposed OPE based algorithm, has facilitated the successful completion of renormalization computations up to five loops on a standard desktop computer.
+These results are derived from over 40,000 two-point propagator-type integrals computed using the HyperlogProcedures package HP , augmented by non-trivial summation protocols inherent to the algorithm.
+We have successfully obtained the 5-loop corrections to the anomalous dimension of the $\phi^{Q}$ operator, as well as the large $N$ expansion of the scaling dimension at the Wilson-Fisher fixed point up to $1/N^{5}$ order.
+It stands for a new record for the scaling dimensions of the $\phi^{Q}$ operator in cubic scalar theory.
+Our results exhibit perfect consistency with lower-loop perturbative calculations reported in the literature and align with corresponding terms derived via semi-classical methods.
+Furthermore, the non-trivial cancellation of $\epsilon$ poles in the final result provides a stringent consistency check, validating its correctness.
+This work further demonstrates the efficiency and versatility of the OPE based algorithm for high-precision renormalization computations in quantum field theories.
+Efforts are currently underway to apply this OPE based algorithm to tackle the challenges presented by more sophisticated problems.
+One promising initiative aims to break the existing 6-loop benchmark in quartic scalar theory by computing the 7-loop scaling dimensions of the $\phi^{Q}$ operator.
+Such research will not only push the precision of renormalization computations but also drive the evolution of modern techniques for multi-loop integral evaluations.
+Also as previously noted, the OPE based algorithm boasts strong generality and extensibility beyond scalar field theories.
+It would be particularly insightful to explore its application to renormalization computations in other theoretical frameworks, such as the Gross-Neveu-Yukawa theory and scalar QED, among others.
+In these theories, Feynman integrals are no longer purely scalar in nature, but instead involve non-trivial numerator structures.
+While future methodological refinements may enable the calculation of such non-trivial master integrals via a modified graphical function method tailored to handle numerator-dependent topologies, the most robust approach available at present relies on computational algorithms Smirnov:2008iw ; Maierhofer:2017gsa ; Wu:2023upw ; Guan:2024byi based on the integration-by-parts (IBP) method Tkachov:1981wb ; Chetyrkin:1981qh ; Laporta:2000dsw .
+Another potential challenge lies in extending this algorithm to compute the anomalous dimensions of more intricate operators, for instance, those containing higher derivatives, Lorentz indices, or in scenarios involving operator mixing Derkachov:1997ch ; Cao:2021cdt .
+Although the conceptual application of the OPE based algorithm remains straightforward in these cases, as demonstrated in Huang:2024hsn , practical computations necessitate more advanced technical implementations.
+Further exploration of the algorithm’s diverse applications in renormalization and targeted enhancements of its computational capabilities will undoubtedly open new avenues for high-precision quantum field theory calculations.
 Acknowledgments
-We would like to thank Oliver Schnetz for stimulating conversations on the graphical function method and the usage of HyperlogProcedures package.
+We would like to thank Oliver Schnetz for stimulating conversations on the graphical function method and the usage of HyperlogProcedures package, and Hugh Osborn for useful comments on this paper.
 This work is supported by the Science Challenge Project (No.
@@
 Appendix C Large $N$ expansion at the Wilson-Fisher fixed point
-Here we present the results of the $1/N^{4}$ and $1/N^{5}$ orders in the large $N$ expansion of $\gamma_{Q}$ at Wilson-Fisher fixed point. We have,
+Here we present the results of the $1/N^{4}$ and $1/N^{5}$ orders in the large $N$ expansion of $\gamma_{\varphi^{Q}}$ at Wilson-Fisher fixed point. We have,
 $\Delta_{Q{\tiny\mbox{FP}}}^{4}$ $=$ $\bigg[-2481664+1861248Q\bigg]\epsilon+\bigg[-\frac{21069760}{3}+8828264Q-3043440Q^{2}\bigg]\epsilon^{2}+\bigg[\frac{259232542}{9}-861840\zeta_{3}+\bigg(-30411798$ (113)